A Complete Bibliography of the Journal of Econometrics (2020–2029)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

28 May 2024 Version 1.30

Title word cross-reference

2 [RW24]. α [LTYZ23]. b [Cas24]. β [BR22]. F [MISW20, PS23a]. g [BC21]. I(1) [BLL21]. ∞ [Roy23]. K [ZL23]. L_{2} [KLSW23]. L_{p} [BKW23]. N [YCK20, CZ20]. p [Kli24, ZD24]. q [CK20]. R [HL20a]. t [HM24, MO23].

-boosting [KLSW23]. -estimation [HL20a]. -fold [ZL23]. -order [BC21]. -prediction [CZ20]. -ratios [MO23]. -stable [LTYZ23]. -values [Kli24, ZD24].

164 [PST22]. 19 [CKS21, GLLZ23, HLM23, HLS21, JZS23, JSY23, KN21, KV23, Kor21, LLSS21, LL21, LMS21, MM21a, PST23, Tam21, Tou21].

2020 [Ano20t, Ano20s, Ano20u, Ano20o, Ano20q, Ano20n, Ano20r, Ano20p]. 2021 [Ano21s, Ano21u, Ano21p, Ano21q, Ano21t, Ano21r]. 2022 [Ano22r, Ano22p, Ano22o, Ano22n, Ano22m, Ano22q]. 2023 [Ano23b, Ano23o, Ano23p, Ano23w, Ano23v, Ano23u, Ano23q, Ano23x, Ano23r, Ano23s, Ano23t, Ano23y, Ano24d, Ano24e, Ano24f]. 2024 [Ano24a, Ano24b, Ano24r, Ano24s, Ano24t, Ano24u, Ano24v]. 212 [CCCM22]. 230 [AACH23]. 2SIF [KS20].

85th [CT20].

ability [CFG24]. abortions [JS24]. Academic [DDH22, BF22]. accumulation [BFLT21]. accuracy [BP20a]. accurate [LMSND22]. ACD [CMRV24]. achievement [BF22]. across [FJ22]. actions [GSV22]. activity [ADDP24, BAFMS20, NP22]. Adaptive [JLZ21, LR20, NP22, SKF23, BG23, CL24b, JPTZ23]. adaptivity [GCT23]. Adjusted [BDSV23, MH20, FKW20, FJS22, HLM^{+}24, JPTZ23, ZD21]. adjusted-range [HLM^{+}24]. adjustment [BJR^{+}24]. administrative [LMSW23]. admissions [ZZ21]. adoption [AI22]. Advances [ADDP24]. Adversarial [AIMM24]. adverse [PG21]. Affine [HNZ22]. after [AKM21, CSV23]. against [HJG23, HHS20]. age [LTZ21, Sch23]. aggregate [GJ23, LST23]. aggregation [DSZ24, GM21, LP22]. ain’t [DKSS23]. Akaike [JM21]. algorithm [CLWY24]. allocation [KW23, PT11, PST22]. alpha [FLLM22]. ALQR [FLS24]. alternatives [FLLM22, HJG23]. ambiguity [BKS22]. America [MGSN24]. among [JS24, KSW24]. Analysis [JLMM21, AHX21, AHZ23, ACL22, ABL22, AI22, Bot20, CMV23, CL24b, CHK22, Che24, CWXZ24, CY22, CK23b, EF24, GHKP21, GK24, Ho23, HT20, HJLP23, JZS23, KHK20, PY20, SW21a, SKY^{+}21, SWZZ24, TPZ23, WY21, WLFF24, ZZ24b]. analytical [NW21]. Analyzing [HW22, SX22]. Anatomy [MM21a]. angular [JC20]. Annals [DvdKWZ22, GH22, NQV21, CT20]. Announcement [Ano20a, Ano23a]. announcements [GSV22]. Annual [Ano20a]. anticipation [DM22]. Application [HS24, BBGS24, BMP23, BK20b, Cal21, CCM21, CNPR22, CYX^{+}23, CKL24, CFG23, DS20, DMP22, DM22, FMM^{+}22, GHM20, GM21, GJ23, Gua21a, GLLZ23, GLLZ24, HKT20, HM24, HL20b, HNZ22, HLS21, JPS24, Kam24, Kim23, KHK20, KZ20, LL22, LH24, NP22, RW24, Wan22, YCK20, ZM24]. Applications [SCC22, BM20, BKW23, FFX20, FKL21, KKIS21, KOEP20, Win24, XP23]. approach [Bai24, BAFMS20, BCPV23, BK20a, CGV22, DHS24, EFS23, FLS24, FH22, FGP22a, HY24, HHO22, HLM^{+}24, HLLO21, HV20, HL23, HKNW23, JP24, JS22, KPS24, KHK20, MZ21, NW21, RC23, SX22, SH23, SW21b, SG21, Tou21, YS21, ZGZ24, vdBJMN21]. approaches [FS21]. appropriate [MNW23b]. Approximate [BN23a, CFR22]. approximately [Kli24]. Approximation [Kno22, MJLS20]. Approximations [EL21, ASLL24]. April [Ano20t, Ano21s, Ano22r, Ano23w, Ano24a, Ano24b]. arbitrage [CCL21, HNZ22]. arbitrage-free [HNZ22]. ARCH [Roy23]. Arctic [DRG^{+}24, DR22b, DRG^{+}23]. area [DRG^{+}23, DRG^{+}24]. Artificial [TPZ23]. ascending [BGM21]. Assessing [WML21]. assessment [AAG20]. assessments [DR22b, DSZ24]. Asset [CLWY24, FFL24, WCLC22, BT24, CFX22, Dal23, GM21, GKX21, LP20a, PT11, PST22, SX22, SG21, Uml24, Win24, YYX24]. asset-pricing [LP20a]. assets [AM22]. assignments [OP24]. Assumption [LMR24, Gof24, KY22]. Assumption-lean [LMR24]. assumptions [AFiK^{+}23, ACS20]. asymmetric [HHLS22, WZLL22]. asymmetry [ACM22, Roy23]. Asymptotic [CY22, DGR20, MISW20, MZ21, PS23a, SWP20, Shi23, ACG20, GGV20]. Asymptotically [Pet22, CL23c]. asynchronous [Che24, CMZ24]. attachment [CCW20]. attention [FLS22]. attitude [LH24]. attrition [GHKOB24]. auction [EFS23, GG20b, JP24, LPV23, LX23]. auction-determined [LPV23]. auctions [BGM21, GG22, GZ23, JZ22, Luo20, Zin24]. audit [YCK20]. Augmented [FKL21, CHL24, DR22a, JLZ20, Wan22]. August [Ano22p, Ano23v]. Australia [IK21]. auto [CFZ23]. auto-regression [CFZ23]. autocorrelation [Cas23]. autocovariance [CCQY24]. autocovariance-based [CCQY24]. Autoencoder [GKX21]. automated [SW21b]. automobile [WY21]. autoregression [GGR24, RLZ^{+}24, YfL21]. Autoregressions [CCM21, Mav21, OSW21, Bog22, CCM19, CCCM22, CGL^{+}22, FS23b, Gua21b, GB21b, Gup23, IK20, KPT23b, MGW23, Pre20, RW24, RR23, RL23]. Autoregressive [BFL24, CXY21, FSU20, FHP23, HT20, HWZW20, JLZ20, JLZ21, KS20, LTYZ23, MH20, MS21a, PDC21, WZL24, ZZLL22, ZHPW20, ZM24, HKST24]. average [CCT23, GdXP22, LMS23, ZD21, Zhu23]. averaging [Boo23, CL23c, FLX22, LZGZ21, PY22, SHL^{+}21, SHWZ23, ZL23, ZZ24a]. aversion [BKS22, JZ22, KRvdK22]. Award [Ano20a]. Awards [Ano23a]. aware [SZ24].

bagging [MGW23]. balance [iSK21, Pre20, ZD24]. bank [GSV22]. based [AFiK^{+}23, AI22, Bei24, BLL24, Boo23, CBN23, CHLZ20, CMV23, CP21, CMRV24, CCQY24, CXW22, Che24, CY22, CLS23, DRT24, FFX20, HLM^{+}24, HL23, KZA20, LPG20, LLZ22, LLYZ22, LZ20, PS21b, RL23, SST21, SKF23, SG21, SXZ23, Uml24, WCLC22, WLFM24, YLC^{+}23, ZZLL22, ZZ24a, ZGZ24, ZD24]. Basket [BDSV23]. Basket-Adjusted [BDSV23]. Bayes [Gal22, OJT20, Tau22]. Bayesian [Bog22, CCCM22, ABL22, ARRW21, BAFMS20, CCM19, Cha23, CGL^{+}22, CK24, DMP22, DTB21, FJS22, FH22, FJ22, FHLL22, GKR22, Ho23, IK22, KZ21, KHK20, Kli24, KS23, MS21b, NP22, NS24, Pet22, Shi23, TPZ23, WFL22, Yu22]. be [HD22]. behavior [CMRV24, CKS21, LPY24]. behavioral [KKKN21]. behaviors [JS24]. Beliefs [GSS22, RSW22, BBRR22, GLWW22, HS21a, KS23, vGW22]. Bellman [Lan24]. benchmark [BDKM23]. benefits [BBRSS23, TZ24]. best [WZZW24]. Beta [GK24, BDKM23]. betas [BFL24, CLY24, ZLTT22]. Better [BMS24]. between [Gua21a, Hor21, ZT22]. Beyond [MB24, BFM23, YZC21]. Bias [HJ24, TZ23, AK24, GHKOB24, PW23, SST23]. bid [JZ22]. Bidding [BGM21]. bids [GG20b]. Big [YCK20]. bilateral [QfLY21]. Binary [GLPY23, LH24, Car23, GR23, Gun23, KPT23a, LTY21, Man23, PF23, Su21, WTH24]. binding [AMSV22]. Bipartite [WLFF24]. bird [HS20]. birthday [CT20]. BLP [HLL21, Wan23]. board [Gua21a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano23c, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano24g, Ano24h, Ano24i, Ano24j, Ano24k, Ano24l, Ano24m, Ano24n, Ano24o, Ano24p, Ano24q]. bond [FKL21, HJL^{+}24, KLL21, WFL22]. bonds [BDKM23, CS22]. Boosting [YN21, KLSW23, YCK20]. Bootstrap [CNPR22, CLRSØ23, HL23, HJLP23, PS21b, PS23b, BHS24, DT20, DS24, FSU20, FL24, GGR24, HV20, LMS23, LP23a, LT21, WZ24]. Bootstrap-based [HL23]. Bootstrapping [BCGR21, GP20, FLZ24]. boundary [CNPR22]. Bounding [Hor21, TZ24]. Bounds [Cal21, AL21, Cha24, Mar24]. Box [ZZ24b]. break [BP20a, CP21, DBH23, DR20, Shi23]. breaks [AKM21, HLM^{+}24, KOEP20, LOW23, MT23a, OW21, PDC21, PT11, PST22]. browser [JLMM21]. bubble [GJM20]. build [AGP20]. bunching [BMS24]. bundled [SZ24]. bundles [AR22]. burst [COR22]. Business [BT24, ABB^{+}22]. Business-cycle [BT24]. buyer [GG20b].

C [CT20]. Can [AMMQ22]. cancer [Kim23]. Canonical [CLS23, Fun24]. capital [BFLT21, CFX22, IK21]. carbon [CMH24, WGH20]. Carlo [AIMM24, FHLL22, KS20]. case [HLT20, HR21, IK21, KV23]. cash [PST23]. catastrophe [BDKM23]. catastrophic [ILMM20]. categorical [HM21]. Causal [CLMZ24, CKS21, DGJOOSB23, GPWY24, Kas22, Sun24, XP23]. Causality [CMPZW20, ATM20, GHM20]. CCE [DS24, JKW21]. Celebrating [SW21a]. celebration [CT20]. Censored [NK22, BHKvD20, CW20, Che23, FVvVV24, WX22]. censoring [CW23]. center [GMM22]. Central [GSV22]. cereal [KHK20]. chain [Gal23]. Chamberlain [GH22]. change [AV22, BHS20, BCS20, CP21, CYZ23, CGGK24, HPP20, JZS23, JPS24, LSK24]. change-point [JZS23]. changes [BDH24, BKW21, CHL24, DR20, FHW23, GHKOB24, HLRW20]. changes-in-changes [GHKOB24]. changing [DGR20, Kas22]. characteristics [GG20a]. chi [DTB21]. chi-squared [DTB21]. child [BBRR22]. children [CEC22, GSS22]. China [HLLO21]. Chinese [ZZ21]. choice [ABL22, AL21, AM20, BKS22, BSX21, CGI21, HMM24, KSSR21, KRvdK22, KMMS21, Lia24, LH24, Lu22, Man23, NS24, PS21a, PVWZ22, STXH23, SY20, Wil20, WY21]. choices [HJ23, KKKN21, LTY21]. circular [HHPT24]. City [BF22]. claims [FLS22]. Clarke [BFM23]. class [PPX24, RW20]. Classical [Kli24]. Climate [Ano24c, MP24a, MP24b, BCS20, CKK^{+}20, CGGK24, DR22b, HPP20, ILMM20, JPS24, MJLS20, PLS20, Pre20, PM24]. clinical [CL24b]. Closed [ASLL21, ASLL24, Gup23]. Closed-form [ASLL21, ASLL24, Gup23]. Cluster [MNW23a, AGP20, CSW23, CK24, Hwa21, KS24, OP24]. Cluster-robust [MNW23a, Hwa21]. cluster-sample [CSW23]. clustered [RV21]. clustering [JLSS24, MNW23b, VL20, YGV24]. clusters [BCL24, KS24, MW20, WZ24]. co [BMPQ22, DLLZ24, HT20, SX22]. co-jump [DLLZ24]. co-movements [SX22]. co-shifting [HT20]. Coefficient [LPG20, Bre21, CFX22, DAM21, DOT22, FL24, FHSW23, JPS24, PW22, PW23, TW22, UWY23]. coefficients [GZW20, HHO22, HN21, JLZ20, LL22, LCL20, LST23, LOW23, SS24b]. cointegrated [BLL21, Pre20, ZGHK24]. Cointegrating [LPG20, Cha20, KP23, PW23, WGH20]. cointegration [iSK21, KASY20, LTY20, PK24, SW21b, Tra21, YfL21]. cointegrations [LT20, TLW22]. collected [PS23a]. collective [Hub23, LL22]. college [PVWZ22, ZZ21]. combination [CGRvD24, CM24, CHL24, LWZ24, MM21b]. combined [CL23c, HL23]. Comment [Bog22]. commodity [HR21]. Common [CMH24, BL21, CGV22, JS22, KOEP20, LCL20, MPS23, SWZZ24]. communal [PW22]. communication [GSV22]. Communities [CHK22]. Community [CFZ23]. comoment [BCV20]. Comparing [Cha23, DR22b, QTZ24, FMM^{+}22]. comparison [HM24, KV23, WAZ24]. comparisons [CCT23, KST22]. competing [Kim23]. complementarity [AR22]. complete [ALR22, JP20]. completion [CKL24]. complex [CFR22]. compliance [BG23, DGJOOSB23]. complier [WTH24]. component [Che24, CY22, LW20]. components [WZ22a]. compositional [ZM24]. computational [HY24]. compute [Woo24]. Computing [AL21]. concentration [MB24]. condition [For24a, Gun23, RW20]. Conditional [Roy23, ZT22, ASZZ24, AF24, BO20, BHS24, BFL24, Car23, CNPR22, CFG23, Dal20, FFX20, FZ20, HW24, HPR24, JC20, JLZ21, Kas22, LWZ24, NP22, PS23b, VW23, WPLL21]. conditionally [CPZ23]. conditioning [LSW23]. conditions [BM21a, CKLZ24, Gal22, HSW24, Kit22]. Confidence [LSZ24, ACG20]. confidentiality [Vil23]. confounded [Kéd23]. confounders [CLMZ24]. confounding [GZW20]. conjugate [Bog22, CCM19, CCCM22]. connectedness [BHSvS21]. connections [EHKS23]. consequences [CMRV24, ZZ21]. Consistency [BH20]. Consistent [AV21]. constancy [DAM21]. Constrained [GHLL22]. Constraints [HS24, AMSV22, CT21, DDH22, FS23a]. constructed [QfLY21]. construction [SZ24]. Consumer [KN21, KKKN21, WML21]. consumption [ABBL24, BT24]. contact [LLSS21]. continuity [BC21]. Continuous [CP21, ALZ22, BLSV23, BK24, BKW23, BCFL21, HSS22b, NP22, uHS23]. continuous-time [BLSV23, BK24, BCFL21]. continuum [ACS20]. contracts [AHZ23, LPV23]. contributes [PW23]. contributions [AAI24]. Control [NS21, AL24, DHS24, Fer21, LY22, LSK24]. convergences [BKW23]. Copula [CXW22, CHY21, FHP23, MW23, NKM22, OP24, SZ24]. Copula-based [CXW22]. copulas [BHK21, FH23, LZ20]. coronavirus [GJ23]. corporate [CK20]. correct [Kle21, LMS23]. corrected [BFM23, DS21, HKL22, HV23]. Correcting [GHKOB24]. correction [SL20]. correlated [LY21]. correlation [CYX^{+}23, Che24, CY22, CLS23, DL24, GLP24, GK24, HLT20, HW24, JC20]. correlation-based [CY22, CLS23]. correlations [BO20, Tsa20]. Corrigendum [AACH23, CCCM22, PST22]. cost [BIJS22]. costs [HMM24]. cotrending [iSK21]. Counterfactual [JP20, ALZ22, KRvdK22, PS21a]. country [LMS21]. country-level [LMS21]. coupled [LW20]. covariance [CHLT24, CK24, DTW22, Gal23, GH23, GKM23, SCC22, WD22, WPLL21, YZC21, ZLLT22, BDSV23, HKST24]. covariance-mean [ZLLT22]. covariances [CBN23, CGQ23, LL23]. Covariate [BJR^{+}24, ZD21, BG23, JPTZ23, KY22, ZD24]. covariate-adaptive [BG23, JPTZ23]. Covariate-adjusted [ZD21]. covariates [BGPS23, BH21, DDF^{+}21, Fen24, GG20b, KKSV21, MW23, NRS23, PW22, TLW22, WZZW24]. coverage [AK24, CFG23]. COVID [GLLZ24, GLLZ23, HLM23, HLS21, JZS23, JSY23, KN21, KV23, Kor21, LLSS21, MM21a, CKS21, LL21, LMS21, PST23, Tou21]. COVID-19 [GLLZ24, GLLZ23, HLM23, HLS21, JZS23, JSY23, KN21, KV23, Kor21, LLSS21, MM21a, CKS21, LL21, LMS21, PST23, Tam21, Tou21]. Cox [PPX24, ZZ24b]. crisis [CGRvD24, HNZ22]. criteria [CGGK24, GCT23]. criterion [LYZ20]. Cross [DS24, ARTT23, BPY21, CFX22, GXZ20, GP20, HS21b, HW22, Hos22, HJPS21, JMS21, KLL21, QTZ24, Uml24, WCWL20, ZL23, uHS23]. Cross-section [DS24, HJPS21, KLL21]. cross-sectional [ARTT23, BPY21, CFX22, GXZ20, HS21b, Hos22, Uml24]. cross-sections [QTZ24, uHS23]. cross-validation [HW22, ZL23]. crowd [DJK21]. CRPS [BZ24]. cure [NK22]. curve [BDFM23, JZS23, WGH20]. curved [Gal23]. curves [VL20]. cycle [BT24, IK21]. cycles [PM24]. Cyclical [CHM23].

D [KKKN21]. Daniel [HKW21]. Data [CFLS24, AA22, ACL22, ATTX24, Ari21, BBGS24, BL21, BCL24, BY23, BM21a, BV23, BLL22, BIJS22, CBN23, CHLZ20, CFX22, Cal21, CSV23, CGRvD24, Cha22, Cha20, CPZ23, CHLT24, CT20, CFVW21, CYX^{+}23, CMZ24, CM24, CHW24, DPS24, DM22, FST23, FGP22a, GXZ20, GLPY23, GJ23, GNP24, HLWZ24, HSJ23, HLLO21, HJW24, HJL^{+}24, Ish20, JYGH21, JLSS24, KPR21, KPS24, LMS23, LMSW23, LP23a, LSTW24, LLV20, LCL20, LY21, LLZ22, LSZ24, LS20b, LS23b, Meh23, MM21b, NSYC21, OW21, PPX24, Pou23, RW24, SW21a, STXH23, SST21, SST23, SH23, SKF23, SXZ23, SWZZ24, Vil23, WX22, WPS24, Yu23, YGV24, ZLB22, ZZZ24]. datasets [KSS21]. date [CP21]. Day [NQV21]. DCS [LW20]. DCS-EGARCH [LW20]. Debiased [Sem23]. debt [CK20, HNZ22, RSW22]. December [Ano23b, Ano24d, Ano24e, Ano24f]. Decision [HS24]. decisions [AHX21, HY23, JS24]. decomposition [MM21b]. defactored [NSYC21]. defined [HSS22a]. degree [CCW20]. demand [BDFM23, JLMM21, LST23, PS21a, RSW22, Wan23, WML21]. densities [Dal20, DSZ24]. density [ALP23, APdAV23, BC21, CGRvD24, GNP24, LSTW24, LNP24, LQ21, ZLB22]. dependence [CFX22, FFX20, GXZ20, GP20, HS21b, HJPS21, KS24, ORS24, SZ24]. dependence-aware [SZ24]. dependency [GLL23]. Dependent [LLV20, BBGS24, CT20, KMS21, LMS23, LPY24, MS21a, Wil20, vdBJMN21]. derivatives [CCT23, RW20]. Design [AI22, AIMM24, Ber20, BC21, DM22]. Design-based [AI22]. designs [BK23, HSW24, Tuv20]. Detecting [BM21b, For24a, GB21b]. Detection [KPR21, CYZ23, FLS22, LS20a, LSK24]. determinants [Woo23]. determined [LPV23]. Determining [Fre22, LS20b]. Deviance [LYZ20]. deviated [LT20]. deviation [WAZ24]. Diagnostic [BPY21, ZZZ24]. dichotomous [FLX22]. Diebold [IKP22]. Difference [AI22, CS21, GB21a, CJS23, Hor21, MW20, RSBP23, SZ20, uHS23]. Difference-In-Differences [AI22, CS21, GB21a, MW20, RSBP23, SZ20, uHS23]. Differences [AI22, CS21, GB21a, LS23a, MW20, RSBP23, SZ20, dCD23, uHS23]. differences-in-differences [dCD23]. differencing [GLX23]. different [BO20]. differential [BS23]. differentiated [BIJS22, Wan23]. Diffusion [BHK21, LS20a, PW21, WZ22a]. diffusions [GS21]. digital [LTZ21]. Digitalization [BDFM23]. dimension [CES20, CJS23]. dimensional [ASW23, BHS20, BKW21, BLL21, BHN22, BMS20, CHLZ20, CLWY24, CPZ23, CCQY24, CHLT24, CFZ23, CGQ23, CCL24a, CCL24b, CMZ24, CM24, CGI20, DDF^{+}21, DLZ21, DGL23, DBH23, FFX20, FHP23, FLLM22, FHW23, GZW20, GLT20, GLLZ23, GLLZ24, HLT20, HHLS22, HJG23, HJW24, KSS21, KASY20, KPT23b, KLSW23, LCW23, LC20, LMNZ24, LS23b, MPS23, MJLS20, PK24, SCC22, Tsa20, UW24, WCWL20, Wan22, WZL24, WLFM24, XP23, YZC21, YN21, YHKZ22, ZZZ24, ZZ24b]. dimensions [CBN23, CMV23, FHLZ20, FLS24, MTWZ24]. dioxide [WGH20]. directions [Tau22]. disappear [DRG^{+}23, DRG^{+}24]. disaster [DN23, SG21]. disaster-type [DN23]. disconnected [MGSN24]. discontinuities [BKL^{+}22]. discontinuity [BK23, Ber20, BC21, HSW24, Tuv20]. discontinuously [KY22]. discount [PVWZ22]. discrepancy [ZGZ24]. Discrete [FS23b, AL21, AM20, Ari21, ABB23, BK24, BSX21, FHW23, HKR20, HJ23, KSSR21, KMMS21, Lia24, Loh23, NS21, NP22, NS24, STXH23, Wil20]. discrete-continuous [NP22]. discrete-time [ABB23]. discriminants [OJT20]. Discussion [Pow23]. Disentangling [PG21, RSV20]. disparity [Par20]. displaced [Woo23]. displacement [Cal21]. distance [AD21b, WD22]. Distinguishing [LPV23]. distortion [JPS24]. Distributed [RLZ^{+}24, FJ22, KY22]. Distribution [BS23, AKOW20, ALZ22, CJ23, Cas24, CJM24, CWXZ24, CFVW24, Hub23, KOPV20, PS23b, YZC21]. Distribution-invariant [BS23]. distributional [ACS20, Cal21, KPV23, Pet22, GG20a, TD20, ZM24]. Distributions [BvBGK24, ACL22, ARTT23, CCW20, CKK^{+}20, HM24, HV20, JMY22, Kit21, LCW23, NP22, TD20, ZGZ24]. diverging [CLMZ24, LZGZ21, LLCW22]. Dividend [PST23]. Do [LMSW23, PS21a]. Does [BP20a]. Domain [GP23, Cha20, CYZ23, KPS24]. dominance [AST20, FMM^{+}22, LLW23, LSW23, LT21, Luo20]. dominant [PY20, PY21]. Double [YCK20, JLZ20, LTYZ23, LTY20, LMR24]. double-machine-learning [LMR24]. double-nonlinear [LTY20]. double-robustness [LMR24]. Doubly [SZ20, HKL22, LCZ23]. draws [HLL21]. drift [COR22, LMSW23, LS20a]. drift-diffusion [LS20a]. driven [AF24, BGK21, BvBKL22, BFL23, BCFL21, CMH24, CKLZ24, GGIS22, GK24, HP24, SWP20, Uml24]. drugs [MB21]. DSGE [KK23]. due [JPS24]. Dufour [CMPZW20]. duration [Bot20, Che23, FOP23, HL20a]. during [CL23a, PST23]. dyadic [GLX23, GNP24]. Dynamic [AHX21, BL21, BLL21, DL24, HS24, HLWZ24, HPR24, JC20, JLSS24, KKSV21, OP24, STXH23, AGL21, AGP20, AA22, AM20, Bai24, BY23, BM21a, BH20, BHSvS21, BHLZ24, BK24, BK20b, BSX21, BD24, DN23, GHKP21, HW24, Han21, HJ23, HJW24, JLMM21, JfL20, KSSR21, KKIS21, KPT23a, Kit22, KZ20, KMMS21, LKLP20, LY21, LXX22, NS24, NSYC21, PS23b, PST23, SCC22, SWX23, SA21, Win24, WY21, YfL21]. Dynamics [HHvR^{+}22, RSV20, ADDP24, BO20, EMS23, NRS23, vGW22].

early [CKS21]. earnings [Bot23, DLCP23, PVWZ22]. Earth [PLS20]. Econometric [GAL20, HPP20, KS23, PY20, PLS20, Pre20, Gua21a, PST23, YS21]. Econometrics [AACH23, Ano24c, BHKL24, CCCM22, PST22, AAI24, BM20, HKW21, LTZ21, Mac23, Pou23, RSVZ20, RSBP23, Tam21, MP24a, MP24b, Ano23a, DvdKWZ22, GAL20]. econometrics/Covid [Tam21]. econometrics/Covid-19 [Tam21]. economic [AV21, ADDP24, DS20, FGP22a, HY22, JPS24]. Economics [DvdKWZ22, Yu22, CFLS24, KKKN21]. economy [ILMM20, LTZ21]. economy-climate [ILMM20]. edges [MZ21]. editor [BHKL24, Vil23, Yu22]. Editorial [IKP22, LTZ21, Mav21, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano23c, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano24g, Ano24h, Ano24i, Ano24j, Ano24k, Ano24l, Ano24m, Ano24n, Ano24o, Ano24p, Ano24q]. editors [HPP20, Ano21o, CMPZW20]. education [MB21]. effect [Cal21, CKL24, CT24, HKR20, HK21, Hos22, HY23, YCK20, ZD21]. Effects [IK21, AX23, ABCR22, Bai24, BKP23, BPQ20, BMP23, BV23, CK23a, CKK^{+}20, CLMZ24, CR24, CFVW24, CO21, DN23, DS21, DGJOOSB23, FOP23, Fus24, GZW20, GGV20, GLPY23, GK24, GL20, Han21, HY24, HSJ23, JLMM21, JPTZ23, Joc23, JSY23, Kam24, KPR21, Kas22, KPS24, Kéd23, Kit22, KLSW23, KPZ23, LMSW23, LPV23, LY21, LSZ24, LP22, LTY21, LH24, LS20b, MMY23, MIMRS24, MLS20, NW21, SU23, SST21, SA21, Sun24, VB23a, WTH24, dCD23]. efficiency [AL21, TPZ23, ZLTT22, ZZ21]. Efficient [CHY21, CCT23, GS21, Gup23, Kle21, KPZ23, LCL20, SCC22, Yan20, BHKvD20, GdXP22]. EGARCH [LW20]. eigenvalue [DTW22, HPR24]. elastic [HHO22]. elasticities [LS24]. elicitation [GLWW22]. elicited [KRvdK22]. elliptical [YZC21]. emissions [WGH20]. Empirical [SG21, CT21, Koh23, MNW23a, MM21b, WAZ24]. employee [CSV23]. Employer [DLCP23, CSV23]. employer-employee [CSV23]. employers [BBRSS23]. Endogeneity [CKT24, CPU23, CW20, KN24]. endogenous [AX23, ABCR22, BHN22, Car23, CCM21, GR23, Hos22, Joc23, QfLY21]. endogenously [HR21, HLM23]. energy [iSK21, Pre20]. engines [GLWW22]. England [KHK20]. Enhanced [SZ24]. enhancement [YYX24]. entry [BK20a]. environmental [WGH20]. epicenter [HLS21]. epidemic [KV23, Kor21]. epidemics [HLS21]. Epilogue [McF21]. episodic [DGRT22]. epistemic [vGW22]. equality [ZGZ24]. equation [iSK21, SW21b, ZHW20]. equations [CK20, CHL21, LM24, LMY20]. equilibria [GJM20, LXX22, Yoo24]. equilibrium [KZ20, JLMM21]. equity [ASKM20, ATU21, GSV22, GLL23]. equivalence [Pre20]. Equivalent [LW23b]. ERP [Cas24]. error [AKOW20, CK20, CSZ22, DOT22, Gel23, KSS21, MM21b, NSYC21, PS21b, Pow23, SL20, Woo24]. errors [BCL24, CMM24, CGI20, Ell20, HHLS22, HM23, HHS20, HSS22b, LPY24, RV21]. Essays [GAL20]. establishing [ACG20]. Establishment [Sch23]. estimate [BCPV23]. estimated [OP24]. estimates [PLS20]. Estimating [BKW21, HLS21, HL23, HHS20, KST22, LP20a, Lu22, MM21a, PDC21, Phi20, RW20, SA21, FZ20, ZHW20]. Estimation [AX23, ALZ22, BHS20, BLL21, BvBGK24, BV23, BS21, BLL22, CSZ22, DOT22, FHP23, FLZ24, For24b, GAL20, HLT20, HR21, HJW24, HHK^{+}24, KSS21, KLL21, KLSW23, LLZ22, LTY20, LOW23, MLG21, MW23, Meh23, MS21b, PY21, QfLY21, RC23, SU23, Tou21, UW24, WLLS24, WX22, WTH24, YfL21, Zin24, AGS24, ASLL24, AA22, AKM21, AD21b, BY23, BAFMS20, BvBKL22, BK24, BLL24, BCV20, BDSV23, BLT21, BMS20, BHK21, BSX21, BIJS22, CHLZ20, CMRV24, CPU23, CP24, Cha20, CHLT24, CSW23, CW20, CHY21, Che21, CCT23, Che23, CR24, CKL24, CGI21, CK24, CHW24, Dal23, DGS21, DPS24, DN23, DMP22, DLZ21, DBH23, EFS23, Erg23, FHLL22, Fun24, GHLL22, GKM21, GdXP22, GNP24, GS21, Gup23, HL20a, HD22, HKT20, HHO22, HM23, HL20b, HLL21, HSJ23, Hos22, Ish20, JfL20, JYGH21]. estimation [JPTZ23, JMS21, JKL24, KN23, KKSV21, KP23, Kor21, LS20a, LP23b, LSTW24, LCZ23, LLV20, LCL20, LWY23, LTYZ23, LNP24, LSZZ20, Loh23, LQ21, LST23, MT23b, MNP20, NRS23, NP22, NS24, NSYC21, PW21, PF23, PK24, RLZ^{+}24, RW24, SKF23, SCC22, Søl20, Tau22, TZ23, VB23a, WGH20, WCWL20, WPLL21, WZLL22, Wan22, Yan20, YS21, Yoo24, YHKZ22, YGV24, ZLB22]. estimator [AF24, DLP21, HKL22, JKW21]. estimators [AD21a, Boo23, CJ23, Can23, CJM24, CCT23, DHIV20, FHLZ20, Fer21, FS23b, GZW20, HSS22a, KSSR21, KPZ23, LMR24, SZ20, dCD23]. ETF [BDSV23]. Euler [CHL21]. euro [HNZ22]. Eurozone [DSZ24]. Evaluating [CKK^{+}20, KKKN21, ORS24]. evaluation [BKS22, BHKvD20, CL23a, HY24, SH23]. event [PPX24, SA21]. events [LKLP20]. Evidence [BF22, CS22, GLWW22, ZZ21, CMM24, LMSW23]. evolutionary [Cas23]. exclusion [DHS24, Kiv20]. Existence [GG20a, FZ22]. Exogeneity [TD20]. exogenous [GK24, MW23]. expectation [GJM20]. Expectations [DvdKWZ22, AHX21, AMMQ22, AL21, CS22, CEC22, DDH22, GMM22, HHvR^{+}22, PVWZ22]. Expected [ILMM20, WTH24, Zin24]. ExpectHill [DGS21]. expectile [DPS24, MTWZ24]. experiment [BKS22]. experiments [BJR^{+}24, DGJOOSB23, GLWW22, HBB24, VB23a]. explainable [FLS22]. explosions [BKN22]. explosive [LPY24]. exponential [ACM22]. exponentially [AD21b, Zhu23]. extended [HMM24, LP23b, RL23]. Extensions [DGRT24]. extent [DRG^{+}23, DRG^{+}24]. external [OSW21]. externalities [WY21]. Extreme [DPS24, ALTZ24, DGS21, HK21, HLPZ24, Tsa20]. eye [HS20].

Factor [AM22, BLL21, CBN23, FKW20, FH22, Fre22, ASKX20, AGP20, BHS20, BN23a, BDH24, BKW21, BH20, BHSvS21, BHLZ24, BK20b, CMV23, CFVW21, CHL24, CLY24, CK23b, DLZ21, DBH23, Erg23, FKL21, FJS22, FLLM22, FHW23, GP20, GCT23, HKTY23, HM23, JMS21, KSS21, KMS24, Lew22, LKLP20, LC20, MLG21, MT23a, MT23b, MW23, NW21, OP24, SCC22, UW24, WLLS24, Wan22, XP23, YHKZ22, YYX24]. Factor-adjusted [FKW20, FJS22]. factor-augmented [CHL24, Wan22]. Factor-based [CBN23]. factors [AM22, ABB23, BLL21, BCV20, CLS23, FKL21, Fre22, Hub23, LP20a, MPS23, SXZ23, AG21]. failure [For24a]. falsification [LMR24]. Fast [LMSND22, DT20, LMS23]. fat [ADDP24, ABB23, KM20]. FDI [HLLO21]. FDR [LSK24]. features [CGV22, WX22]. February [Ano20s, Ano22o, Ano23u, Ano24r]. Feedback [Cha22]. few [Fer21, MW20, WZ24]. fields [JM21]. filter [LLSS21]. filtered [CHY21, CXW22]. filtering [CKLZ24, Dal20, GS21, Lan24]. filters [BN23b]. Finance [Yu22, CFLS24, HM24]. financial [ACL22, ATTX24, BHSvS21, CGRvD24, CCL21, LP22, LTZ21, RSVZ20, SKF23, SCC22]. Financially [CL24b]. Financing [HS24]. Finding [LLSS21]. fingerprint [CKK^{+}20]. Finite [HV23, Sen24, LNP24, SS24b, TD20]. Finite-sample [HV23, TD20]. Firm [EMS23, HS24, APdAV23, DKSS23, LMSW23]. Firms [CSV23, EHKS23, Gua21a]. first [AGS24, EFS23, GG20b, GG22, GZ23, JZ22, KSW24, Zin24]. first-order [KSW24]. first-price [EFS23, GG20b, GG22, GZ23, JZ22, Zin24]. first-stage [AGS24]. Fisher [OJT20, YYX24, ZD21]. fit [KKSV21, LZ20, WD22]. fixed [BMP23, CK23a, Cas24, DS21, GGV20, GLPY23, HSJ23, KPS24, Kit22, LSZ24, MLS20, dCD23]. fixed- [Cas24]. Flexible [DHIV20, CGRvD24]. flow [HKNW23]. flows [PST23]. focus [Man23]. fold [ZL23]. forecast [BP20a, CM24, CHL24, HL23, ORS24]. forecasting [ARRS23, BLSV23, FKL21, HW22, LKLP20, PS21b, QTZ24, WXY23, ZT22]. forecasts [ALP23, CL23c, LMS21]. foreword [BHKL24]. form [ASLL21, ASLL24, CK24, Gup23]. formation [CU23, CEC22, Gao20, GLX23, Gua21a]. Forward [SH23, ZZ24a]. Forward-selected [SH23]. forward-validation [ZZ24a]. Fourier [FHW23]. fraction [HLS21]. fractional [PM24, WXY23]. fractionally [MNP20]. framework [ABB23, CCQY24, Koh23]. Francis [IKP22]. fraud [FLS22]. free [BIJS22, DR22b, HNZ22, JS22, LSTW24, VB23b]. frequencies [HV20]. Frequency [Cha20, GP23, JLP20, ASKX20, ASB20, ASS24, BKS22, CHLZ20, CHLT24, Che24, CMZ24, CHW24, GHM20, HKO^{+}23, KPS24, LLV20, LLZ22, MMF20, PS23a, SKF23, SXZ23, WCLC22, ZLB22, ZLTT22]. Frequentist [KZ21, Tau22]. frictions [BGM21, Par20]. Frisch [Cha24]. frontier [CPU23, CP24, CWXZ24, KHK20, TPZ23]. full [DTPP23]. Fully [KP23, WGH20]. function [DHS24, GdXP22, RW20]. function-valued [RW20]. Functional [PW22, SX22, ALP23, CFX22, CCQY24, EFS23, HLWZ24, LLCW22, PW23, TW22]. functional-coefficient [CFX22, TW22]. functionally [BP20b]. functionals [CT21, Yan20, ZHW20]. Functions [BLL21, AKOW20, ALZ22, AL24, FGP23, HHS20, KOPV20, LS24, LY22, NS21]. fund [AAG20, FJ22, HJLP23]. funding [AG21]. fused [MT23a]. future [SW21a]. fuzzy [HSW24].

G [ZZLL22]. G-GARCH [ZZLL22]. game [EF24, KS23]. games [ALR22, BK24, BK20a, CU23, JP20, Koh23, LXX22, Yoo24]. gap [BMW23, DLCP23]. GARCH [ALP23, BPQ20, CHY21, DW20, FZ22, GL20, HHM22, HPR24, SKY^{+}21, WZLL22, ZZLL22]. GARCH-type [GL20]. Gary [GH22]. gases [CGV22]. Gaussian [CPZ23, FHP23, LM24, TD20]. GB2 [HM24]. Gender [LS23a, HMM24]. gender-specific [HMM24]. general [BHSvS21, CLRSØ23, CLMZ24, GLP24, KZ20, MIMRS24, vdBJMN21]. Generalized [BH20, Can23, GM21, BHLZ24, CZ20, GH23, HM24, JLZ21, LSK24, RW24, Wan22, Zhu23]. Generative [AIMM24]. Generic [ACG20]. geoadditive [KHK20]. geometric [BK20a]. George [CT20]. get [KW23]. Gini [Cha24]. Global [CGV22, Ho23, HT20, BCS20, CMH24, CK23b, KOEP20, KK23, GG20a]. GMM [BCPV23, CR20, FGP^{+}22b, HSJ23, Hwa21, HKL22, HV23]. go [HLM23]. good [HLT21, ZD24]. Goodness [WD22, LZ20]. Goodness-of-fit [WD22, LZ20]. government [RSW22]. gradient [YCK20]. Granger [GHM20]. granular [BM21b]. graphical [FFX20]. greenhouse [CGV22]. gross [LS24]. gross-output [LS24]. Group [MT23a, HJPS21, KPZ23, LCW23, LOW23, Meh23, MSW20, SWX23, WS21, WPS24, YGV24]. groups [GB21b, LSZZ20]. Growth [CFG23, FGP22a, HY22]. Growth-at-Risk [CFG23]. guest [HPP20]. guide [KN23, MNW23a].

HAR [Cas24, SY20]. Hard [MZ21]. Hartz [Woo23]. Hawkes [CLRSØ23]. hazard [EF24, PG21, vdBJMN21]. hazard-rate [EF24]. Head [Kam24]. Health [GMM22, BBRR22, JS24, PG21]. heavy [BBGS24, DGS21, HHLS22, SL20, ZZLL22]. heavy-tailed [BBGS24, HHLS22, SL20, ZZLL22]. hedge [AAG20]. hedging [ABB23]. Hellinger [AD21b]. hemispheric [HT20, KOEP20]. hero [BMPQ22]. Heterogeneity [ABBL24, vGW22, AGS24, APdAV23, AGL21, ABL22, ALL23, Bot20, BMP23, Bot23, CGI21, DKSS23, Esc23, FLZ24, FW23, Fus24, GLP24, HHvR^{+}22, JS24, KST22, LS23b, Luo20, LXX22, Tra21]. Heterogeneous [GXZ20, OW21, ADDP24, BS21, CSZ22, CO21, GLPY23, GH23, HLLO21, KSS21, KS23, LCW23, LCL20, LS24, LTY21, PVWZ22, PPX24, SA21, Sun24]. heteroscedastic [CZ20, CGI20]. heteroskedastic [HHLS22, LMY20]. heteroskedasticity [AX23, AS23, Cas23, JLZ21]. Hierarchical [CD21, FH22, KSS21]. High [ASKX20, ASB20, ASS24, BHN22, CHLZ20, CSV23, CPZ23, Che24, DLZ21, DGL23, FLLM22, JLP20, KASY20, MMF20, MPS23, PK24, WZL24, ASW23, BHS20, BKW21, BMS20, CMV23, CJS23, CCQY24, CHLT24, CFZ23, CGQ23, CCL24a, CCL24b, CMZ24, CM24, CGI20, CHW24, DDF^{+}21, DBH23, FFX20, FHP23, GZW20, GLT20, GLLZ23, GLLZ24, HHLS22, HJG23, HJW24, HV20, KPT23b, KLSW23, LLV20, LLZ22, LC20, LMNZ24, MTWZ24, MJLS20, PS23a, SKF23, SCC22, SXZ23, Tsa20, UW24, WCLC22, Wan22, WLFM24, YZC21, YN21, ZLB22, ZLTT22, ZZ24b]. High-dimensional [BHN22, CHLZ20, CPZ23, FLLM22, KASY20, MPS23, PK24, WZL24, ASW23, BHS20, BMS20, CCQY24, CHLT24, CFZ23, CGQ23, CM24, CGI20, DDF^{+}21, DBH23, FFX20, FHP23, GZW20, GLT20, GLLZ23, GLLZ24, HHLS22, HJW24, KPT23b, KLSW23, LC20, LMNZ24, MJLS20, SCC22, Tsa20, WLFM24, YZC21, ZZ24b]. High-frequency [ASKX20, MMF20, CHLZ20, CHLT24, CHW24, LLV20, SKF23, SXZ23, ZLB22, ZLTT22]. Higher [RR23, AD21a, ACL22, ACM22, Gua21b, LMS23]. higher-moment [ACM22]. Higher-order [RR23, AD21a, LMS23]. highly [WZ22b]. Hill [DHIV20]. Hiring [DKSS23]. Historical [FZ20]. hitting [AS21]. homoskedastic [Kle21]. homoskedasticity [BPY21]. Honor [CMPZW20, GH22, IKP22]. Honoring [NQV21, HKW21]. Honour [GAL20]. horizon [DRT24, KMS24]. household [Hub23, vGW22]. households [LL22]. How-To [Ano23n]. HP [LLSS21]. human [BFLT21, CKK^{+}20, IK21]. Hybrid [LP20b, LMNZ24, WZLL22]. hypotheses [DTB21, Hor21]. Hypothesis [CCL24a, KZA20, CJS23, COR22, Kli24, LLYZ22, CCL24b]. ice [DR22b, DRG^{+}23, DRG^{+}24]. ice-free [DR22b]. ID [AK24]. Identifiability [Fun24]. Identification [AL23, BDKM23, BMP23, BHK21, CU23, CMPZW20, Fus24, Gua21b, HHM22, HKR20, Han21, HJ23, HSS22b, Hub23, Ish20, KPT23a, Kor21, LL22, LS24, Lia24, LSZZ20, LXX22, LX23, PF23, VB23a, Yoo24, AFiK^{+}23, ACF24, AR20, Ari21, BIJS22, CCM21, Che21, CGQ23, DW23, Esc23, FST23, For24a, Gao20, GG20b, GR23, Gun23, HY24, HK21, HSW24, JP20, Kit21, KK23, LP23b, LCZ23, Loh23, MISW20, Meh23, NS21, STXH23, TD20, Tou21, Wil20, Kas22]. Identification-robust [AL23, BDKM23]. identified [AL24, BK20a, DHK20, JP24, LCZ23, LMY20, OSW21, Sem23]. identify [BGPS23]. Identifying [AM20, BKP23, DGJOOSB23, Kam24, Kéd23, Kiv20, SWX23, SXZ23, WS21, GZ23, Kim23]. identity [JM21]. idiosyncratic [CLY24]. Ill [BMS20]. Ill-posed [BMS20]. Illegal [MB21]. Illuminating [HY22]. immigrant [DLCP23]. immigrant-native [DLCP23]. impact [CKS21, CGGK24, Kas22, Phi20]. impacts [JPS24]. imperfect [BG23]. Implementation [KKIS21]. implied [ASLL21, CR20, CCL21, GLL23, LQ21]. importance [LWY23]. impossibility [KS24]. Impossible [BM20, Kiv20]. improvability [PY22]. improve [BP20a]. Improved [LWY23, LT21]. Improving [OJT20]. Impulse [BLL21, GHKP21, IK20, IK22]. Impulse-Response [BLL21]. imputation [CBN23]. inattention [GGIS22, Lia24]. Incentive [GGIS22, LPV23]. Incentive-driven [GGIS22]. Incentives [GLWW22]. incidence [FOP23]. incidental [JS22]. Incoherent [HS24]. income [ABBL24]. Incomplete [HS24, FST23, STXH23, TD20, Yoo24]. Incorporating [DW20, CHW24]. increasing [FHLZ20, FLS24]. independence [ASZZ24, Car23, DW23, Kit21, WLFM24]. index [Gao20, LLCW22, Mar24, NRS23, ZGHK24]. indexed [CS22]. indices [LP22]. Indirect [BY23, FK21, KS20]. individual [KST22, LS20b, MMY23]. Individualized [KW23]. inducing [LMT22]. industry [FJ22]. inefficiency [KHK20]. inequality [Bei24, CT21, FS23a, KZ21]. infection [JZS23, MM21a]. infections [HLS21, LMS21]. Inference [AKOW20, AKM21, ALR22, AL24, ARRW21, BLSV23, BG23, CMPZW20, CCW20, CKL24, DHK20, GLT20, HS21b, HV20, HSS22a, KOEP20, LPG20, LMY20, MMY23, OSW21, RV21, SL20, WZZW24, Xu20, ZLLT22, ASW23, AD21a, ALZ22, AV21, AL23, BHS20, BY23, BM21a, Bei24, BM20, BK20a, Boo23, BLL22, CCL23, CU23, CP21, Cas23, CNPR22, CLRSØ23, CSZ22, CLMZ24, CLY24, DTPP23, DGRT24, FHP23, FST23, FGP23, For24b, FK21, FSU20, FL24, GKR22, GLP24, GZ23, GLLZ23, GLLZ24, HHLS22, HPTZ23, HK21, Hwa21, HV23, IK20, IK22, JLZ20, JLZ21, JMS21, JS22, KSS21, KRW22, KPS24, KN23, KS20, KN24, Kle21, KS23, KS24, KPT23b, KLSW23, LP23a, LR20, LM24, LL20, LT20, LP23c, LMSND22, LS23b, MLG21, MW20, MNW23a, MM21a, MW23, PY21, Pet22, PK24, RLZ^{+}24, RR23]. inference [SU23, SST21, Shi23, SY20, TLW22, TW22, UW24, VW23, VB23b, WGH20, WX22, Wan22, WZ24, XP23, YLC^{+}23, ZHW20, Zhu23, Zin24, vdBJMN21]. inferences [KLP20, ZZLL22]. Inferential [BHLZ24, Tra21]. Infinite [JMY22, ATM20, HL20b]. inflation [AMMQ22, CS22, DSZ24]. inflation-indexed [CS22]. influence [WLFF24]. Influencers [CHK22]. informality [BFLT21]. Information [GCT23, AFiK^{+}23, ASZZ24, ALR22, CHW24, JP20, JP24, Koh23, KLL21, LYZ20, SST23, Yoo24, YGV24]. information-Theoretic [JP24]. initial [BM21a]. Innovate [HS24]. instability [AV22, HW22]. institutions [FMM^{+}22]. Instrument [BIJS22, KN23, Sun24, AK24, Kit21, OSW21, VW23]. Instrument-free [BIJS22]. Instrumental [AGS24, EL21, NSYC21, BMS20, Che21, CLY24, GKM21, GLT20, Hor21, Kle21, WZ24, WTH24]. instruments [CSW23, Fen24, Gof24, Gun23, LWZ24, Loh23, MO23, NS21, Søl20]. insurance [BBRSS23, PG21, SZ24]. integer [BLSV23, CD21]. integer-valued [BLSV23, CD21]. Integrated [SST21, DTW22, FGP22a, LLV20, MNP20, SWP20]. integration [TW22]. interaction [Hos22, HY23]. interactions [AL21, Joc23, LH24, LLCW22]. interactive [Bai24, CK23a, GLPY23, HSJ23, KPS24, LSZ24, MLS20, NW21]. interdependent [GG20b]. interest [DSZ24, GZW20]. interference [HBB24]. interlocks [Gua21a]. Internal [HKNW23]. international [BK20b]. intertemporal [JfL20]. intervals [BH20, LSZ24]. interventions [MIMRS24]. Intraday [ARTT23, DW20, LW20]. Introducing [Ano23n]. Introduction [CSV23, CMPZW20, CFLS24, CT20, DvdKWZ22, GH22, HPP20, MP24a, MP24b, Tam21, Yu22, RSVZ20]. Invariance [TD20, Bai23]. invariant [BS23, Yan20]. invertible [Fun24]. investing [FH22]. investment [CK20]. investments [BBRR22, CEC22, DDH22]. involving [LT20]. Irregular [Esc23, HK21]. Isotonic [BK23]. Issue [CSV23, DvdKWZ22, GAL20, GH22, Mav21, MP24a, MP24b, NQV21, CMPZW20, CT20, IKP22, LTZ21, RSVZ20]. Issues [MNP20]. Itô [SKY^{+}21]. iterated [DT20]. IV [AK24, AL23, CSW23, CKR23, Hos22, KSSR21, KN23, PK24]. IVX [DR22a, DGRT24].

J [AACH23, CCCM22, PST22]. Jackknife [CSW23, CT21]. January [Ano23q, Ano24s, Ano24t]. Japan [CS22]. Japanese [ATU21]. Jean [CMPZW20]. Jean-Marie [CMPZW20]. job [APdAV23, BFLT21, Cal21]. job-title [APdAV23]. JoE [RSVZ20]. joined [KOEP20]. Joint [Boo23, IK22, Din23, GK24]. Journal [Ano23a, DvdKWZ22, GAL20]. July [Ano20o, Ano21p, Ano22m, Ano23r]. jump [BKL^{+}22, DLLZ24, GS21, LS20a, MMF20, PW21, Tod22, WZ22a]. jump-diffusions [GS21]. jumps [ASLL21, Che24, LLZ22]. June [Ano20u, Ano21u, Ano22n, Ano23x]. just [AK24, BCL23]. just-ID [AK24].

Kernel [GNP24, LPG20, BAFMS20, SY20]. Kernel-based [LPG20]. kernels [Kno22, KSW24]. King [GAL20]. kink [SWZZ24]. kinks [LLSS21]. knockoff [LSK24]. known [HLS21]. Kolmogorov [HLM^{+}24]. Kotlarski [Lew22]. Kronecker [GKM23]. kurtosis [LP22]. Kuznets [WGH20].

Labor [BFLT21, CHM23, FOP23, GSS22, HKNW23, IK21, MB21]. LADE [ZZLL22]. LADE-based [ZZLL22]. lag [BCL23]. lagged [Wil20]. Langevin [LMNZ24]. Laplace [CP21, HLL21]. Laplace-based [CP21]. Large [BLL21, Bog22, CCM19, CCCM22, CK23b, CHKP23, XP23, ABL22, BM21b, CBN23, CGRvD24, CES20, Cha23, CGL^{+}22, FZ20, FHW23, GHM20, GB21b, Gup23, HLT20, HPTZ23, HS21b, Ho23, HWZW20, JYGH21, KPR21, KS23, KS24, LMT22, MT23a, Mac23, RLZ^{+}24, RW24, SS24a, SKF23, Tra21, WPLL21, YHKZ22, ZHPW20]. Large-dimensional [BLL21, YHKZ22]. large-scale [ABL22, HPTZ23, HWZW20]. Lasso [ASW23, MT23a, LSG22]. LATE [HSW24]. Latent [AR22, ASLL24, BGPS23, Dal23, GCT23, HJPS21, LCW23, LP20a, LYZ20, LMSND22, LMNZ24, Mar24, Meh23, MSW20, PPX24, SWX23, SXZ23, WLLS24, WS21, WPS24, XP23]. later [CEC22]. law [SS24a]. LDV [HS24]. leads [CMM24]. lean [LMR24]. Learner [VB23b]. learning [BBGS24, BZ24, CCQY24, CT20, FJ22, LMR24, ON21, Phi20, Sem23, WFL22, YCK20]. least [AF24, KP23, RR23]. Lebanese [FMM^{+}22]. Level [HKT20, LW23b, CGQ23, LMS21, MNW23b, YS21]. Levels [vGW22, FMM^{+}22]. leverage [BPQ20, CT24, GK24, HKR20]. life [IK21]. life-cycle [IK21]. Likelihood [Bai24, BM21a, BvBGK24, KZ20, AS21, ASLL24, AA22, ABL22, BDH24, BvBKL22, BK24, CMRV24, CPU23, CT21, CFR22, DS21, DBH23, FS23b, HN21, LWY23, LTYZ23, RC23, SST21, VW23, Wan22, YLC^{+}23]. likelihood-based [CMRV24]. likelihoods [SST23]. Limit [KMS21, BCFL21, PW23]. limiting [Cas24]. Linear [EL21, FW23, KSSR21, AL23, BHN22, Boo23, Can23, Che21, CXZC21, FS23a, FLLM22, FL24, GdXP22, GLLZ23, GLLZ24, HKL22, Kle21, KZ20, LM24, LWZ24, LLCW22, LS23b, MNW23b, NSYC21, SWP20, Sem23, Shi23, UWY23, Win24]. linearization [Bei24]. link [Gao20]. linkages [GLL23]. linked [CSV23]. links [CU23]. Liquidity [NEFG20, GH23]. literature [RSBP23]. loading [Lew22]. loadings [BN23a]. Local [AACH22, AACH23, Bei24, CJM24, BKN22, BP20b, BD24, Fre22, GLL23, HJG23, HJ24, HV20, LP20b, LP23c, OJT20, Xu20]. Locally [LM24, BN23b, FK21]. locally-nonstationary [BN23b]. Location [BvBGK24, HLLO21]. lockdowns [JSY23]. log [HW24, HM24]. log- [HM24]. Logical [GLX23]. logit [AGL21, HHL20, HN21, Kit22, LST23]. lognormal [LNP24]. lognormal-Weibull [LNP24]. London [JSY23]. Long [HKST24, LKLP20, ACM22, BCL23, CR24, DRT24, Erg23, FHLL22, KPS24, KMS24, PS23a]. long-horizon [DRT24, KMS24]. long-run [FHLL22]. Long-term [LKLP20, CR24]. longitudinal [BV23, HLWZ24, LP23a]. look [ZD24]. loses [Kas22]. loss [BKS22]. loudly [GSV22]. low [CKL24, WZL24]. low-rank [CKL24, WZL24].

Machine [BBGS24, ON21, LMR24, Phi20, Sem23, YCK20]. macro [HJL^{+}24]. Macroeconomic [ARRS23, HS21a, GSV22]. major [HMM24, PVWZ22]. making [ASS24]. management [FLS22, SZ24, SCC22]. manifold [Gal23]. many [AM22, AS23, BHN22, Ber20, CSW23, Fer21, HK21, LWZ24, LSW23, LMSND22, MO23, Søl20, WZZW24]. mapping [SCC22]. March [Ano20q, Ano21q, Ano23s, Ano24u, Ano24v]. Marginal [MB24, BKP23, LWY23, TW22]. Marie [CMPZW20]. marker [vdBJMN21]. marker-dependent [vdBJMN21]. market [ASS24, ATU21, BAFMS20, BFLT21, CYX^{+}23, CHM23, FKL21, FOP23, GLL23, HR21, HHvR^{+}22, MB21, NEFG20, NP22, Par20, vGW22]. markets [CCL21, DJK21, HKNW23, JLMM21]. Markov [ABCR22, ASLL24, CD21, GJ23, JLMM21, JMY22, KvD23]. Markov-switching [CD21, KvD23]. Markowitz [AST20, PKHZ24]. Marriage [GSS22]. martingale [CJS23, LZ20]. masks [CKS21]. matched [BJR^{+}24, BV23]. Matching [Fen24, Fer21, Woo23, ZZ21]. matchings [SS24a]. Maternal [CEC22]. mathematical [HSS22a]. matrices [CK24, WPLL21, YZC21]. matrix [ACM22, CHLT24, CXY21, CYX^{+}23, Che24, CK23b, GH23, GKM23, HW24, HKTY23, SKF23, WCWL20, YHKZ22]. matrix-exponential [ACM22]. matrix-valued [CXY21, CYX^{+}23]. matrix-variate [GH23]. matters [ABB23]. Max [CXZC21, WLFM24]. Max-linear [CXZC21]. max-sum [WLFM24]. maximizing [Su21]. Maximum [ASLL24, BvBKL22, BvBGK24, CPU23, LTYZ23, Wan22, YLC^{+}23, ABL22, CFR22, DBH23, FS23b, RC23, ZGZ24]. Maxwell [GAL20]. May [Ano21t, Ano23t]. McFadden [HKW21]. MCMC [GHLL22]. mean [AF24, CHKP23, DGR20, DMP22, FS21, HHLS22, HT20, ZGZ24, ZLLT22, HKST24]. mean-variance [FS21]. measure [AMMQ22, FFX20]. Measurement [CK20, AKOW20, ALTZ24, BHSvS21, CMM24, DOT22, HSS22b]. measurements [WML21]. measures [GH23, Tod22]. Measuring [DHPW24, SSW22]. mechanism [CCL21]. mediation [GLLZ23, GLLZ24]. mediators [GLLZ23, GLLZ24]. Medicare [KKKN21]. medicine [Man23]. medium [ZZ21]. medium-run [ZZ21]. membership [APdAV23, JKL24]. memberships [LOW23]. memory [ACM22, BCL23, Erg23, KPS24]. Mental [JS24]. method [BKL^{+}22, HL20a, YYX24]. Methodology [KKIS21]. Methods [Yu22, CKR23, DGRT24, GG22, KMMS21, LMNZ24]. metric [ZGZ24]. Metropolis [Gal23]. Mexico [BF22]. microstructural [AACH22, AACH23]. microstructure [Che24, LLV20]. MIDAS [KKSV21, MS21b]. migration [KRvdK22]. mild [HLRW20]. mimicking [BDKM23]. Minimax [DM22, Tuv20, Bai23]. Minimax-regret [DM22]. minimum [CHLZ20, DLZ21, FWYZ24]. Mining [WLLS24]. mis [MNP20]. mis-specified [MNP20]. misaligned [Pou23]. misclassification [LH24]. misclassified [CU23]. mispricing [AACH22, AACH23]. misreported [TZ24]. Missing [BGK21, BH21, CBN23, CPZ23, DM22, JMS21, XP23]. misspecification [Pet22]. misspecified [Cas23, GM21, LYZ20]. mitigation [MB24]. Mixed [JKL24, AS21, Cha20, FLS24, GHM20, HHL20, HKO^{+}23].

Mixture [BvBGK24, MS21a]. mixtures [DSZ24, FS23b, HJ23, LNP24]. modal [UWY23]. mode [HWZW20, WLFF24]. Model [CL23c, LZGZ21, Su21, VB23b, ZL23, ALL23, AL23, BM21a, BKN22, Boo23, Bot20, BK20b, BFM23, BCS20, BD24, Car23, iSK21, CSW23, CW20, CLS23, CGGK24, DTPP23, DR22b, Din23, FKW20, FHP23, FLX22, GZW20, GG20b, GH23, Gua21a, HW24, HKTY23, HKST24, HHO22, HMM24, HLM23, HN21, HWZW20, Hub23, ILMM20, JLZ21, KLL21, Kor21, LP23b, LL22, Lia24, LW20, LST23, Mar24, PPX24, PY22, PST23, QfLY21, RLZ^{+}24, RSV20, SHL^{+}21, SHWZ23, SWZZ24, Uml24, UW24, Wan23, YLC^{+}23, ZZ24a, ZZ24b, ZHPW20, Zhu23]. Model-free [VB23b]. modeled [BCL23]. Modeling [GH23, HI20, WXY23, ATTX24, BP20a, CL24a, HLWZ24, PW22, WZL24, XP23]. Modelling [HHPT24, KHK20, PM24, FGP22a, Pre20, SW21b]. Models [BLL21, BvBGK24, CSV23, Fun24, HS24, AFiK^{+}23, AGL21, AD21a, ALZ22, ASKX20, ASLL21, ASLL24, AF24, AR22, AGP20, AA22, AM22, ABL22, AL21, AL24, AM20, ARRS23, Ari21, BKL^{+}22, BHS20, BL21, BN23a, Bai24, BDH24, BCL24, BPY21, BKW21, BGPS23, BY23, BH20, BHSvS21, BHLZ24, Bei24, BHN22, BGK21, BvBKL22, BFL23, BPQ20, BMP23, BMS20, Bre21, BCFL21, BSX21, BLL22, BIJS22, CFX22, CK23a, Can23, CMV23, CP21, Cas23, CD21, CNPR22, CMRV24, CPU23, CP24, Cha22, CPZ23, CZ20, CXY21, CHY21, CFVW21, Che21, CCT23, Che23, CKT24, CLY24, CKR23, CGI20, CK23b, CFG24, CXZC21, CFR22, DL24, Dal23, DAM21, DW20, DS21, DLZ21, DLP21, DOT22, DGL23, DHK20, DBH23, DR20, EFS23, EF24, Erg23, Esc23, FFX20, FKL21, FST23, FLLM22]. models [FLZ24, Fen24, FVvVV24, For24a, FK21, Fre22, FL24, FHSW23, FHW23, FHLL22, GXZ20, Gao20, GLPY23, GLX23, GHM20, GP20, GHKP21, GK24, GM21, GJM20, GKX21, GR23, GP23, Gun23, GCT23, GLLZ23, GLLZ24, HHM22, HHL20, HL20a, Han21, HP24, HMM24, HPP20, Ho23, HNZ22, HSJ23, Hor21, Hos22, HY23, HJW24, HSS22b, HJPS21, IK22, Ish20, JfL20, JLZ20, JYGH21, JMS21, JP24, KSSR21, KPR21, KRW22, KPS24, KKIS21, KPT23a, KN24, Kim23, Kit22, KK23, KvD23, KZ20, KMMS21, Lan24, LR20, LM24, LYZ20, LCL20, LY21, LTYZ23, LSZ24, LZGZ21, LSZZ20, LC20, LLCW22, LY22, LMSND22, LMNZ24, LMT22, LS20b, Lu22, LS23b, LOW23, LX23, LMY20, MLG21, MT23a, MMY23, MT23b, MNW23b, MNP20, MW23, MS21a, MLS20, MJLS20, NKM22, NK22, NS24, NSYC21, OP24, OW21, PDC21, PW21, PS21a, PF23]. models [PS21b, PS23b, PY21, Pet22, Pre20, RC23, RV21, RW20, Roy23, STXH23, SST21, Sem23, SL20, SKY^{+}21, TD20, TW22, Tuv20, WD22, WLLS24, WZLL22, Wan22, WZ22a, WPS24, Wil20, Win24, YS21, YHKZ22, YYX24, ZZLL22, ZGHK24, ZT22, ZM24, vdBJMN21]. moderately [LT20]. modified [KP23, WGH20]. moment [ACM22, Bei24, CKLZ24, DGL23, FST23, For24a, Gal22, Kit22, RW20]. Moments [KvD23, ACS20, ACL22, FZ22, Gua21b]. Monetary [CS22]. monitoring [HLRW20]. monotone [LY22]. monotonic [HSS22b]. monotonicity [BGPS23, CKT24, Gof24]. Monte [AIMM24, FHLL22, KS20]. monthly [HKST24]. moral [PG21]. Most [HJG23]. movements [SX22]. moving [LMS23, Zhu23]. moving-average [LMS23]. Multi [LCW23, CGQ23, Fus24, KSS21, SG21, SWZZ24, YS21, YYX24]. Multi-dimensional [LCW23, KSS21]. multi-factor [KSS21, YYX24]. multi-kink [SWZZ24]. multi-level [CGQ23, YS21]. multi-period [SG21]. multi-valued [Fus24]. multicointegrated [KP23]. multicointegration [BCS20]. multifactor [CSZ22, NSYC21]. multilevel [CLS23]. multinomial [HN21, Lu22]. Multiple [BH23, ABB23, BKW21, CS21, CL24a, CK20, Gof24, LXX22, MT23a, MT23b, PDC21, Yoo24]. multiplicative [HLT20, PS21b]. Multiscale [VL20]. Multivariate [BPQ20, ZHPW20, ARRS23, CD21, CHY21, DTPP23, DW20, DHIV20, GPWY24, GS21, Gun23, HHM22, HLWZ24, JLSS24, KOEP20, NKM22, YfL21]. mutual [ASZZ24, FJ22, HJLP23, WLFM24].

naive [OJT20]. national [CK23b]. native [DLCP23]. near [NW21, SWP20]. Nearest [BCV20]. Nested [BK24]. net [HHO22, HNZ22, IK21]. Network [CHK22, CFVW24, JSY23, BV23, BK20b, CU23, CFVW21, CFZ23, CCT23, FLS22, GLX23, Gua21a, HM23, HHK^{+}24, HWZW20, KW23, KMS21, WLFF24]. networks [CCW20, CLMZ24, DLLZ24, FFL24, HHK^{+}24, HWZW20, JKL24, PY20, RLZ^{+}24, ZHPW20, AIMM24, TPZ23]. Neural [TPZ23, CCT23, CLMZ24, FFL24, WML21]. neutral [LNP24]. neutrality [BDKM23]. Newey [AAI24, KSW24]. News [GLL23, SSW22, ZD24]. News-implied [GLL23]. nicer [BMS24]. Niño [LKLP20]. No [ZD24, RW24]. nodes [CCW20]. nodewise [CMV23]. noise [AACH22, AACH23, CHLT24, Che24, LLV20, Tuv20, ZZLL22]. noisy [CMZ24, LLZ22, WML21, ZLB22]. nominal [Hor21]. Non [BvBGK24, HHK^{+}24, JLZ20, Bog22, BCGR21, CCM19, CCCM22, DDH22, DGJOOSB23, Fun24, GL20, HKT20, HLRW20, HKO^{+}23, LM24, TD20]. non-academic [DDH22]. non-compliance [DGJOOSB23]. non-conjugate [Bog22, CCM19, CCCM22]. non-Gaussian [LM24, TD20]. non-invertible [Fun24]. non-parametric [HKO^{+}23]. Non-representative [HHK^{+}24]. Non-standard [JLZ20]. non-stationarity [HLRW20]. Non-Stationary [BvBGK24, BCGR21, GL20, HKT20]. Noncausal [DS20]. nonclassical [HSS22b]. nondifferentiable [KOPV20]. Nonlinear [CFVW21, RSVZ20, ABBL24, BN23b, BLT21, CYZ23, DS21, GHKP21, KN24, LTY20, SST21, Tau22, WS21]. Nonlinearities [BO20]. Nonlinearity [KKSV21]. Nonparametric [AAG20, Bot20, BH21, CP24, Cha24, CHW24, Dal20, Gal22, Gao20, GG20b, HHO22, HL20b, KV23, LP23b, Loh23, PW21, Tod22, TLW22, WPLL21, Wil20, uHS23, AL23, BLL24, BKW23, CHY21, CHL21, DLP21, Esc23, FJ22, FSU20, Han21, Hos22, HSS22b, KLL21, LW23a, LSTW24, LL20, LST23, VL20, Wan23, ZZ24b]. nonresponse [FMM^{+}22]. Nonseparable [FVvVV24, GR23, Ish20, MMY23]. nonstationarity [Cas24, CXW22]. Nonstationary [HJPS21, BN23b, Cas23, DLP21, PDC21]. nontransferable [GLX23]. Normal [BvBGK24]. normalised [Fun24]. normality [GGV20]. normalization [HLM^{+}24]. normals [FS23b]. notching [BMS24]. Note [Ano21o]. November [Ano20n, Ano23o]. Nowcasting [BMW23, CGL^{+}22, HKO^{+}23, ADDP24]. NPIV [CCT23]. nuisance [CO21, Xu20, ZHW20]. null [BLT21, JLZ20]. number [CK23a, CLMZ24, Fre22, LZGZ21, LSZZ20, LLCW22]. numbers [SS24a].

objective [HD22]. observables [KY22]. Observation [CKLZ24, AF24, BGK21, Che24, GK24, MS21a]. observation-dependent [MS21a]. Observation-driven [CKLZ24, AF24, BGK21, GK24]. observational [DDF^{+}21]. observations [BGK21, CYX^{+}23, Fer21, HI20, PS23a, XP23]. Observed [MZ21, GJ23]. occasionally [AMSV22]. occasionally-binding [AMSV22]. Occupation [ZLB22]. October [Ano20r, Ano21r, Ano22q, Ano23p]. off [AM20]. offer [Kam24]. older [BDFM23]. OLS [WGH20]. on-the-job [BFLT21]. One [AK24, HKTY23, BCL23, CEC22]. One-way [HKTY23]. online [GLWW22]. Optimal [BKS22, BLL24, CHLT24, EL21, ZZ24a, BP20b, Che21, CL23c, JSY23, KSW24, SY20]. optimality [Bai23]. optimization [For24b, JfL20]. Option [BAFMS20, CL24b, LQ21]. option-implied [LQ21]. options [CT24, Tod22, TZ23]. order [AD21a, BC21, DS21, DHK20, HL20b, KSW24, LMS23, LX23, MO23, RR23]. ordered [ALR22]. ordering [ARRS23]. origins [DKSS23]. Ornstein [WXY23]. Out-of-sample [CFG23]. outcome [Kit21]. outcomes [GSS22, Koh23, MB21]. outliers [JPS24]. output [BMW23, LS24]. Over-identified [LCZ23]. Overbidding [GZ23]. overidentified [Gal22]. Overlap [DDF^{+}21]. overlapping [CFG24, HLL21]. overnight [DW20, LW20]. overspecified [LSZZ20]. Overview [ACL22, HKW21, KKIS21].

Pages [Ano20t, Ano20s, Ano20u, Ano20o, Ano20q, Ano20n, Ano20r, Ano20p, Ano21s, Ano21u, Ano21p, Ano21q, Ano21t, Ano21r, Ano22r, Ano22p, Ano22o, Ano22n, Ano22m, Ano22q, Ano23w, Ano23v, Ano23u, Ano23q, Ano23x, Ano23r, Ano23s, Ano23t]. pair [SZ24]. pairs [BJR^{+}24]. Pairwise [OJT20, KST22, YLC^{+}23]. pairwise-rank-likelihood-based [YLC^{+}23]. pandemic [CL23a, CKS21, GLLZ23, HKO^{+}23, KN21, LL21, PST23, Tam21, GLLZ24]. Panel [HLPZ24, LMS21, MLS20, MSW20, WPS24, ABCR22, AA22, ABL22, ALL23, Ari21, BBGS24, BKL^{+}22, BL21, Bai24, BCL24, BPY21, BY23, BM21a, CBN23, CFX22, Cal21, Cha22, CFVW21, CFVW24, DS21, DM22, FGP22a, FW23, GXZ20, GLPY23, HBB24, HSJ23, HLLO21, HJW24, HJPS21, Ish20, JYGH21, JLSS24, KPR21, KSS21, KPS24, KHK20, LCL20, LY21, LSZ24, LSZZ20, LS20b, LS23b, LOW23, Meh23, NSYC21, OW21, PW22, SW21a, SST21, SST23, SH23, SWZZ24, Win24, YfL21, YGV24]. panels [AM20, BS21, BM21b, CSZ22, HS21b, JS22, KS20, KKSV21, NW21, SWX23, Tra21, WS21]. panic [KN21]. papers [Ano23n]. parameter [AV22, BKN22, CNPR22, CO21, DR20, HD22, KP23, LSTW24, MH20, ON21, RL23]. parameter-free [LSTW24]. Parameters [BHKL24, BKL^{+}22, Cal21, CKLZ24, FS23a, HK21, HLM23, HN21, JS22, LZGZ21, RW20, Xu20, YN21]. Parametric [Erg23, LNP24, AACH22, AACH23, HKO^{+}23, ZGHK24]. parametrisation [Fun24]. Parental [BBRR22]. Parsimony [LMT22]. Part [KKKN21]. Partial [FST23, GR23, BMPQ22, Che21, DL24, JP20, RC23, Tou21]. Partially [BHKvD20, Kim23, AL24, BGPS23, Che21, GJ23, JP24, LMY20, UWY23]. participation [CMM24, TZ24]. pass [AM22, BGS24]. Past [SW21a]. pathways [MB24]. patterns [KHK20]. pay [EMS23]. peak [LL21]. Peer [Joc23, KPZ23, LH24]. PELVE [LW23b]. penalization [GHLL22]. Penalized [SHWZ23, BGS24, CHL21, HN21, LP23a, MS21b]. Penetrating [TX23]. pensions [IK21]. perceptions [BF22]. Perfect [JLMM21]. performance [AAG20, HJLP23, ORS24, QTZ24, Uml24]. period [SG21]. periods [CS21, CK23a, CGRvD24]. permanent [Phi20]. Permutation [CO21, Bai23]. Perron [NQV21]. persistence [ALTZ24, ABBL24, CPRR23]. persistent [AV21, AV22, Ell20, WZ22b]. personalized [Man23]. perspective [AACH22, AACH23, JZS23, WAZ24]. pervasive [KPR21]. pervasiveness [GCT23]. PI [NQV21]. PI-Day [NQV21]. Pierre [NQV21]. planned [FOP23]. plans [KKKN21]. plight [MGSN24]. pockets [ALTZ24]. Point [BP20b, JP20, CLRSØ23, CYZ23, DBH23, JZS23]. points [BHS20, Fen24]. policies [CKS21, DLCP23, WZZW24]. Policy [CL23a, HY24, MIMRS24, SU23]. polynomial [WGH20]. pooled [JKW21]. pooling [JMY22]. Population [HBB24, SS24b]. portfolio [BKS22, CHLZ20, CLWY24, DLZ21, FWYZ24]. portfolios [BDKM23, FZ20]. posed [BMS20]. possibly [Cas23, CFG24, Fun24, GL20, MISW20]. Post [LL23, HHLS22, ZZZ24]. Post-processed [LL23]. post-screening [ZZZ24]. post-selection [HHLS22]. Posterior [KOPV20, LLYZ22, BHKvD20, DMP22, Kli24]. Posterior-based [LLYZ22]. posteriors [LL23, LWY23]. potential [Kit21]. Power [YYX24, HLT21, LWY23, VW23, WZLL22, Roy23]. powerful [HJG23]. practice [JYGH21, KN23, MNW23a, Yu23]. precision [WCWL20]. predict [BGS24, CEC22]. Predictability [PT11, PST22, ATU21, DGRT22, DGRT24, DRT24, FS21, GL20, HLT21, KMS24, TX23, WFL22]. predictable [HJL^{+}24, KMS24]. Predicting [AG21, Uml24]. prediction [BH20, CZ20, DJK21, FLX22, JP20, LZGZ21, Man23, MS21b, SZ24, Su21, WAZ24, ZL23]. predictions [PS21a]. Predictive [ATTX24, CFG24, FLS24, LLCW22, AV21, AV22, CCL23, CGRvD24, CLY24, DR22a, DSZ24, JMY22, KASY20, LSG22, LP23c, YN21, ZGHK24]. predictors [WZ22b]. preference [CGI21]. preferences [DDH22, Hub23, PVWZ22]. preferential [CCW20]. premia [ASKM20, BGS24, FKL21, RSV20, Uml24]. premium [AG21]. presence [ABBL24, BKP23, HKT20, Kéd23, KASY20, NRS23, Pet22, WZ22a, Xu20]. present [SW21a]. prevalence [Tou21]. price [ASB20, Dal20, EFS23, GG20b, GG22, GZ23, HR21, JZ22, KLP20, LQ21, Phi20, RSV20, Zin24]. Prices [AKA23, BT24, GSV22, HSHS20, HS21a, KLL21]. pricing [BAFMS20, BDKM23, CFX22, CL24b, Dal23, FFL24, FLLM22, GM21, GKX21, LP20a, SZ24, SG21, Uml24, Win24, YYX24]. principal [Che24, CY22]. priors [Bog22, CCM19, CCCM22, LMT22]. privacy [BS23, Vil23]. private [PG21]. Probabilistic [Man23, GMM22]. Probabilities [DvdKWZ22, CR20, HSHS20, Hor21, KRvdK22]. Probability [DR22b, LW23b, DSZ24, Kli24, PS21b]. problem [HKT20, KK23, MM21a]. procedure [AGP20, AM22, LSK24]. process [ASB20, GJ23, PM24, WXY23]. processed [LL23]. processes [ATM20, BLSV23, CLRSØ23, DS20, FZ22, GPWY24, LS20a, LT20, MNP20, SWP20]. Product [GKM23]. production [AKA23, HHS20, KHK20, LS24, PY20]. products [BIJS22, Wan23]. Professor [CT20]. Profile [HSJ23]. Profiling [MGSN24, MGW23]. profits [AKA23]. program [CMM24, Kam24, SH23, TZ24]. programming [HSS22a]. Projected [YHKZ22]. projection [FFX20]. projection-based [FFX20]. Projections [DRG^{+}24, DR22b, HJ24, DRG^{+}23]. propensity [HK21]. properties [BN23b, CY22, DMP22, HLT21, HHK^{+}24, KZ21, Shi23]. property [CLWY24]. proportion [NK22]. proxies [AKA23, ACF24, WLLS24, WAZ24]. Proxy [ARRW21, ACF24, GKR22, HHS20]. Proxy-SVARs [ARRW21, ACF24]. Pseudo [BK24, FS23b]. public [FMM^{+}22, RSW22]. pure [LR20]. puzzle [HKR20].

QLR [FS23a]. QMLE [MH20]. quality [YCK20]. Quantile [CW23, EF24, GG22, ALZ22, ALL23, CCL23, CL23b, Che23, CCL24a, CCL24b, CFVW24, CFG23, FLS24, FGP^{+}22b, GGV20, HPTZ23, HLPZ24, JPTZ23, LP23a, MLG21, NK22, SWZZ24, TLW22, WZLL22, Zhu23]. quantiles [CL24a]. Quasi [BFL23, DBH23, MB24]. Quasi-maximum [DBH23]. quasi-representative [MB24].

random [Bre21, HHO22, HN21, JM21, JMS21, KMS21, LY21, LST23, WLFM24]. Randomization [MW20, BG23, ZD21]. randomizations [JPTZ23]. randomize [Bai23]. randomized [VB23a]. range [BLL24, HLM^{+}24]. range-based [BLL24]. Rank [WLFM24, CKL24, FHLZ20, WZL24, YLC^{+}23]. Rank-based [WLFM24]. rate [EF24, LLSS21, LMR24, MM21a, vdBJMN21]. rates [BH20, DSZ24, PVWZ22]. Ratio [CMV23, BDH24, KZ20, VW23, WCLC22]. rational [AL21, GJM20, Lia24]. ratios [MO23]. RCP8.5 [MB24]. reaction [GLLZ23, GLLZ24]. Real [WFL22, DSZ24, HJL^{+}24]. Real-time [WFL22, HJL^{+}24]. Realized [ACM22, BMPQ22, CMZ24, BPQ20, GK24, GH23, HP24, SKY^{+}21, WXY23]. recommendation [KPV23]. record [CP21]. recovery [CYX^{+}23]. recurrent [BLT21]. Recursive [JYGH21]. Reducing [CES20]. reduction [TZ23]. refinements [MO23]. Refining [DW23]. refitted [WCWL20]. Reflections [Vil23]. reforms [CS22, Woo23]. regime [MS21a, UW24]. regimes [BO20]. region [Kit21]. regional [KHK20]. regions [ON21]. Regression [Ber20, JPTZ23, LPG20, Tuv20, BK23, BGS24, Boo23, BLT21, BH21, BC21, CJ23, CCL23, CMV23, CJM24, CSW23, CZ20, CW20, CFZ23, CW23, CL23b, Che23, CCL24a, CCL24b, CMZ24, CFVW24, CGI20, CXZC21, CYZ23, DR22a, DRT24, FJS22, FLS24, FGP^{+}22b, GGV20, GHM20, GG22, GLT20, GdXP22, HHLS22, HPTZ23, HL20b, HLPZ24, HSW24, HSS22b, KSSR21, Kle21, KKS23, KASY20, LP23a, LSG22, LW23a, MNW23b, MTWZ24, MZ21, NK22, PW23, Shi23, SS24b, SWZZ24, TW22, UWY23, VL20, Xu20, ZZ24b, ZGHK24, ZLLT22]. Regression-adjusted [JPTZ23]. regression-based [DRT24]. regressions [ASKX20, AV21, AV22, BBGS24, BKW23, CHL24, DS24, FW23, KM20, LT20, LP23c, MSW20, MS21b, PS23a, WGH20, Wan22, WZ24, YN21]. regressors [Che23, GHKP21, GK24, GR23, GL20, HM21, LP23c, Loh23, NSYC21, ZGHK24]. regret [DM22]. regular [CGRvD24, CT21, Kno22]. regularization [CXZC21]. Regularized [DSZ24, FKW20]. rejection [Hor21]. related [KOEP20, VW23]. Relaxing [Car23, DHS24]. Relevant [DR20, Fre22, SU23]. reliable [GZW20]. repeated [uHS23]. reports [CMM24]. representative [GLWW22, HHK^{+}24, MB24]. Reprint [CCL24b, DRG^{+}24, GLLZ24]. Reproducibility [Vil23]. rerandomization [ZD24]. Residual [CMV23, DR22a, BHS24]. Residual-augmented [DR22a]. Residual-based [CMV23]. resource [BMP23]. Response [BLL21, ALR22, Ari21, Car23, FLX22, GLPY23, GSV22, GHKP21, GR23, IK20, KPT23a]. responses [ABBL24, IK22]. restricted [LT21, Luo20]. restriction [DHS24, DGL23, GZ23]. restrictions [AS23, BGPS23, GHM20, Kiv20, Yan20]. results [ACG20, KS24]. Retire [MTWZ24]. Retirement [GMM22]. return [ATU21, ALTZ24, DGRT24, GL20, HLT21, HL20b, NRS23, TX23, WFL22]. returns [BGS24, BBRR22, DGRT22, DW20, Din23, HJL^{+}24, KMS24, PT11, PST22, SX22]. revealed [HMM24]. revenue [Zin24]. revisited [HKR20]. Revisiting [HLLO21]. Risk [CFG23, ASKM20, ATU21, ARTT23, AG21, BGS24, BT24, BHKvD20, DGS21, DHPW24, FKL21, FHLL22, HS20, HL20b, HLPZ24, ILMM20, JZ22, LS23a, LNP24, NRS23, PVWZ22, RSV20, SCC22, SG21, Uml24, BHS24]. risk-neutral [LNP24]. risk-return [HL20b]. risks [Kim23, SZ24]. risky [JS24]. Robust [ATM20, AA22, AD21b, BLT21, Che21, GKR22, GLP24, GZ23, HHLS22, KPS24, LT20, LP23c, LPY24, MTWZ24, Søl20, AGP20, AL23, BDKM23, BHKvD20, CCL23, Cas23, Che24, DTB21, Ho23, Hwa21, HKL22, LM24, MNW23a, SZ20, SKF23, TW22, VW23, WAZ24, LCZ23]. Robustifying [PKHZ24]. robustness [HHS20, JKW21, LMR24]. Role [HMM24, APdAV23, ASS24, AG21, ADDP24, BBRSS23, BM21a, DDH22, PVWZ22, SST23]. rolling [CJ23]. root [HKT20, LT20, LP23c, NW21]. roots [BP20b, FLS24, LP20b]. rotation [ZT22]. roughness [BCPV23]. rounding [GMM22]. Roy [HMM24, LP23b]. rule [AK24]. rules [FVvVV24]. run [AG21, FHLL22, ZZ21].

safety [IK21]. Sample [LY22, BKP23, Cha20, CSW23, CW23, CGQ23, CKL24, CFG23, DM22, FVvVV24, GL20, HV23, LW23a, RC23, TD20]. sampled [HR21, HHK^{+}24]. samples [Mac23]. Sampling [DMP22, HV20, LWY23]. SAR [QfLY21]. satellite [RW24]. saving [CK20]. savings [IK21]. Scalable [DTPP23]. scale [ABL22, HPTZ23, HWZW20, LMT22, RLZ^{+}24, ZHPW20]. school [PS21a, ZZ21]. science [CFLS24, Yu23]. Score [CR20, HP24, Uml24, BvBKL22, BFL23, BCFL21, FS23a, HW24, SST23]. Score-driven [HP24, Uml24, BvBKL22, BFL23, BCFL21]. scores [HK21]. screening [ZZZ24]. sea [DRG^{+}23, DRG^{+}24]. sealed [JZ22]. sealed-bid [JZ22]. search [BFLT21, GLWW22]. Seasonal [HKST24]. Second [DS21, MO23, DHK20]. Second-order [DS21, MO23, DHK20]. section [DS24, HJPS21, KLL21]. sectional [ARTT23, BPY21, CFX22, GXZ20, GP20, HS21b, Hos22, Uml24]. sections [QTZ24, uHS23]. seemingly [WGH20]. segmented [KOEP20]. SEIRD [Kor21]. select [HN21]. selected [SH23]. selection [BKP23, BFM23, CLWY24, CW23, CGI20, CLS23, CGGK24, FKW20, FVvVV24, HHLS22, HJW24, KY22, LPV23, LY22, Mar24, MS21b, PY22, PG21, RC23, Su21, WCLC22, Woo23]. selectively [BH21]. Self [BF22, HLM^{+}24]. self-normalization [HLM^{+}24]. Self-perceptions [BF22]. seller [Zin24]. Semi [LST23, ZGHK24, CHY21, Wan23]. Semi-nonparametric [LST23, CHY21, Wan23]. Semi-parametric [ZGHK24]. semicovariance [BPQ20]. Semiparametric [Ari21, BSX21, CL24a, CW20, CR24, Dal23, FLX22, NS24, UWY23, WZ22b, AL21, BKW23, BLL22, DGL23, FHP23, HL20a, JLZ21, LL22, LLCW22, MLG21, YLC^{+}23, vdBJMN21]. Semiparametrically [GdXP22]. sensitivity [PLS20]. sentiment [SSW22]. separability [CP24]. separable [CKT24, ZGZ24]. September [Ano20p, Ano23y]. sequence [HJG23]. sequences [HKTY23]. Sequential [HLRW20, FHLL22, Yoo24]. serial [Tsa20]. series [ASW23, AF24, ACL22, BGK21, BLT21, BM21b, CD21, CKK^{+}20, CCQY24, CXY21, CXW22, CFZ23, CGI20, CYZ23, DGR20, DS20, DN23, DLP21, HR21, HI20, HHPT24, HKST24, HV23, JZS23, KLL21, LL20, LZGZ21, LC20, NKM22, PM24, RV21, SX22, Tsa20, WD22, WZL24, ZM24]. serology [Tou21]. set [AFiK^{+}23, BK20a, CGI21, DW23, GHM20, Sem23]. set-identification [DW23]. set-identified [BK20a, Sem23]. sets [ACG20, CGRvD24, Lu22]. settings [AI22, KLSW23]. several [BLL22, ZGZ24, dCD23]. shares [BMP23]. Sharp [Mar24, STXH23]. Sharpe [CMV23, WCLC22]. shift [HKT20]. Shifting [HKST24, HT20]. shifting-mean [HT20]. shocks [ABBL24, BL21, Bot23, CMH24, DN23, HLLO21, JS22, KvD23, LCL20]. short [AG21, AM20, DPS24]. short-run [AG21]. short-tailed [DPS24]. shortfall [WTH24]. should [HD22, KW23, MMF20, Woo24]. Shrinkage [HM21, HM23, MT23b]. Sieve [FL24, Hos22, LQ21, DLP21, KMMS21, Wan23]. sign [Kno22]. sign-regular [Kno22]. signal [CGQ23]. Significance [FFL24]. SII [HS24]. similarity [RL23]. similarity-based [RL23]. Simple [AD21a, GZW20, HL20a, HLT21, Hwa21, AHZ23, iSK21, Din23, DTB21]. Simulation [CCT23, HLL21, SG21, FZ20]. simulation-based [SG21]. simulations [AIMM24, LKLP20]. Simultaneous [KRW22, LM24, LMY20]. Simultaneously [HS24]. single [LLCW22, SW21b, ZGHK24]. single-equation [SW21b]. single-index [LLCW22, ZGHK24]. singular [Gal23]. sinusoidal [PM24]. SIR [KW23]. size [ACG20, HLT21, Kle21, WY21]. skewed [HM24]. skill [CEC22, FJ22]. slope [BS21]. Small [GL20, CK23a]. Small-sample [GL20]. Smirnov [HLM^{+}24]. smooth [CHL24]. Smoothed [HPTZ23, RV21]. smoothed-clustered [RV21]. smoothing [HS21b, KMMS21, Lan24, PW22]. Social [EHKS23, KKS23, BBRSS23, JKL24, Joc23, LTY21, LH24, ZHPW20, CHK22]. solution [KK23]. Solving [CHL21, KMMS21]. Some [KS24, HI20]. SONIC [CHK22]. Sorting [Woo23, CHM23, EHKS23, LS23a]. sources [BLL22]. space [CNPR22, CES20, CPZ23, GH23, KZ20, Lan24, LMT22, RL23]. spaces [ZGZ24]. span [PS23a]. Spanning [AST20, BDKM23]. Sparse [CL23b, LLSS21, MGW23, RW24, FJS22, FLLM22, KPT23b, LL23, SW21b]. sparsity [Can23, WPLL21]. Spatial [JfL20, LY21, Pou23, RL23, ALL23, BL21, BPY21, CKK^{+}20, CSZ22, Gup23, HJW24, LR20, MH20, PY21, QfLY21, RC23, RLZ^{+}24, RR23, SWX23, YfL21, ZHPW20]. spatio [MGW23, RW24]. spatio-temporal [MGW23, RW24]. spatiotemporal [FLZ24]. speak [GSV22]. Special [CSV23, CMPZW20, GAL20, Mav21, IKP22, LTZ21, RSVZ20]. specific [HMM24]. Specification [FHSW23, HHL20, JfL20, CWXZ24, PS23b, Yan20]. specifications [Cha23]. specified [MNP20]. spectra [Cas23]. Spectral [YGV24]. speculation [HR21]. speed [ASS24]. spending [RSW22]. Spherical [ZM24]. spillover [BV23, VB23a]. spillovers [ACM22, DGJOOSB23, HM23, KvD23]. splines [CHL21]. splitting [CLWY24, CKL24, LW23a]. sporadic [TX23]. spot [DTW22, TZ23]. spread [JSY23]. spreads [AG21]. Spurious [TW22, LT20]. squared [DTB21]. squares [AF24, KP23, RR23]. Stability [YfL21]. Stable [Koh23, LTYZ23, SS24a]. stage [AGS24, AF24, CHL21, KS20]. staggered [AI22]. staleness [BPR24, KLP20]. Standard [BCL24, Gel23, JLZ20, OJT20, Pow23, RV21, Woo24]. star [ZD24]. Start [Kam24]. State [CYZ23, CES20, CPZ23, Dal20, GH23, KZ20, Lan24, LMT22, LQ21, ORS24]. State-domain [CYZ23]. state-price [Dal20]. state-space [GH23, Lan24, LMT22]. Stationarity [JLP20, HLRW20]. Stationary [BvBGK24, GJM20, NKM22, BCGR21, FK21, GL20, HKT20, TLW22].

Statistical [iSK21, CT20, GLLZ23, KLP20, MJLS20, DR22b, DLZ21, Zhu23, GLLZ24]. statistics [AGL21, BC21, KZA20, LLYZ22, LX23]. stayers [Ish20]. steel [HR21]. Stein [Boo23]. Stein-type [Boo23]. STEM [HMM24]. step [Che23, HV23]. Stochastic [BN23b, AD21a, ASLL21, ARRS23, AST20, ACM22, Bog22, BCPV23, BCGR21, Bot20, CCM19, CCCM22, CPU23, CP24, Cha23, CWXZ24, CHKP23, DTPP23, FMM^{+}22, For24b, ILMM20, KHK20, KZ20, LLW23, LP20b, LSW23, LP23c, LT21, Luo20]. Stock [DLLZ24, NRS23, BGS24, CYX^{+}23, DGRT22, GL20, GLLZ23, GLLZ24, HLT21, HHvR^{+}22, KMS24, NP22, PT11, PST22, vGW22]. strategic [AL21, BH23, HY23]. strategic-interactions [AL21]. strategy [ACF24]. strength [AR20, KN23]. Strong [AM22]. strongly [LPY24]. Structural [GP23, HKW21, KPT23b, OSW21, AFiK^{+}23, AGL21, AHX21, AHZ23, AV22, BDH24, BKW21, BP20a, CP21, CHL24, CFR22, DR20, Esc23, FS23b, FHW23, GHKP21, Gua21b, HHM22, HW22, HLM^{+}24, HHK^{+}24, KSSR21, KKIS21, LSK24, LOW23, MT23a, NS21, OW21, PS21a, PT11, PST22, Shi23, Tau22]. structure [ASKM20, AGP20, CSZ22, DTW22, HLT20, KSS21, LOW23, NSYC21, SWZZ24, YGV24, GKM23]. structured [Can23, HM23]. structures [CYX^{+}23, HJPS21, LCW23, Meh23, MSW20, SWX23, WS21, WPS24]. studies [CGGK24, DDF^{+}21, SA21]. Study [GMM22, GCT23, KKKN21, ZZZ24]. studying [GLLZ23, GLLZ24]. subject [FS23a, Gal22]. Subjective [DvdKWZ22, AHX21, CEC22, GLWW22, HHvR^{+}22]. subset [Kle21]. substitutes [BH23]. subvector [Bei24]. Sufficient [AGL21, WLLS24]. sum [WLFM24]. summary [LSTW24]. Supplementing [Fen24]. supply [GSS22, IK21]. support [CYX^{+}23]. surfaces [ASLL21]. Survey [ZHW20, CGI21, FMM^{+}22, GLWW22, MM21b]. Surveying [ABB^{+}22]. surveys [CMM24]. survival [NK22]. suspensions [PST23]. SVARMA [Fun24]. SVARs [ACF24, ARRW21, AMSV22, GKR22]. switching [ABCR22, CD21, KvD23, MS21a, UW24]. synchronization [ABCR22]. synthesis [RSBP23]. Synthetic [VB23b, iSK21]. system [KOEP20]. Systematic [BPR24, ARTT23, BS21]. systems [AV21, KP23, Pre20].

tabulated [LSTW24]. Tail [ATU21, CMRV24, GMM22, NRS23, DHPW24, WX22]. tailed [BBGS24, DPS24, HHLS22, SL20, ZZLL22]. tailored [HD22]. tails [ADDP24, ABB23, BR22, DGS21, KM20]. Taking [KMS24]. targets [KPV23]. taxes [IK21]. technology [CEC22]. temperature [HKST24]. temperatures [CGV22, HT20, KOEP20]. tempered [SWP20]. temporal [HS21b, LP22, MGW23, RW24]. tensor [WZL24]. tenuous [HS21a]. term [ASKM20, CR24, LKLP20]. test [ATM20, BDH24, BFM23, CO21, DTB21, FMM^{+}22, FLLM22, GKM23, HHL20, HJG23, KY22, LWZ24, LT21, LZ20, MMF20, PS23a]. Testing [ASZZ24, ACS20, AS23, AV22, AR20, BC21, CFX22, CJS23, CGQ23, CYX^{+}23, CWXZ24, DHS24, DAM21, DGRT22, DTW22, Ell20, FZ22, FHW23, GAL20, GHM20, HSW24, JLP20, JPS24, JZ22, Kiv20, LLW23, LSW23, MNW23b, MS21a, SY20, Tsa20, WZ22a, Win24, YZC21, ZGZ24, ACF24, BKW21, BP20b, CCL24a, CCL24b, DTB21, FS21, GGR24, HKT20, HLM^{+}24, KZA20, KZ20, LLYZ22, LPY24, WD22, WZ22b, YYX24]. Testing-optimal [SY20]. tests [ACG20, AST20, BPY21, iSK21, Cas24, CR20, CFG23, CFG24, DRT24, FFL24, FS23a, FHSW23, GL20, HLT21, Hor21, KZ21, KMS24, LLZ22, LMR24, MMF20, MISW20, PS23b, TD20, Tou21, VW23, WLFM24, Yan20, ZD21, ZD24]. them [AK24]. Themed [MP24a, MP24b]. theorems [KMS21]. Theoretic [JP24]. theoretical [CLWY24]. Theory [BBGS24, BM20, Cas23, JYGH21, BHLZ24, DGR20, KN23, PW23, SWP20, TD20, Tra21, Tsa20]. thickness [DRG^{+}23, DRG^{+}24]. three [CMM24]. Threshold [LW23a, LC20, GGR24, KKS23, MT23b, MLS20, MSW20]. thresholding [CGQ23]. thresholds [Ber20]. Tiao [CT20]. tilted [AD21b]. Time [ACL22, BHSvS21, BHKL24, Bot23, CM24, CHL24, DN23, FWYZ24, GPWY24, GKM21, GJ23, JZS23, LPG20, SHL^{+}21, Yan20, ASW23, AF24, ABB23, BGS24, BLSV23, BGK21, BKN22, BK24, BMP23, BLT21, BM21b, BKW23, BCFL21, CS21, CK23a, CCM21, CD21, CKK^{+}20, CCQY24, CXY21, CXW22, CFZ23, Che23, CGI20, CKLZ24, CYZ23, DGR20, DS20, DLP21, FL24, FHSW23, HR21, HKR20, HI20, HHPT24, HLM23, HJL^{+}24, HV23, Ish20, JS24, KRW22, KV23, KLL21, LLW23, LL20, LZGZ21, LC20, LS20b, NKM22, PPX24, RV21, SX22, SHWZ23, Tsa20, Uml24, VB23b, WD22, WFL22, WZL24, WPS24, YN21, ZM24]. Time-invariant [Yan20]. time-series [CD21]. time-to-event [PPX24]. Time-Varying [BHKL24, LPG20, BHSvS21, Bot23, CM24, CHL24, FWYZ24, GPWY24, GKM21, SHL^{+}21, BGS24, BKN22, BMP23, CCM21, Che23, CKLZ24, FL24, FHSW23, HLM23, Ish20, JS24, KRW22, SHWZ23, Uml24, WPS24]. times [AS21, Che24]. timing [AFiK^{+}23, GB21a]. title [APdAV23]. Tobin [CK20]. Tobit [BD24, CKR23]. today [Yu23]. total [MM21b]. trade [BK20b]. tradeoff [HL20b]. traders [ASB20]. trading [BAFMS20, CHW24, NP22]. training [FOP23]. transform [FHW23]. transformation [BMP23, CZ20, HLL21, LZ20, YLC^{+}23]. Transformations [Kit22]. Transformed [DRT24]. transient [PLS20]. transition [CMH24]. transitions [FOP23]. transparency [Vil23]. trawl [BLSV23]. Treasury [NEFG20]. treated [Fer21, MW20]. Treatment [CK23a, HY23, KPV23, AX23, ALZ22, BKP23, Cal21, CLMZ24, CR24, CKL24, CO21, Fus24, GB21a, Han21, HY24, HK21, JPTZ23, Kéd23, KLSW23, LSZ24, MMY23, Man23, PF23, SU23, SA21, WTH24, ZD21]. treatments [BH23, VB23b, dCD23, uHS23]. trend [Ell20]. trending [RSBP23]. Trends [GG20a, CKK^{+}20, FSU20, KV23, KOEP20]. trials [CL24b]. triangular [Fen24, MMY23, PF23]. True [PW23, Hor21, Kli24]. trust [FMM^{+}22]. trustworthy [Hwa21]. Tuning [LSTW24]. TVP [CES20].

TVP-VAR [CES20]. Twenty [CSV23]. Twisted [HSHS20]. Twisting [BDFM23]. Twitter [AMMQ22]. Two [AF24, Che23, HWZW20, dCD23, AM22, BGS24, CGQ23, CHL21, FW23, HKTY23, HV23, KS20, LS23b, OJT20, WLFF24]. two-dimensional [LS23b]. Two-mode [HWZW20, WLFF24]. two-pass [AM22, BGS24]. two-sample [CGQ23]. Two-stage [AF24, CHL21, KS20]. Two-step [Che23, HV23]. Two-way [dCD23, FW23, HKTY23]. type [Boo23, DN23, GL20, HLM^{+}24, LLYZ22]. types [Kéd23].

U.S. [CKS21, NEFG20]. Uhlenbeck [WXY23]. Ultrahigh [WCWL20, CLWY24, ZZZ24]. unadjusted [LMNZ24]. unbiased [GGV20]. uncertainty [ABB^{+}22, CCM21, HSHS20, HS21a, Yu23, vGW22]. Unconditional [MIMRS24, ASZZ24, Gua21b]. Uncovering [LTY21]. Under-identification [AFiK^{+}23]. underidentification [Sen24, Win24]. Understanding [KRvdK22, LP22]. undirected [FFX20, GNP24]. Unequal [BBRSS23]. unfair [ZZ21]. unified [ZD24]. Uniform [BKW23, CLY24, FGP23, LL20, LS23b, IK20, KN24]. Union [APdAV23]. unit [HKT20, LP20b, LT20, LP23c, NW21]. units [KPR21, PY20, PY21]. unity [BP20b, BD24]. university [DDH22]. unknown [BCL24, CK24]. Unobserved [Luo20, AGL21, ABL22, ALL23, Bot20, BMP23, Bot23, BCV20, CGI21, Esc23, FLZ24, FW23, Fus24, Hub23, JS24, KST22, Lu22, LXX22, SST21]. unrelated [WGH20]. unreported [HLS21]. use [BBRSS23, CR20, MMF20]. Using [AIMM24, BGPS23, CCM21, HN21, Mac23, ASLL24, ACS20, AMMQ22, BLL22, BIJS22, CFX22, Cal21, CLMZ24, CK23b, CKLZ24, FHLL22, GHLL22, GHKOB24, HLL21, HKO^{+}23, KY22, KRvdK22, KST22, KMMS21, LX23, MM21b, MB24, RV21, SZ24]. utilities [GLX23]. utility [ILMM20, Su21]. utility-maximizing [Su21].

vaccinated [KW23]. vaccines [KW23]. Valid [HK21, Pet22]. validating [FKL21]. validation [HW22, JMS21, WCWL20, ZL23, ZZ24a]. validity [GGR24, IK20, JM21, Sun24]. value [FGP23, GG20b, Tsa20, BHS24]. Value-at-Risk [BHS24]. valued [BLSV23, CD21, CXY21, CYX^{+}23, Fus24, RW20]. values [CBN23, Kli24, ZD24]. VAR [CES20, GP23, IK22, KvD23, Pet22, FZ20, LW23b]. Variable [CGI20, AGS24, ARRS23, Dal23, GKM21, HJW24, LYZ20, LMNZ24, NSYC21, WTH24]. Variables [EL21, BHN22, BMS20, Che21, DLP21, GLT20, Hor21, HHS20, Kle21, KMS21, LSW23, LMSND22, NS21, QfLY21, WZ24, Wil20]. Variance [Gal23, HS20, MZ21, Par20, ASKM20, ATM20, CHLZ20, DGR20, DLZ21, FWYZ24, FS21, HL23, HKL22, PW23, RSV20, SS24b, YGV24]. Variance-covariance [Gal23]. variances [BMPQ22]. variate [GH23]. Variation [ZLTT22, GB21a, Tod22]. variational [LMSND22]. variations [BS21]. VARs [Cha23, CHKP23, DW23, HKO^{+}23, MPS23]. Varying [BHKL24, Bre21, LPG20, BGS24, BHSvS21, BKN22, BMP23, Bot23, CCM21, Che23, CM24, CHL24, CKLZ24, DAM21, DOT22, FWYZ24, FL24, FHSW23, GPWY24, GKM21, GJ23, HLM23, Ish20, JS24, KRW22, SHL^{+}21, SHWZ23, UWY23, Uml24, WPS24, YN21]. Vector [CCM21, FH23, Mav21, OSW21, Bog22, CCM19, CCCM22, CGL^{+}22, DS20, FHP23, FS23b, Gof24, Gua21b, GB21b, HT20, KZA20, KPT23b, Pre20, SL20, YfL21, HKST24]. vectors [Cha20, WLFM24]. Vehicle [WY21]. versus [Vil23]. via [ASZZ24, BC21, CL24b, CGQ23, CFVW24, CHL21, FHW23, LKLP20, LWY23, LNP24, Phi20, Tsa20, WD22, WCWL20, WAZ24, YYX24]. view [HS20]. vine [NKM22]. viral [HLM23]. Virtual [FZ20]. VIX [AG21]. volatilities [BH20]. Volatility [ALTZ24, CT24, KM20, LS20a, LLZ22, SKY^{+}21, WAZ24, AD21a, ASLL21, AG21, ARRS23, ACM22, Bog22, BCPV23, BLL24, BCGR21, BCFL21, CMH24, CCM19, CCM21, CCCM22, CNPR22, Cha23, CK23b, CHKP23, DTPP23, DW20, Din23, HKT20, HP24, JLZ20, LLV20, LW20, NEFG20, SKF23, TZ23, WXY23, Yan20, AG21]. volume [DRG^{+}23, DRG^{+}24].

Wage [CSV23]. wages [APdAV23, CPRR23, DKSS23, Kas22, LS23a, Sch23, Woo23]. Wald [FS23a, LLYZ22]. Wald-type [LLYZ22]. Wales [KHK20]. Walker [RW24]. WALS [DMP22]. war [JLMM21, Kim23]. warming [GG20a]. Washington [LMSW23]. Wasserstein [AIMM24]. waveform [PM24]. wavelet [BKL^{+}22]. way [FW23, HKTY23, dCD23]. weak [AM22, ACF24, CP24, CSW23, LWZ24, MISW20, TD20, VW23, WZ24]. weak-instrument [VW23]. weaker [BN23a]. weakly [CKT24]. Weibull [LNP24]. weighted [AF24, DLP21, ZHW20, Zhu23]. weighting [HHK^{+}24]. weights [QfLY21, RL23]. welfare [Kas22, KKKN21]. well [PS21a]. West [KSW24]. where [DKSS23]. WHF [Fun24]. Which [MMF20]. white [Tuv20]. Whitney [AAI24]. Who [Kas22, KW23]. Wild [LP23a, WZ24, FSU20]. will [DRG^{+}23, DRG^{+}24, LL21]. wins [Kas22]. wisdom [DJK21]. Wishart [GH23]. without [CKT24, CKL24, HS21b, Ish20]. women [JS24]. Words [GSV22]. work [PS21a]. worker [APdAV23, HKNW23, Woo23]. Workers [CSV23, EHKS23]. workforce [BDFM23].

X [IKP22].

year [CEC22]. years [CSV23, SW21a]. yield [HNZ22]. young [JS24]. you’re [DKSS23]. youth [MGSN24]. Yule [RW24].

zero [BMPQ22, CCW20, GHM20, HI20]. zero-degree [CCW20]. zoo [WLLS24].

References

Alvarez:2022:RLE

[AA22] Javier Alvarez and Manuel Arellano. Robust likelihood estimation of dynamic panel data models. Journal of Econometrics, 226(1):21–61, January 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407621000956.

Andersen:2022:LMM

[AACH22] Torben G. Andersen, Ilya Archakov, Gökhan Cebiroglu, and Nikolaus Hautsch. Local mispricing and microstructural noise: a parametric perspective. Journal of Econometrics, 230(2):510– 534, October 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621001780. See corri-

gendum [AACH23].

Andersen:2023:CLM

[AACH23] Torben G. Andersen, Ilya Archakov, Gökhan Cebiroglu, and Nikolaus Hautsch. Corrigendum to “Local mispricing and microstructural noise: a parametric perspective” [J. Econometrics 230 (2022) 510–534]. Journal of Econometrics, 232(2):598–603, February 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623000027. See [AACH22].

Almeida:2020:NAH

[AAG20] Caio Almeida, Kym Ardison, and René Garcia. Nonparametric assessment of hedge fund performance. Journal of Econometrics, 214(2):349–378, February 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619301617.

Abadie:2024:WNC

[AAI24] Alberto Abadie, Joshua Angrist, and Guido Imbens. Whitney Newey’s contributions to econometrics. Journal of Econometrics, 240(2):??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407624000344.

Altig:2022:SBU

[ABB^{+}22] David Altig, Jose Maria Barrero, Nicholas Bloom, Steven J. Davis, Brent Meyer, and Nicholas Parker. Surveying business uncertainty. Journal of Econometrics, 231(1):282–303, November 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620302785.

Augustyniak:2023:DTH

[ABB23] Maciej Augustyniak, Alexandru Badescu, and Jean-François Bégin. A discrete-time hedging framework with multiple factors and fat tails: On what matters. Journal of Econometrics, 232(2):416–444, February 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002049.

Arellano:2024:HCR

[ABBL24] Manuel Arellano, Richard Blundell, Stéphane Bonhomme, and Jack Light. Heterogeneity of consumption responses to income shocks in the presence of nonlinear persistence. Journal of Econometrics, 240(2):??, March 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623001434.

Agudze:2022:MSP

[ABCR22] Komla M. Agudze, Monica Billio, Roberto Casarin, and Francesco Ravazzolo. Markov switching panel with endogenous synchronization effects. Journal of Econometrics, 230(2):281– 298, October 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621001251.

Ando:2022:BML

[ABL22] Tomohiro Ando, Jushan Bai, and Kunpeng Li. Bayesian and maximum likelihood analysis of large-scale panel choice models with unobserved heterogeneity. Journal of Econometrics, 230(1):20–38, September 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440762100083X.

Angelini:2024:ITS

[ACF24] Giovanni Angelini, Giuseppe Cavaliere, and Luca Fanelli. An identification and testing strategy for proxy-SVARs with weak proxies. Journal of Econometrics, 238(2):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623003202.

Andrews:2020:GRE

[ACG20] Donald W. K. Andrews, Xu Cheng, and Patrik Guggenberger. Generic results for establishing the asymptotic size of confidence sets and tests. Journal of Econometrics, 218(2):496–531, October 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301470.

Andersen:2022:OTS

[ACL22] Torben G. Andersen, Chia-Lin Chang, and Shiqing Ling. Overview: Time series analysis of higher moments and distributions of financial data. Journal of Econometrics, 227(1):1– 3, March 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621002335.

Asai:2022:RME

[ACM22] Manabu Asai, Chia-Lin Chang, and Michael McAleer. Realized matrix-exponential stochastic volatility with asymmetry, long memory and higher-moment spillovers. Journal of Econometrics, 227(1):285–304, March 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001809.

Amengual:2020:TDA

[ACS20] Dante Amengual, Marine Carrasco, and Enrique Sentana. Testing distributional assumptions using a continuum of moments. Journal of Econometrics, 218(2):655–689, October 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301536.

Ahsan:2021:SEI

[AD21a] Md. Nazmul Ahsan and Jean-Marie Dufour. Simple estimators and inference for higher-order stochastic volatility models. Journal of Econometrics, 224(1):181–197, September 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001007.

Antoine:2021:REE

[AD21b] Bertille Antoine and Prosper Dovonon. Robust estimation with exponentially tilted Hellinger distance. Journal of Econometrics, 224(2):330–344, October 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303705.

Antolin-Diaz:2024:ANE

[ADDP24] Juan Antolín-Díaz, Thomas Drechsel, and Ivan Petrella. Advances in nowcasting economic activity: the role of heterogeneous dynamics and fat tails. Journal of Econometrics, 238(2): ??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623003500.

Aknouche:2024:TSW

[AF24] Abdelhakim Aknouche and Christian Francq. Two-stage weighted least squares estimator of the conditional mean of observation-driven time series models. Journal of Econometrics, 237(2B):??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S030440762100213X.

Ackerberg:2023:UIS

[AFiK^{+}23] Daniel A. Ackerberg, Garth Frazer, Kyoo il Kim, Yao Luo, and Yingjun Su. Under-identification of structural models based on timing and information set assumptions. Journal of Econometrics, 237(1):??, November 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623001574.

Andreou:2021:PVV

[AG21] Elena Andreou and Eric Ghysels. Predicting the VIX and the volatility risk premium: the role of short-run funding spreads Volatility Factors. Journal of Econometrics, 220(2):366–398, February 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620301251.

Aguirregabiria:2021:SSU

[AGL21] Victor Aguirregabiria, Jiaying Gu, and Yao Luo. Sufficient statistics for unobserved heterogeneity in structural dynamic logit models. Journal of Econometrics, 223(2):280–311, August 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303286.

Alonso:2020:RPB

[AGP20] Andrés M. Alonso, Pedro Galeano, and Daniel Peña. A robust procedure to build dynamic factor models with cluster structure. Journal of Econometrics, 216(1):35–52, May 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620300099.

Abadie:2024:IVE

[AGS24]  Alberto Abadie, Jiaying Gu, and Shu Shen.  Instru-

mental variable estimation with first-stage heterogeneity. Journal of Econometrics, 240(2):??, March 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623000702.

An:2021:DDU

[AHX21] Yonghong An, Yingyao Hu, and Ruli Xiao. Dynamic decisions under subjective expectations: a structural analysis. Journal of Econometrics, 222(1):645–675, May 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407620302414.

An:2023:SAS

[AHZ23] Yonghong An, Shengjie Hong, and Daiqiang Zhang. A structural analysis of simple contracts. Journal of Econometrics, 236(2):??, October 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407623001501.

Athey:2022:DBA

[AI22] Susan Athey and Guido W. Imbens. Design-based analysis in difference-in-differences settings with staggered adoption. Journal of Econometrics, 226(1):62–79, January 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621000488.

Athey:2024:UWG

[AIMM24] Susan Athey, Guido W. Imbens, Jonas Metzger, and Evan Munro. Using Wasserstein Generative Adversarial Networks for the design of Monte Carlo simulations. Journal of Econometrics, 240(2):??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407621000440.

Angrist:2024:OIR

[AK24] Joshua Angrist and Michal Kolesár. One instrument to rule them all: the bias and coverage of just-ID IV. Journal of Econometrics, 240(2):??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407623000295.

Aguiar:2023:PPP

[AKA23] Victor H. Aguiar, Nail Kashaev, and Roy Allen. Prices, profits, proxies, and production. Journal of Econometrics, 235(2):666– 693, August 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407622001300.

Andrews:2021:IAE

[AKM21] Isaiah Andrews, Toru Kitagawa, and Adam McCloskey. Inference after estimation of breaks. Journal of Econometrics, 224(1):39–59, September 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620302591.

Adusumilli:2020:IDF

[AKOW20] Karun Adusumilli, Daisuke Kurisu, Taisuke Otsu, and Yoon-Jae Whang. Inference on distribution functions under measurement error. Journal of Econometrics, 215(1):131–164, March 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761930199X.

Aradillas-Lopez:2021:CSE

[AL21] Andrés Aradillas-López. Computing semiparametric efficiency bounds in discrete choice models with strategic-interactions and rational expectations. Journal of Econometrics, 221(1):25– 42, March 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620300452.

Antoine:2023:IRN

[AL23] Bertille Antoine and Pascal Lavergne. Identification-robust nonparametric inference in a linear IV model. Journal of Econometrics, 235(1):1–24, July 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440762200046X.

Aradillas-Lopez:2024:IMP

[AL24] Andres Aradillas-Lopez. Inference in models with partially identified control functions. Journal of Econometrics, 238(1):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623002695.

Ando:2023:SPQ

[ALL23] Tomohiro Ando, Kunpeng Li, and Lina Lu. A spatial panel quantile model with unobserved heterogeneity. Journal of Econometrics, 232(1):191–213, January 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002323.

Abadir:2023:GDF

[ALP23] Karim M. Abadir, Alessandra Luati, and Paolo Paruolo. GARCH density and functional forecasts. Journal of Econometrics, 235(2):470–483, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407622001154.

Aradillas-Lopez:2022:IOR

[ALR22] Andrés Aradillas-López and Adam M. Rosen. Inference in ordered response games with complete information. Journal of Econometrics, 226(2):451–476, February 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002347.

Andersen:2024:VMP

[ALTZ24] Torben G. Andersen, Yingying Li, Viktor Todorov, and Bo Zhou. Volatility measurement with pockets of extreme return persistence. Journal of Econometrics, 237(2C):??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407620303924.

Ai:2022:EIC

[ALZ22] Chunrong Ai, Oliver Linton, and Zheng Zhang. Estimation and inference for the counterfactual distribution and quantile functions in continuous treatment models. Journal of Econometrics, 228(1):39–61, May 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621000543.

Arcidiacono:2020:IDD

[AM20] Peter Arcidiacono and Robert A. Miller. Identifying dynamic discrete choice models off short panels. Journal of Econometrics, 215(2):473–485, April 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619302064.

Anatolyev:2022:FMM

[AM22] Stanislav Anatolyev and Anna Mikusheva. Factor models with many assets: Strong factors, weak factors, and the two-pass procedure. Journal of Econometrics, 229(1):103–126, July 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621000130.

Angelico:2022:CWM

[AMMQ22] Cristina Angelico, Juri Marcucci, Marcello Miccoli, and Filippo Quarta. Can we measure inflation expectations using Twitter? Journal of Econometrics, 228(2):259–277, June 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622000227.

Aruoba:2022:SOB

[AMSV22] S. Boragan Aruoba, Marko Mlikota, Frank Schorfheide, and Sergio Villalvazo. SVARs with occasionally-binding constraints. Journal of Econometrics, 231(2):477–499, December 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002487.

Anonymous:2020:AAA

[Ano20a] Anonymous. Annual award announcement. Journal of Econometrics, 214(2):iii, February 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407619302684.

Anonymous:2020:EBa

[Ano20b] Anonymous. Editorial Board. Journal of Econometrics, 214(1): ii, January 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407619302313.

Anonymous:2020:EBb

[Ano20c] Anonymous. Editorial Board. Journal of Econometrics, 214(2): ii, February 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407619302659.

Anonymous:2020:EBc

[Ano20d] Anonymous. Editorial Board. Journal of Econometrics, 215(1): ii, March 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620300312.

Anonymous:2020:EBd

[Ano20e] Anonymous. Editorial Board. Journal of Econometrics, 215 (2):ii, April 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030440762030049X.

Anonymous:2020:EBe

[Ano20f] Anonymous. Editorial Board. Journal of Econometrics, 216 (1):ii, May 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620300695.

Anonymous:2020:EBf

[Ano20g] Anonymous. Editorial Board. Journal of Econometrics, 216 (2):ii, June 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620301147.

Anonymous:2020:EBg

[Ano20h] Anonymous. Editorial Board. Journal of Econometrics, 217 (1):ii, July 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620301627.

Anonymous:2020:EBh

[Ano20i] Anonymous. Editorial Board. Journal of Econometrics, 217(2): ii, August 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620301779.

Anonymous:2020:EBi

[Ano20j] Anonymous. Editorial Board. Journal of Econometrics, 218(1): ii, September 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620301871.

Anonymous:2020:EBj

[Ano20k] Anonymous. Editorial Board. Journal of Econometrics, 218(2): ii, October 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030440762030289X.

Anonymous:2020:EBk

[Ano20l] Anonymous. Editorial Board. Journal of Econometrics, 219(1): ii, November 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620303225.

Anonymous:2020:EBl

[Ano20m] Anonymous. Editorial Board. Journal of Econometrics, 219(2): ii, December 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620303596.

Anonymous:2020:PN

[Ano20n] Anonymous. Pages 1–200 (November 2020). Journal of Econometrics, 219(1):??, November 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2020:PJb

[Ano20o] Anonymous. Pages 1–202 (July 2020). Journal of Econometrics, 217(1):??, July 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2020:PS

[Ano20p] Anonymous. Pages 1–242 (September 2020). Journal of Econometrics, 218(1):??, September 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2020:PM

[Ano20q] Anonymous. Pages 1–304 (March 2020). Journal of Econometrics, 215(1):??, March 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2020:PO

[Ano20r] Anonymous. Pages 243–770 (October 2020). Journal of Econometrics, 218(2):??, October 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2020:PFa

[Ano20s] Anonymous. Pages 295–540 (February 2020). Journal of Econometrics, 214(2):??, February 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2020:PA

[Ano20t] Anonymous. Pages 305–632 (April 2020). Journal of Econometrics, 215(2):??, April 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2020:PJa

[Ano20u] Anonymous. Pages 327–536 (June 2020). Journal of Econometrics, 216(2):??, June 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2021:EBa

[Ano21a] Anonymous. Editorial Board. Journal of Econometrics, 220(1): ii, January 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620303845.

Anonymous:2021:EBb

[Ano21b] Anonymous. Editorial Board. Journal of Econometrics, 220(2): ii, February 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621000051.

Anonymous:2021:EBc

[Ano21c] Anonymous. Editorial Board. Journal of Econometrics, 221(1): ii, March 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030440762100021X.

Anonymous:2021:EBd

[Ano21d] Anonymous. Editorial Board. Journal of Econometrics, 221 (2):ii, April 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407621000324.

Anonymous:2021:EBe

[Ano21e] Anonymous. Editorial Board. Journal of Econometrics, 222(1): ii, ???? 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407621000634.

Anonymous:2021:EBf

[Ano21f] Anonymous. Editorial Board. Journal of Econometrics, 222(1): ii, ???? 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407621000774.

Anonymous:2021:EBg

[Ano21g] Anonymous. Editorial Board. Journal of Econometrics, 222 (1):ii, May 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407621000919.

Anonymous:2021:EBh

[Ano21h] Anonymous. Editorial Board. Journal of Econometrics, 222 (2):ii, June 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621001068.

Anonymous:2021:EBi

[Ano21i] Anonymous. Editorial Board. Journal of Econometrics, 223 (1):ii, July 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621001196.

Anonymous:2021:EBj

[Ano21j] Anonymous. Editorial Board. Journal of Econometrics, 223(2): ii, August 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030440762100138X.

Anonymous:2021:EBk

[Ano21k] Anonymous. Editorial Board. Journal of Econometrics, 224(1): ii, September 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621001676.

Anonymous:2021:EBl

[Ano21l] Anonymous. Editorial Board. Journal of Econometrics, 224(2): ii, October 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621002001.

Anonymous:2021:EBm

[Ano21m] Anonymous. Editorial Board. Journal of Econometrics, 225(1): ii, November 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621002098.

Anonymous:2021:EBn

[Ano21n] Anonymous. Editorial Board. Journal of Econometrics, 225(2): ii, December 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621002384.

Anonymous:2021:NE

[Ano21o] Anonymous. A note from the Editors. Journal of Econometrics, 225(2):131, December 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407621002396.

Anonymous:2021:PJb

[Ano21p] Anonymous. Pages 1–276 (July 2021). Journal of Econometrics, 223(1):??, July 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2021:PMa

[Ano21q] Anonymous. Pages 1–336 (March 2021). Journal of Econometrics, 221(1):??, March 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2021:PO

[Ano21r] Anonymous. Pages 245–466 (October 2021). Journal of Econometrics, 224(2):??, October 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2021:PA

[Ano21s] Anonymous. Pages 337–676 (April 2021). Journal of Econometrics, 221(2):??, April 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2021:PMb

[Ano21t] Anonymous. Pages 601–860 (May 2021). Journal of Econometrics, 222(1):??, May 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2021:PJa

[Ano21u] Anonymous. Pages 861–1108 (June 2021). Journal of Econometrics, 222(2):??, June 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2022:EBa

[Ano22a] Anonymous. Editorial Board. Journal of Econometrics, 226(1): ii, January 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621002694.

Anonymous:2022:EBb

[Ano22b] Anonymous. Editorial Board. Journal of Econometrics, 226(2): ii, February 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621002979.

Anonymous:2022:EBc

[Ano22c] Anonymous. Editorial Board. Journal of Econometrics, 227(1): ii, March 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407622000148.

Anonymous:2022:EBd

[Ano22d] Anonymous. Editorial Board. Journal of Econometrics, 227 (2):ii, April 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407622000264.

Anonymous:2022:EBe

[Ano22e] Anonymous. Editorial Board. Journal of Econometrics, 228 (1):ii, May 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407622000513.

Anonymous:2022:EBf

[Ano22f] Anonymous. Editorial Board. Journal of Econometrics, 228 (2):ii, June 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407622000744.

Anonymous:2022:EBg

[Ano22g] Anonymous. Editorial Board. Journal of Econometrics, 229 (1):ii, July 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407622000823.

Anonymous:2022:EBh

[Ano22h] Anonymous. Editorial Board. Journal of Econometrics, 229(2): ii, August 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030440762200104X.

Anonymous:2022:EBi

[Ano22i] Anonymous. Editorial Board. Journal of Econometrics, 230(1): ii, September 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407622001105.

Anonymous:2022:EBj

[Ano22j] Anonymous. Editorial Board. Journal of Econometrics, 230(2): ii, October 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030440762200135X.

Anonymous:2022:EBk

[Ano22k] Anonymous. Editorial Board. Journal of Econometrics, 231(1): ii, November 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407622001415.

Anonymous:2022:EBl

[Ano22l] Anonymous. Editorial Board. Journal of Econometrics, 231(2): ii, December 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407622001749.

Anonymous:2022:PJb

[Ano22m] Anonymous. Pages 1–218 (July 2022). Journal of Econometrics, 229(1):??, July 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2022:PJa

[Ano22n] Anonymous. Pages 177–398 (June 2022). Journal of Econometrics, 228(2):??, June 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2022:PF

[Ano22o] Anonymous. Pages 205–498 (February 2022). Journal of Econometrics, 226(2):??, February 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2022:PAb

[Ano22p] Anonymous. Pages 219–452 (August 2022). Journal of Econometrics, 229(2):??, August 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2022:PO

[Ano22q] Anonymous. Pages 221–558 (October 2022). Journal of Econometrics, 230(2):??, October 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2022:PAa

[Ano22r] Anonymous. Pages 305–518 (April 2022). Journal of Econometrics, 227(2):??, April 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2023:JEA

[Ano23a] Anonymous. Journal of Econometrics awards announcement. Journal of Econometrics, 234(1):1–2, May 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623000799.

Anonymous:2023:D

[Ano23b] Anonymous. December 2023. Journal of Econometrics, 237(2): ??, ???? 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2023:EBa

[Ano23c] Anonymous. Editorial Board. Journal of Econometrics, 232(1): ii, January 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407622001981.

Anonymous:2023:EBb

[Ano23d] Anonymous. Editorial Board. Journal of Econometrics, 232(2): ii, February 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030440762300009X.

Anonymous:2023:EBc

[Ano23e] Anonymous. Editorial Board. Journal of Econometrics, 233(1): ii, March 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623000520.

Anonymous:2023:EBd

[Ano23f] Anonymous. Editorial Board. Journal of Econometrics, 233 (2):ii, April 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623000672.

Anonymous:2023:EBe

[Ano23g] Anonymous. Editorial Board. Journal of Econometrics, 234 (1):ii, May 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623000891.

Anonymous:2023:EBf

[Ano23h] Anonymous. Editorial Board. Journal of Econometrics, 234 (2):ii, June 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623001239.

Anonymous:2023:EBg

[Ano23i] Anonymous. Editorial Board. Journal of Econometrics, 235 (1):ii, July 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623001367.

Anonymous:2023:EBh

[Ano23j] Anonymous. Editorial Board. Journal of Econometrics, 235(2): ii, August 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030440762300177X.

Anonymous:2023:EBi

[Ano23k] Anonymous. Editorial Board. Journal of Econometrics, 236(1): ??, September 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623002014.

Anonymous:2023:EBj

[Ano23l] Anonymous. Editorial Board. Journal of Econometrics, 236(2): ??, October 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623002221.

Anonymous:2023:EBk

[Ano23m] Anonymous. Editorial Board. Journal of Econometrics, 237(1): ??, November 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623002555.

Anonymous:2023:IHP

[Ano23n] Anonymous. Introducing how-to papers. Journal of Econometrics, 232(2):271, February 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407623000015.

Anonymous:2023:N

[Ano23o] Anonymous. November 2023. Journal of Econometrics, 237(1): ??, November 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2023:O

[Ano23p] Anonymous. October 2023. Journal of Econometrics, 236(2): ??, October 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2023:PJa

[Ano23q] Anonymous. Pages 1–270 (January 2023). Journal of Econometrics, 232(1):??, January 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2023:PJc

[Ano23r] Anonymous. Pages 1–324 (July 2023). Journal of Econometrics, 235(1):??, July 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2023:PMa

[Ano23s] Anonymous. Pages 1–332 (March 2023). Journal of Econometrics, 233(1):??, March 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2023:PMb

[Ano23t] Anonymous. Pages 1–370 (May 2023). Journal of Econometrics, 234(1):??, May 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2023:PF

[Ano23u] Anonymous. Pages 271–604 (February 2023). Journal of Econometrics, 232(2):??, February 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2023:PAb

[Ano23v] Anonymous. Pages 325–2294 (August 2023). Journal of Econometrics, 235(2):??, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2023:PAa

[Ano23w] Anonymous. Pages 333–714 (April 2023). Journal of Econometrics, 233(2):??, April 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2023:PJb

[Ano23x] Anonymous. Pages 371–776 (June 2023). Journal of Econometrics, 234(2):??, June 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2023:S

[Ano23y] Anonymous. September 2023. Journal of Econometrics, 236 (1):??, September 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2024:Aa

[Ano24a] Anonymous. April 2024. Journal of Econometrics, 241(1):??, April 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2024:Ab

[Ano24b] Anonymous. April 2024. Journal of Econometrics, 241(2):??, April 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2024:CE

[Ano24c] Anonymous. Climate econometrics. Journal of Econometrics, 239(??):??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2024:Da

[Ano24d] Anonymous. December 2023. Journal of Econometrics, 237 (2A):??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2024:Db

[Ano24e] Anonymous. December 2023. Journal of Econometrics, 237 (2B):??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2024:Dc

[Ano24f] Anonymous. December 2023. Journal of Econometrics, 237 (2C):??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2024:EBa

[Ano24g] Anonymous. Editorial Board. Journal of Econometrics, 237 (2A):??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0304407623003263.

Anonymous:2024:EBb

[Ano24h] Anonymous. Editorial Board. Journal of Econometrics, 237 (2B):??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0304407623003093.

Anonymous:2024:EBc

[Ano24i] Anonymous. Editorial Board. Journal of Econometrics, 237 (2C):??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0304407623003159.

Anonymous:2024:EBd

[Ano24j] Anonymous. Editorial Board. Journal of Econometrics, 238(1): ??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623003354.

Anonymous:2024:EBe

[Ano24k] Anonymous. Editorial Board. Journal of Econometrics, 238(2): ??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407624000058.

Anonymous:2024:EBf

[Ano24l] Anonymous. Editorial Board. Journal of Econometrics, 239(??): ??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407624000113.

Anonymous:2024:EBg

[Ano24m] Anonymous. Editorial Board. Journal of Econometrics, 239(2): ??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407624000526.

Anonymous:2024:EBh

[Ano24n] Anonymous. Editorial Board. Journal of Econometrics, 240(1): ??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407624000599.

Anonymous:2024:EBi

[Ano24o] Anonymous. Editorial Board. Journal of Econometrics, 240(2): ??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407624000447.

Anonymous:2024:EBj

[Ano24p] Anonymous. Editorial Board. Journal of Econometrics, 241(1): ??, April 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407624001040.

Anonymous:2024:EBk

[Ano24q] Anonymous. Editorial Board. Journal of Econometrics, 241(2): ??, April 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407624001246.

Anonymous:2024:F

[Ano24r] Anonymous. February 2024. Journal of Econometrics, 239(2): ??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2024:Ja

[Ano24s] Anonymous. January 2024. Journal of Econometrics, 238(1): ??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2024:Jb

[Ano24t] Anonymous. January 2024. Journal of Econometrics, 238(2): ??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2024:Ma

[Ano24u] Anonymous. March 2024. Journal of Econometrics, 240(1): ??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2024:Mb

[Ano24v] Anonymous. March 2024. Journal of Econometrics, 240(2): ??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Addison:2023:UMD

[APdAV23] John T. Addison, Pedro Portugal, and Hugo de Almeida Vilares. Union membership density and wages: the role of worker, firm, and job-title heterogeneity. Journal of Econometrics, 233(2):612–632, April 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407621003006.

Antoine:2020:TIS

[AR20] Bertille Antoine and Eric Renault. Testing identification strength. Journal of Econometrics, 218(2):271–293, October 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301378.

Allen:2022:LCB

[AR22] Roy Allen and John Rehbeck. Latent complementarity in bundles models. Journal of Econometrics, 228(2):322–341, June 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002438.

Aristodemou:2021:SIP

[Ari21] Eleni Aristodemou. Semiparametric identification in panel data discrete response models. Journal of Econometrics, 220(2):253– 271, February 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620301202.

Arias:2023:MFV

[ARRS23] Jonas E. Arias, Juan F. Rubio-Ramírez, and Minchul Shin. Macroeconomic forecasting and variable ordering in multivariate stochastic volatility models. Journal of Econometrics, 235(2):1054–1086, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407622001695.

Arias:2021:IBP

[ARRW21] Jonas E. Arias, Juan F. Rubio-Ramírez, and Daniel F. Waggoner. Inference in Bayesian proxy-SVARs. Journal of Econometrics, 225(1):88–106, November 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303985.

Andersen:2023:ICS

[ARTT23] Torben G. Andersen, Raul Riva, Martin Thyrsgaard, and Viktor Todorov. Intraday cross-sectional distributions of systematic risk. Journal of Econometrics, 235(2):1394–1418, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622002032.

Abbring:2021:LMH

[AS21] Jaap H. Abbring and Tim Salimans. The likelihood of mixed hitting times. Journal of Econometrics, 223(2):361–375, August 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001093.

Anatolyev:2023:TMR

[AS23] Stanislav Anatolyev and Mikkel Sølvsten. Testing many restrictions under heteroskedasticity. Journal of Econometrics, 236(1): ??, September 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623001677.

Ait-Sahalia:2020:HFT

[ASB20] Yacine Aït-Sahalia and Celso Brunetti. High frequency traders and the price process. Journal of Econometrics, 217(1):20– 45, July 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407619302428.

Ait-Sahalia:2020:TSE

[ASKM20] Yacine Aït-Sahalia, Mustafa Karaman, and Loriano Mancini. The term structure of equity and variance risk premia. Journal of Econometrics, 219(2):204–230, December 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762030097X.

Ait-Sahalia:2020:HFF

[ASKX20] Yacine Aït-Sahalia, Ilze Kalnina, and Dacheng Xiu. Highfrequency factor models and regressions. Journal of Econometrics, 216(1):86–105, May 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620300129.

Ait-Sahalia:2021:CFI

[ASLL21] Yacine Aït-Sahalia, Chenxu Li, and Chen Xu Li. Closedform implied volatility surfaces for stochastic volatility models with jumps. Journal of Econometrics, 222(1):364–392, ???? 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-

6895 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0304407620301962.

Ait-Sahalia:2024:MLE

[ASLL24] Yacine Aït-Sahalia, Chenxu Li, and Chen Xu Li. Maximum likelihood estimation of latent Markov models using closed-form approximations. Journal of Econometrics, 240(2):??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303389.

Ait-Sahalia:2024:HFM

[ASS24] Yacine Aït-Sahalia and Mehmet Saglam. High frequency market making: the role of speed. Journal of Econometrics, 239(2):??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623000581.

Arvanitis:2020:STM

[AST20] Stelios Arvanitis, Olivier Scaillet, and Nikolas Topaloglou. Spanning tests for Markowitz stochastic dominance. Journal of Econometrics, 217(2):291–311, August 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619302519.

Adamek:2023:LIH

[ASW23] Robert Adamek, Stephan Smeekes, and Ines Wilms. Lasso inference for high-dimensional time series. Journal of Econometrics, 235(2):1114–1143, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407622001804.

Ai:2024:TUC

[ASZZ24] Chunrong Ai, Li-Hsien Sun, Zheng Zhang, and Liping Zhu. Testing unconditional and conditional independence via mutual information. Journal of Econometrics, 240(2):??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001609.

Akashi:2020:RCT

[ATM20] Fumiya Akashi, Masanobu Taniguchi, and Anna Clara Monti. Robust causality test of infinite variance processes. Journal of Econometrics, 216(1):235–245, May 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762030021X.

Andersen:2024:PMF

[ATTX24] Torben G. Andersen, Robert Taylor, Allan Timmermann, and Dacheng Xiu. Predictive modeling of financial data. Journal of Econometrics, 237(2C):??, December 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407623002129.

Andersen:2021:TRR

[ATU21] Torben G. Andersen, Viktor Todorov, and Masato Ubukata. Tail risk and return predictability for the Japanese equity market. Journal of Econometrics, 222(1):344–363, ???? 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407620301950.

Andersen:2021:CIP

[AV21] Torben G. Andersen and Rasmus T. Varneskov. Consistent inference for predictive regressions in persistent economic systems. Journal of Econometrics, 224(1):215–244, September 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303547.

Andersen:2022:TPI

[AV22] Torben G. Andersen and Rasmus T. Varneskov. Testing for parameter instability and structural change in persistent predictive regressions. Journal of Econometrics, 231(2):361–386, December 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621002529.

Abrevaya:2023:ETE

[AX23] Jason Abrevaya and Haiqing Xu. Estimation of treatment effects under endogenous heteroskedasticity. Journal of Econometrics, 234(2):451–478, June 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001500.

Barone-Adesi:2020:OMT

[BAFMS20] Giovanni Barone-Adesi, Nicola Fusari, Antonietta Mira, and Carlo Sala. Option market trading activity and the estimation of the pricing kernel: a Bayesian approach. Journal of Econometrics, 216(2):430–449, June 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761930226X.

Bai:2023:WRM

[Bai23] Yuehao Bai. Why randomize? Minimax optimality under permutation invariance. Journal of Econometrics, 232(2):565–575, February 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621002566.

Bai:2024:LAD

[Bai24] Jushan Bai. Likelihood approach to dynamic panel models with interactive effects. Journal of Econometrics, 240(1):??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623003524.

Babii:2024:MLP

[BBGS24] Andrii Babii, Ryan T. Ball, Eric Ghysels, and Jonas Striaukas. Machine learning panel data regressions with heavytailed dependent data: Theory and application. Journal of Econometrics, 237(2C):??, December 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407622001282.

Biroli:2022:PBA

[BBRR22] Pietro Biroli, Teodora Boneva, Akash Raja, and Christopher Rauh. Parental beliefs about returns to child health investments. Journal of Econometrics, 231(1):33–57, November 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620302712.

Bana:2023:UUS

[BBRSS23] Sarah Bana, Kelly Bedard, Maya Rossin-Slater, and Jenna Stearns. Unequal use of social insurance benefits: the role of employers. Journal of Econometrics, 233(2):633–660, April 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622000628.

Bugni:2021:TCD

[BC21] Federico A. Bugni and Ivan A. Canay. Testing continuity of a density via g-order statistics in the regression discontinuity design. Journal of Econometrics, 221(1):138–159, March 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620300579.

Buccheri:2021:CTL

[BCFL21] Giuseppe Buccheri, Fulvio Corsi, Franco Flandoli, and Giulia Livieri. The continuous-time limit of score-driven volatility models. Journal of Econometrics, 221(2):655–675, April 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-

6895 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0304407620302669.

Boswijk:2021:BNS

[BCGR21] H. Peter Boswijk, Giuseppe Cavaliere, Iliyan Georgiev, and Anders Rahbek. Bootstrapping non-stationary stochastic volatility. Journal of Econometrics, 224(1):161–180, September 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621000282.

Bauwens:2023:WML

[BCL23] Luc Bauwens, Guillaume Chevillon, and Sébastien Laurent. We modeled long memory with just one lag! Journal of Econometrics, 236(1):??, September 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623001616.

Bai:2024:SEP

[BCL24] Jushan Bai, Sung Hoon Choi, and Yuan Liao. Standard errors for panel data models with unknown clusters. Journal of Econometrics, 240(2):??, March 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303341.

Bolko:2023:GAE

[BCPV23] Anine E. Bolko, Kim Christensen, Mikko S. Pakkanen, and Bezirgen Veliyev. A GMM approach to estimate the roughness of stochastic volatility. Journal of Econometrics, 235(2):745– 778, August 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407622001476.

Bruns:2020:MMG

[BCS20] Stephan B. Bruns, Zsuzsanna Csereklyei, and David I. Stern. A multicointegration model of global climate change. Journal of Econometrics, 214(1):175–197, January 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619301137.

Boudt:2020:NCE

[BCV20] Kris Boudt, Dries Cornilly, and Tim Verdonck. Nearest comoment estimation with unobserved factors. Journal of Econometrics, 217(2):381–397, August 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619302556.

Bykhovskaya:2024:LUD

[BD24] Anna Bykhovskaya and James A. Duffy. The local to unity dynamic tobit model. Journal of Econometrics, 241(2):??, April 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407624001106.

Barth:2023:TDC

[BDFM23] Erling Barth, James C. Davis, Richard B. Freeman, and Kristina McElheran. Twisting the demand curve: Digitalization and the older workforce. Journal of Econometrics, 233(2):443– 467, April 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621003018.

Bai:2024:LRT

[BDH24] Jushan Bai, Jiangtao Duan, and Xu Han. The likelihood ratio test for structural changes in factor models. Journal of Econometrics, 238(2):??, January 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623003470.

Beaulieu:2023:IRB

[BDKM23] Marie-Claude Beaulieu, Jean-Marie Dufour, Lynda Khalaf, and Olena Melin. Identification-robust beta pricing, spanning, mimicking portfolios, and the benchmark neutrality of catastrophe bonds. Journal of Econometrics, 236(1):??, September 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623001586.

Boudt:2023:EBA

[BDSV23] Kris Boudt, Kirill Dragun, Orimar Sauri, and Steven Vanduffel. ETF Basket-Adjusted Covariance estimation. Journal of Econometrics, 235(2):1144–1171, August 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001816.

Bei:2024:LLB

[Bei24] Xinyue Bei. Local linearization based subvector inference in moment inequality models. Journal of Econometrics, 238(1): ??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623002658.

Bertanha:2020:RDD

[Ber20] Marinho Bertanha. Regression discontinuity design with many thresholds. Journal of Econometrics, 218(1):216–241, September 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620300361.

Bobba:2022:SPA

[BF22] Matteo Bobba and Veronica Frisancho. Self-perceptions about academic achievement: Evidence from Mexico City. Journal of Econometrics, 231(1):58–73, November 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620302724.

Blasques:2023:QSD

[BFL23] F. Blasques, Christian Francq, and Sébastien Laurent. Quasi score-driven models. Journal of Econometrics, 234(1):251– 275, May 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030440762200001X.

Blasques:2024:ACB

[BFL24] F. Blasques, Christian Francq, and Sébastien Laurent. Autoregressive conditional betas. Journal of Econometrics, 238(2): ??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623003469.

Bobba:2021:LMS

[BFLT21] Matteo Bobba, Luca Flabbi, Santiago Levy, and Mauricio Tejada. Labor market search, informality, and on-the-job human capital accumulation. Journal of Econometrics, 223(2):433– 453, August 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620303304.

Bruck:2023:CCT

[BFM23] Florian Brück, Jean-David Fermanian, and Aleksey Min. A corrected Clarke test for model selection and beyond. Journal of Econometrics, 235(1):105–132, July 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622000446.

Bugni:2023:IUC

[BG23] Federico A. Bugni and Mengsi Gao. Inference under covariateadaptive randomization with imperfect compliance. Journal of Econometrics, 237(1):??, November 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002130.

Blasques:2021:MOO

[BGK21] F. Blasques, P. Gorgi, and S. J. Koopman. Missing observations in observation-driven time series models. Journal of Econometrics, 221(2):542–568, April 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0304407620302670.

Barkley:2021:BFA

[BGM21] Aaron Barkley, Joachim R. Groeger, and Robert A. Miller. Bidding frictions in ascending auctions. Journal of Econometrics, 223(2):376–400, August 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620303250.

Bang:2023:UMR

[BGPS23] Minji Bang, Wayne Yuan Gao, Andrew Postlewaite, and Holger Sieg. Using monotonicity restrictions to identify models with partially latent covariates. Journal of Econometrics, 235(2):892–921, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407622001555.

Bakalli:2024:PTP

[BGS24] Gaetan Bakalli, Stéphane Guerrier, and Olivier Scaillet. A penalized two-pass regression to predict stock returns with timevarying risk premia. Journal of Econometrics, 237(2C):??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407622002147.

Barigozzi:2020:GDF

[BH20] Matteo Barigozzi and Marc Hallin. Generalized dynamic factor models and volatilities: Consistency, rates, and prediction intervals. Journal of Econometrics, 216(1):4–34, May 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620300087.

Breunig:2021:NRS

[BH21] Christoph Breunig and Peter Haan. Nonparametric regression with selectively missing covariates. Journal of Econometrics, 223(1):28–52, July 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620303183.

Balat:2023:MTS

[BH23] Jorge F. Balat and Sukjin Han. Multiple treatments with strategic substitutes. Journal of Econometrics, 234(2):732–757, June 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001671.

Bu:2021:DCI

[BHK21] Ruijun Bu, Kaddour Hadri, and Dennis Kristensen. Diffusion copulas: Identification and estimation. Journal of Econometrics, 221(2):616–643, April 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0304407620302104.

Blasques:2024:TVP

[BHKL24] F. Blasques, A. C. Harvey, S. J. Koopman, and A. Lucas. Time-varying parameters in econometrics: the editor’s foreword. Journal of Econometrics, 237(2B):??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-

6895 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0304407623001173.

Borowska:2020:PCP

[BHKvD20] Agnieszka Borowska, Lennart Hoogerheide, Siem Jan Koopman, and Herman K. van Dijk. Partially censored posterior for robust and efficient risk evaluation. Journal of Econometrics, 217(2):335–355, August 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407619302532.

Barigozzi:2024:ITG

[BHLZ24] Matteo Barigozzi, Marc Hallin, Matteo Luciani, and Paolo Zaffaroni. Inferential theory for generalized dynamic factor models. Journal of Econometrics, 239(2):??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623000593.

Belloni:2022:HDL

[BHN22] Alexandre Belloni, Christian Hansen, and Whitney Newey. High-dimensional linear models with many endogenous variables. Journal of Econometrics, 228(1):4–26, May 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002220.

Bai:2020:EIC

[BHS20] Jushan Bai, Xu Han, and Yutang Shi. Estimation and inference of change points in high-dimensional factor models. Journal of Econometrics, 219(1):66–100, November 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301172.

Beutner:2024:RBC

[BHS24] Eric Beutner, Alexander Heinemann, and Stephan Smeekes. A residual bootstrap for conditional Value-at-Risk. Journal of Econometrics, 238(2):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0304407623002701.

Barigozzi:2021:TVG

[BHSvS21] Matteo Barigozzi, Marc Hallin, Stefano Soccorsi, and Rainer von Sachs. Time-varying general dynamic factor models and the measurement of financial connectedness. Journal of Econometrics, 222(1):324–343, ???? 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407620301949.

Byrne:2022:IFI

[BIJS22] David P. Byrne, Susumu Imai, Neelam Jain, and Vasilis Sarafidis. Instrument-free identification and estimation of differentiated products models using cost data. Journal of Econometrics, 228(2):278–301, June 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622000021.

Bai:2024:CAE

[BJR^{+}24] Yuehao Bai, Liang Jiang, Joseph P. Romano, Azeem M. Shaikh, and Yichong Zhang. Covariate adjustment in experiments with matched pairs. Journal of Econometrics, 241(1):??, April 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407624000861.

Bontemps:2020:GAI

[BK20a] Christian Bontemps and Rohit Kumar. A geometric approach to inference in set-identified entry games. Journal of Econometrics, 218(2):373–389, October 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762030141X.

Brauning:2020:DFN

[BK20b] Falk Bräuning and Siem Jan Koopman. The dynamic factor network model with an application to international trade. Journal of Econometrics, 216(2):494–515, June 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619302398.

Babii:2023:IRD

[BK23] Andrii Babii and Rohit Kumar. Isotonic regression discontinuity designs. Journal of Econometrics, 234(2):371–393, June 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621000506.

Blevins:2024:NPL

[BK24] Jason R. Blevins and Minhae Kim. Nested pseudo likelihood estimation of continuous-time dynamic discrete games. Journal of Econometrics, 238(2):??, January 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002920.

Bada:2022:WMP

[BKL^{+}22] O. Bada, A. Kneip, D. Liebl, T. Mensinger, J. Gualtieri, and R. C. Sickles. A wavelet method for panel models with jump discontinuities in the parameters. Journal of Econometrics, 226(2):399–422, February 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002189.

Blasques:2022:TVP

[BKN22] Francisco Blasques, Siem Jan Koopman, and Marc Nientker. A time-varying parameter model for local explosions. Journal of Econometrics, 227(1):65–84, March 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001846.

Bartalotti:2023:IMT

[BKP23] Otávio Bartalotti, Désiré Kédagni, and Vitor Possebom. Identifying marginal treatment effects in the presence of sample selection. Journal of Econometrics, 234(2):565–584, June 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002797.

Bellemare:2022:OFP

[BKS22] Charles Bellemare, Sabine Kröger, and Kouamé Marius Sossou. Optimal frequency of portfolio evaluation in a choice experiment with ambiguity and loss aversion. Journal of Econometrics, 231(1):248–264, November 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303900.

Baltagi:2021:ETH

[BKW21] Badi H. Baltagi, Chihwa Kao, and Fa Wang. Estimating and testing high dimensional factor models with multiple structural changes. Journal of Econometrics, 220(2):349–365, February 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762030124X.

Bu:2023:UCN

[BKW23] Ruijun Bu, Jihyun Kim, and Bin Wang. Uniform and L_{p} convergences for nonparametric continuous time regressions with semiparametric applications. Journal of Econometrics, 235(2):1934– 1954, August 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623000726.

Bai:2021:DSP

[BL21] Jushan Bai and Kunpeng Li. Dynamic spatial panel data models with common shocks. Journal of Econometrics, 224(1):134–160, September 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620303961.

Barigozzi:2021:LDD

[BLL21] Matteo Barigozzi, Marco Lippi, and Matteo Luciani. Largedimensional dynamic factor models: Estimation of impulseresponse functions with I(1) cointegrated factors. Journal of Econometrics, 221(2):455–482, April 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0304407620302219.

Buchinsky:2022:EIS

[BLL22] Moshe Buchinsky, Fanghua Li, and Zhipeng Liao. Estimation and inference of semiparametric models using data from several sources. Journal of Econometrics, 226(1):80–103, January 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621000385.

Bollerslev:2024:ONR

[BLL24] Tim Bollerslev, Jia Li, and Qiyuan Li. Optimal nonparametric range-based volatility estimation. Journal of Econometrics, 238(1):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407623002646.

Bennedsen:2023:IFC

[BLSV23] Mikkel Bennedsen, Asger Lunde, Neil Shephard, and Almut E. D. Veraart. Inference and forecasting for continuous-time integer-valued trawl processes. Journal of Econometrics, 236(2): ??, October 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623001926.

Bravo:2021:RNR

[BLT21] Francesco Bravo, Degui Li, and Dag Tjøstheim. Robust nonlinear regression estimation in null recurrent time series. Journal of Econometrics, 224(2):416–438, October 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303766.

Bertanha:2020:IIE

[BM20] Marinho Bertanha and Marcelo J. Moreira. Impossible inference in econometrics: Theory and applications. Journal of Econometrics, 218(2):247–270, October 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301366.

Barbosa:2021:LIR

[BM21a] José Diogo Barbosa and Marcelo J. Moreira. Likelihood inference and the role of initial conditions for the dynamic panel data model. Journal of Econometrics, 221(1):160–179, March 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301652.

Brownlees:2021:DGT

[BM21b] Christian Brownlees and Geert Mesters. Detecting granular time series in large panels. Journal of Econometrics, 220(2):544– 561, February 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620301329.

Botosaru:2023:ITV

[BMP23] Irene Botosaru, Chris Muris, and Krishna Pendakur. Identification of time-varying transformation models with fixed effects, with an application to unobserved heterogeneity in resource shares. Journal of Econometrics, 232(2):576–597, February 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002633.

Bollerslev:2022:ZHR

[BMPQ22] Tim Bollerslev, Marcelo C. Medeiros, Andrew J. Patton, and Rogier Quaedvlieg. From zero to hero: Realized partial (co)variances. Journal of Econometrics, 231(2):348–360, December 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621002517.

Breunig:2020:IPE

[BMS20] Christoph Breunig, Enno Mammen, and Anna Simoni. Illposed estimation in high-dimensional models with instrumental variables. Journal of Econometrics, 219(1):171–200, November 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620302335.

Bertanha:2024:BBN

[BMS24] Marinho Bertanha, Andrew H. McCallum, and Nathan Seegert. Better bunching, nicer notching. Journal of Econometrics, 237(2A):??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0304407623002282.

Berger:2023:NOG

[BMW23] Tino Berger, James Morley, and Benjamin Wong. Nowcasting the output gap. Journal of Econometrics, 232(1):18–34, January 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303523.

Bai:2023:AFM

[BN23a] Jushan Bai and Serena Ng. Approximate factor models with weaker loadings. Journal of Econometrics, 235(2):1893–1916, August 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030440762300060X.

Blasques:2023:SPN

[BN23b] Francisco Blasques and Marc Nientker. Stochastic properties of nonlinear locally-nonstationary filters. Journal of Econometrics, 235(2):2082–2095, August 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623001264.

Bauwens:2020:NRC

[BO20] Luc Bauwens and Edoardo Otranto. Nonlinearities and regimes in conditional correlations with different dynamics. Journal of Econometrics, 217(2):496–522, August 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619302611.

Bognanni:2022:CLB

[Bog22] Mark Bognanni. Comment on “Large Bayesian vector autoregressions with stochastic volatility and non-conjugate priors”. Journal of Econometrics, 227(2):498–505, April 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002554. See [CCM19].

Boot:2023:JIB

[Boo23] Tom Boot. Joint inference based on Stein-type averaging estimators in the linear regression model. Journal of Econometrics, 235(2):1542–1563, August 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623000155.

Botosaru:2020:NAD

[Bot20] Irene Botosaru. Nonparametric analysis of a duration model with stochastic unobserved heterogeneity. Journal of Econometrics, 217(1):112–139, July 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761930243X.

Botosaru:2023:TVU

[Bot23] Irene Botosaru. Time-varying unobserved heterogeneity in earnings shocks. Journal of Econometrics, 235(2):1378–1393, August 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407622002020.

Boot:2020:DMS

[BP20a] Tom Boot and Andreas Pick. Does modeling a structural break improve forecast accuracy? Journal of Econometrics, 215(1):

35–59, March 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407619301824.

Bykhovskaya:2020:POT

[BP20b] Anna Bykhovskaya and Peter C. B. Phillips. Point optimal testing with roots that are functionally local to unity. Journal of Econometrics, 219(2):231–259, December 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620300981.

Bollerslev:2020:MLE

[BPQ20] Tim Bollerslev, Andrew J. Patton, and Rogier Quaedvlieg. Multivariate leverage effects and realized semicovariance GARCH models. Journal of Econometrics, 217(2):411–430, August 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619302581.

Bandi:2024:SS

[BPR24] Federico M. Bandi, Davide Pirino, and Roberto Renò. Systematic staleness. Journal of Econometrics, 238(1):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002385.

Baltagi:2021:DTH

[BPY21] Badi H. Baltagi, Alain Pirotte, and Zhenlin Yang. Diagnostic tests for homoskedasticity in spatial cross-sectional or panel models. Journal of Econometrics, 224(2):245–270, October 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303444.

Bandi:2022:T

[BR22] Federico M. Bandi and Roberto Renò. β in the tails. Journal of Econometrics, 227(1):134–150, March 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620302128.

Breunig:2021:VRC

[Bre21] Christoph Breunig. Varying random coefficient models. Journal of Econometrics, 221(2):381–408, April 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S030440762030244X.

Breitung:2021:EHP

[BS21] Jörg Breitung and Nazarii Salish. Estimation of heterogeneous panels with systematic slope variations. Journal of Econometrics, 220(2):399–415, February 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301263.

Bi:2023:DID

[BS23] Xuan Bi and Xiaotong Shen. Distribution-invariant differential privacy. Journal of Econometrics, 235(2):444–453, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762200121X.

Buchholz:2021:SED

[BSX21] Nicholas Buchholz, Matthew Shum, and Haiqing Xu. Semiparametric estimation of dynamic discrete choice models. Journal of Econometrics, 223(2):312–327, August 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303274.

Bandi:2024:BCC

[BT24] Federico M. Bandi and Andrea Tamoni. Business-cycle consumption risk and asset prices. Journal of Econometrics, 237 (2C):??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0304407623001410.

Braun:2023:ESE

[BV23] Martin Braun and Valentin Verdier. Estimation of spillover effects with matched data or longitudinal network data. Journal of Econometrics, 233(2):689–714, April 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002827.

Blasques:2024:MLE

[BvBGK24] Francisco Blasques, Janneke van Brummelen, Paolo Gorgi, and Siem Jan Koopman. Maximum likelihood estimation for non-stationary location models with mixture of normal distributions. Journal of Econometrics, 238(1):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002919.

Blasques:2022:MLE

[BvBKL22] Francisco Blasques, Janneke van Brummelen, Siem Jan Koopman, and André Lucas. Maximum likelihood estimation for score-driven models. Journal of Econometrics, 227(2):325– 346, April 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621001743.

Bao:2023:IIE

[BY23] Yong Bao and Xuewen Yu. Indirect inference estimation of dynamic panel data models. Journal of Econometrics, 235(2):1027–1053, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407622001683.

Berrisch:2024:CL

[BZ24] Jonathan Berrisch and Florian Ziel. CRPS learning. Journal of Econometrics, 237(2C):??, December 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407621002724.

Callaway:2021:BDT

[Cal21] Brantly Callaway. Bounds on distributional treatment effect parameters using panel data with an application on job displacement. Journal of Econometrics, 222(2):861–881, June 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620302839.

Caner:2023:GLM

[Can23] Mehmet Caner. Generalized linear models with structured sparsity estimators. Journal of Econometrics, 236(2):??, October 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762300194X.

Carlson:2023:RCI

[Car23] Alyssa Carlson. Relaxing conditional independence in an endogenous binary response model. Journal of Econometrics, 232(2):490–500, February 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440762100230X.

Casini:2023:TES

[Cas23] Alessandro Casini. Theory of evolutionary spectra for heteroskedasticity and autocorrelation robust inference in possibly misspecified and nonstationary models. Journal of Econometrics, 235(2):372–392, August 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622000999.

Casini:2024:FLD

[Cas24] Alessandro Casini. The fixed-b limiting distribution and the ERP of HAR tests under nonstationarity. Journal of Econometrics, 238(2):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440762300341X.

Cahan:2023:FBI

[CBN23] Ercument Cahan, Jushan Bai, and Serena Ng. Factor-based imputation of missing values and covariances in panel data of large dimensions. Journal of Econometrics, 233(1):113–131, March 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622000215.

Carriero:2022:CLB

[CCCM22] Andrea Carriero, Joshua Chan, Todd E. Clark, and Massimiliano Marcellino. Corrigendum to “Large Bayesian vector autoregressions with stochastic volatility and non-conjugate priors” [j. econometrics 212 (1) (2019) 137–154]. Journal of Econometrics, 227(2):506–512, April 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002773. See [CCM19].

Chen:2021:IAM

[CCL21] Shiyi Chen, Michael T. Chng, and Qingfu Liu. The implied arbitrage mechanism in financial markets. Journal of Econometrics, 222(1):468–483, ???? 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0304407620302013.

Cai:2023:NRI

[CCL23] Zongwu Cai, Haiqiang Chen, and Xiaosai Liao. A new robust inference for predictive quantile regression. Journal of Econometrics, 234(1):227–250, May 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762100302X.

Chen:2024:HTH

[CCL24a]  Zhao Chen, Vivian Xinyi Cheng, and Xu Liu.  Hy-

pothesis testing on high dimensional quantile regression. Journal of Econometrics, 238(1):??, January 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002415.

Chen:2024:RHT

[CCL24b] Zhao Chen, Vivian Xinyi Cheng, and Xu Liu. Reprint: Hypothesis testing on high dimensional quantile regression. Journal of Econometrics, 239(2):??, February 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623003676.

Carriero:2019:LBV

[CCM19] Andrea Carriero, Todd E. Clark, and Massimiliano Marcellino. Large Bayesian vector autoregressions with stochastic volatility and non-conjugate priors. Journal of Econometrics, 212 (1):137–154, September 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440761930079X.

See comment [Bog22] and corrigendum [CCCM22].

Carriero:2021:UTV

[CCM21] Andrea Carriero, Todd E. Clark, and Massimiliano Marcellino. Using time-varying volatility for identification in vector autoregressions: an application to endogenous uncertainty. Journal of Econometrics, 225(1):47–73, November 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001858.

Chang:2024:ABL

[CCQY24] Jinyuan Chang, Cheng Chen, Xinghao Qiao, and Qiwei Yao. An autocovariance-based learning framework for high-dimensional functional time series. Journal of Econometrics, 239(2):??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623000167.

Chen:2023:EEA

[CCT23] Jiafeng Chen, Xiaohong Chen, and Elie Tamer. Efficient estimation of average derivatives in NPIV models: Simulation comparisons of neural network estimators. Journal of Econometrics, 235(2):1848–1875, August 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623000349.

Chan:2020:IDD

[CCW20] N. H. Chan, Simon K. C. Cheung, and Samuel P. S. Wong. Inference for the degree distributions of preferential attachment networks with zero-degree nodes. Journal of Econometrics, 216(1): 220–234, May 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620300208.

Catania:2021:HMS

[CD21] Leopoldo Catania and Roberto Di Mari. Hierarchical Markovswitching models for multivariate integer-valued time-series. Journal of Econometrics, 221(1):118–137, March 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620300531.

Cunha:2022:MSE

[CEC22] Flávio Cunha, Irma Elo, and Jennifer Culhane. Maternal subjective expectations about the technology of skill formation predict investments in children one year later. Journal of Econometrics, 231(1):3–32, November 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620302700.

Chan:2020:RSS

[CES20] Joshua C. C. Chan, Eric Eisenstat, and Rodney W. Strachan. Reducing the state space dimension in a large TVP-VAR. Journal of Econometrics, 218(1):105–118, September 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620300348.

Corradi:2023:STC

[CFG23] Valentina Corradi, Jack Fosten, and Daniel Gutknecht. Outof-sample tests for conditional quantile coverage an application to growth-at-risk. Journal of Econometrics, 236(2):??, October 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002063.

Corradi:2024:PAT

[CFG24] Valentina Corradi, Jack Fosten, and Daniel Gutknecht. Predictive ability tests with possibly overlapping models. Journal of Econometrics, 241(1):??, April 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407624000629.

Cattaneo:2024:DSE

[CFLS24] Matias D. Cattaneo, Yingying Fan, Runze Li, and Rui Song. Data science in economics and finance: Introduction. Journal of Econometrics, 239(2):??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623003433.

Czellar:2022:AML

[CFR22] Veronika Czellar, David T. Frazier, and Eric Renault. Approximate maximum likelihood for complex structural models. Journal of Econometrics, 231(2):432–456, December 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002463.

Chen:2021:NFM

[CFVW21] Mingli Chen, Iván Fernández-Val, and Martin Weidner. Nonlinear factor models for network and panel data. Journal of Econometrics, 220(2):296–324, February 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301238.

Chernozhukov:2024:NPQ

[CFVW24] Victor Chernozhukov, Iván Fernández-Val, and Martin Weidner. Network and panel quantile effects via distribution regression. Journal of Econometrics, 240(2):??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303390.

Cai:2022:TCA

[CFX22] Zongwu Cai, Ying Fang, and Qiuhua Xu. Testing capital asset pricing models using functional-coefficient panel data models with cross-sectional dependence. Journal of Econometrics, 227(1):114–133, March 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620302086.

Chen:2023:CNA

[CFZ23] Elynn Y. Chen, Jianqing Fan, and Xuening Zhu. Community network auto-regression for high-dimensional time series. Journal of Econometrics, 235(2):1239–1256, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001890.

Cui:2024:MSC

[CGGK24] Xiaomeng Cui, Bulat Gafarov, Dalia Ghanem, and Todd Kuffner. On model selection criteria for climate change impact studies. Journal of Econometrics, 239(??):??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002270.

Chiou:2020:VSH

[CGI20] Hai-Tang Chiou, Meihui Guo, and Ching-Kang Ing. Variable selection for high-dimensional regression models with time series and heteroscedastic errors. Journal of Econometrics, 216(1): 118–136, May 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620300142.

Crawford:2021:SPE

[CGI21] Gregory S. Crawford, Rachel Griffith, and Alessandro Iaria. A survey of preference estimation with unobserved choice set heterogeneity. Journal of Econometrics, 222(1):4–43, ???? 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-

6895 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0304407620302463.

Cimadomo:2022:NLB

[CGL^{+}22] Jacopo Cimadomo, Domenico Giannone, Michele Lenza, Francesca Monti, and Andrej Sokol. Nowcasting with large Bayesian vector autoregressions. Journal of Econometrics, 231(2):500–519, December 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407621002499.

Chen:2023:TSI

[CGQ23] Song Xi Chen, Bin Guo, and Yumou Qiu. Testing and signal identification for two-sample high-dimensional covariances via multi-level thresholding. Journal of Econometrics, 235(2):1337– 1354, August 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407622001944.

Casarin:2024:FPD

[CGRvD24] Roberto Casarin, Stefano Grassi, Francesco Ravazzolo, and Herman K. van Dijk. A flexible predictive density combination for large financial data sets in regular and crisis periods. Journal of Econometrics, 237(2C):??, December 2024.

CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407622002093.

Chen:2022:GTG

[CGV22] Li Chen, Jiti Gao, and Farshid Vahid. Global temperatures and greenhouse gases: a common features approach. Journal of Econometrics, 230(2):240–254, October 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001159.

Chambers:2020:FDE

[Cha20] Marcus J. Chambers. Frequency domain estimation of cointegrating vectors with mixed frequency and mixed sample data. Journal of Econometrics, 217(1):140–160, July 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762030004X.

Chamberlain:2022:FPD

[Cha22] Gary Chamberlain. Feedback in panel data models. Journal of Econometrics, 226(1):4–20, January 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001937.

Chan:2023:CSV

[Cha23] Joshua C. C. Chan. Comparing stochastic volatility specifications for large Bayesian VARs. Journal of Econometrics, 235(2):1419–1446, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407622002056.

Chalak:2024:NGF

[Cha24] Karim Chalak. Nonparametric Gini–Frisch bounds. Journal of Econometrics, 238(1):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0304407623002762.

Chen:2021:ROE

[Che21] Qihui Chen. Robust and optimal estimation for partially linear instrumental variables models with partial identification.

Journal of Econometrics, 221(2):368–380, April 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407620302293.

Chen:2023:TSE

[Che23] Songnian Chen. Two-step estimation of censored quantile regression for duration models with time-varying regressors. Journal of Econometrics, 235(2):1310–1336, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001920.

Chen:2024:HFP

[Che24] Dachuan Chen. High frequency principal component analysis based on correlation matrix that is robust to jumps, microstructure noise and asynchronous observation times. Journal of Econometrics, 240(1):??, March 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407624000472.

Chen:2022:SSN

[CHK22] Cathy Yi-Hsuan Chen, Wolfgang Karl Härdle, and Yegor Klochkov. SONIC: SOcial Network analysis with Influencers and Communities. Journal of Econometrics, 228(2):177–220, June 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621000816.

Cross:2023:LSV

[CHKP23] Jamie L. Cross, Chenghan Hou, Gary Koop, and Aubrey Poon. Large stochastic volatility in mean VARs. Journal of Econometrics, 236(1):??, September 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762300163X.

Cui:2021:SEE

[CHL21] Liyuan Cui, Yongmiao Hong, and Yingxing Li. Solving Euler equations via two-stage nonparametric penalized splines. Journal of Econometrics, 222(2):1024–1056, June 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620302323.

Chen:2024:TVFb

[CHL24] Qitong Chen, Yongmiao Hong, and Haiqi Li. Time-varying forecast combination for factor-augmented regressions with smooth structural changes. Journal of Econometrics, 240(1):??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407624000393.

Chang:2024:OCM

[CHLT24] Jinyuan Chang, Qiao Hu, Cheng Liu, and Cheng Yong Tang. Optimal covariance matrix estimation for high-dimensional noise in high-frequency data. Journal of Econometrics, 239(2): ??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407622001543.

Cai:2020:HDM

[CHLZ20] T. Tony Cai, Jianchang Hu, Yingying Li, and Xinghua Zheng. High-dimensional minimum variance portfolio estimation based on high-frequency data. Journal of Econometrics, 214(2):482– 494, February 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407619301630.

Crane:2023:CLM

[CHM23] Leland D. Crane, Henry R. Hyatt, and Seth M. Murray. Cyclical labor market sorting. Journal of Econometrics, 233(2):524– 543, April 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407622000616.

Cui:2024:NEH

[CHW24] Wenhao Cui, Jie Hu, and Jiandong Wang. Nonparametric estimation for high-frequency data incorporating trading information. Journal of Econometrics, 240(1):??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407624000368.

Chen:2021:EEM

[CHY21] Xiaohong Chen, Zhuo Huang, and Yanping Yi. Efficient estimation of multivariate semi-nonparametric GARCH filtered copula models. Journal of Econometrics, 222(1):484–501, ???? 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-

6895 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0304407620302025.

Cai:2023:DRR

[CJ23] Zongwu Cai and Ted Juhl. The distribution of rolling regression estimators. Journal of Econometrics, 235(2):1447–1463, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622002068.

Cattaneo:2024:LRD

[CJM24] Matias D. Cattaneo, Michael Jansson, and Xinwei Ma. Local regression distribution estimators. Journal of Econometrics, 240 (2):??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621000427.

Chang:2023:TMD

[CJS23] Jinyuan Chang, Qing Jiang, and Xiaofeng Shao. Testing the martingale difference hypothesis in high dimension. Journal of Econometrics, 235(2):972–1000, August 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001658.

Chalak:2020:MEM

[CK20] Karim Chalak and Daniel Kim. Measurement error in multiple equations: Tobin’s q and corporate investment, saving, and debt. Journal of Econometrics, 214(2):413–432, February 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619301575.

Callaway:2023:TEI

[CK23a] Brantly Callaway and Sonia Karami. Treatment effects in interactive fixed effects models with a small number of time periods. Journal of Econometrics, 233(1):184–208, March 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762200029X.

Choi:2023:LVM

[CK23b] Sung Hoon Choi and Donggyu Kim. Large volatility matrix analysis using global and national factor models. Journal of Econometrics, 235(2):1917–1933, August 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623000635.

Creal:2024:BEC

[CK24] Drew Creal and Jaeho Kim. Bayesian estimation of cluster covariance matrices of unknown form. Journal of Econometrics, 241(1):??, April 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440762400071X.

Chang:2020:ETT

[CKK^{+}20] Yoosoon Chang, Robert K. Kaufmann, Chang Sik Kim, J. Isaac Miller, Joon Y. Park, and Sungkeun Park. Evaluating trends in time series of distributions: a spatial fingerprint of human effects on climate. Journal of Econometrics, 214(1):274–294, January 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619301216.

Choi:2024:ILR

[CKL24] Jungjun Choi, Hyukjun Kwon, and Yuan Liao. Inference for low-rank completion without sample splitting with application to treatment effect estimation. Journal of Econometrics, 240(1): ??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407624000289.

Creal:2024:ODF

[CKLZ24] Drew Creal, Siem Jan Koopman, André Lucas, and Marcin Zamojski. Observation-driven filtering of time-varying parameters using moment conditions. Journal of Econometrics, 238(2): ??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623003512.

Chesher:2023:IMT

[CKR23] Andrew Chesher, Dongwoo Kim, and Adam M. Rosen. IV methods for Tobit models. Journal of Econometrics, 235(2):1700– 1724, August 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623000258.

Chernozhukov:2021:CIM

[CKS21] Victor Chernozhukov, Hiroyuki Kasahara, and Paul Schrimpf. Causal impact of masks, policies, behavior on early Covid-19 pandemic in the U.S. Journal of Econometrics, 220(1):23–62, January 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620303468.

Chen:2024:EWS

[CKT24] Songnian Chen, Shakeeb Khan, and Xun Tang. Endogeneity in weakly separable models without monotonicity. Journal of Econometrics, 238(1):??, January 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762300283X.

Callaway:2023:PED

[CL23a] Brantly Callaway and Tong Li. Policy evaluation during a pandemic. Journal of Econometrics, 236(1):??, September 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623001483.

Chen:2023:SQR

[CL23b] Le-Yu Chen and Sokbae Lee. Sparse quantile regression. Journal of Econometrics, 235(2):2195–2217, August 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623001306.

Chen:2023:MAA

[CL23c] Yi-Ting Chen and Chu-An Liu. Model averaging for asymptotically optimal combined forecasts. Journal of Econometrics, 235(2):592–607, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407622001245.

Catania:2024:SMM

[CL24a] Leopoldo Catania and Alessandra Luati. Semiparametric modeling of multiple quantiles. Journal of Econometrics, 237(2B): ??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407622002044.

Chaudhuri:2024:FAC

[CL24b] Shomesh E. Chaudhuri and Andrew W. Lo. Financially adaptive clinical trials via option pricing analysis. Journal of Econometrics, 240(2):??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440762030364X.

Chen:2024:CIG

[CLMZ24] Xiaohong Chen, Ying Liu, Shujie Ma, and Zheng Zhang. Causal inference of general treatment effects using neural networks with a diverging number of confounders. Journal of Econometrics, 238(1):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407623002713.

Cavaliere:2023:BIH

[CLRSØ23] Giuseppe Cavaliere, Ye Lu, Anders Rahbek, and Jacob Stærk-Østergaard. Bootstrap inference for Hawkes and general point processes. Journal of Econometrics, 235(1):133–165, July 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622000574.

Choi:2023:CCB

[CLS23] In Choi, Rui Lin, and Yongcheol Shin. Canonical correlationbased model selection for the multilevel factors. Journal of Econometrics, 233(1):22–44, March 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002207.

Cai:2024:ASA

[CLWY24] Zhanrui Cai, Changcheng Li, Jiawei Wen, and Songshan Yang. Asset splitting algorithm for ultrahigh dimensional portfolio selection and its theoretical property. Journal of Econometrics, 239(2):??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407622000902.

Cheng:2024:UPI

[CLY24] Mingmian Cheng, Yuan Liao, and Xiye Yang. Uniform predictive inference for factor models with instrumental and idiosyncratic betas. Journal of Econometrics, 237(2C):??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407622002123.

Chen:2024:TVFa

[CM24] Bin Chen and Kenwin Maung. Time-varying forecast combination for high-dimensional data. Journal of Econometrics, 237(2C):??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0304407623000556.

Campos-Martins:2024:CVS

[CMH24] Susana Campos-Martins and David F. Hendry. Common volatility shocks driven by the global carbon transition. Journal of Econometrics, 239(??):??, February 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623001665.

Celhay:2024:WLM

[CMM24] Pablo Celhay, Bruce D. Meyer, and Nikolas Mittag. What leads to measurement errors? Evidence from reports of program participation in three surveys. Journal of Econometrics, 238(2): ??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030440762300297X.

Carrasco:2020:EIS

[CMPZW20] Marine Carrasco, Marcelo Moreira, Benoit Perron, and Victoria Zinde-Walsh. Editors’ introduction: Special issue in honor of Jean-Marie Dufour on identification, inference, and causality. Journal of Econometrics, 218(2):243–246, October 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620302645.

Cavaliere:2024:TBA

[CMRV24] Giuseppe Cavaliere, Thomas Mikosch, Anders Rahbek, and Frederik Vilandt. Tail behavior of ACD models and consequences for likelihood-based estimation. Journal of Econometrics, 238(2):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407623003299.

Caner:2023:SRA

[CMV23] Mehmet Caner, Marcelo Medeiros, and Gabriel F. R. Vasconcelos. Sharpe ratio analysis in high dimensions: Residual-based nodewise regression in factor models. Journal of Econometrics, 235(2):393–417, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407622000926.

Chen:2024:RRA

[CMZ24] Dachuan Chen, Per A. Mykland, and Lan Zhang. Realized regression with asynchronous and noisy high frequency and high dimensional data. Journal of Econometrics, 239(2):??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762300132X.

Cavaliere:2022:BIB

[CNPR22] Giuseppe Cavaliere, Heino Bohn Nielsen, Rasmus Søndergaard Pedersen, and Anders Rahbek. Bootstrap inference on the boundary of the parameter space, with application to conditional volatility models. Journal of Econometrics, 227(1):241– 263, March 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620302232.

Chung:2021:PTH

[CO21] EunYi Chung and Mauricio Olivares. Permutation test for heterogeneous treatment effects with a nuisance parameter. Journal of Econometrics, 225(2):148–174, December 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001561.

Christensen:2022:DBH

[COR22] Kim Christensen, Roel Oomen, and Roberto Renò. The drift burst hypothesis. Journal of Econometrics, 227(2):461–497, April 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620303912.

Casini:2021:CRL

[CP21] Alessandro Casini and Pierre Perron. Continuous record Laplace-based inference about the break date in structural change models. Journal of Econometrics, 224(1):3–21, September 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762030261X.

Centorrino:2024:NES

[CP24] Samuele Centorrino and Christopher F. Parmeter. Nonparametric estimation of stochastic frontier models with weak separability. Journal of Econometrics, 238(2):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623003573.

Carneiro:2023:PW

[CPRR23] Anabela Carneiro, Pedro Portugal, Pedro Raposo, and Paulo M. M. Rodrigues. The persistence of wages. Journal of Econometrics, 233(2):596–611, April 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002839.

Centorrino:2023:MLE

[CPU23] Samuele Centorrino and María Pérez-Urdiales. Maximum likelihood estimation of stochastic frontier models with endogeneity. Journal of Econometrics, 234(1):82–105, May 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002761.

Chan:2023:HDC

[CPZ23] Joshua C. C. Chan, Aubrey Poon, and Dan Zhu. Highdimensional conditionally Gaussian state space models with missing data. Journal of Econometrics, 236(1):??, September 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623001628.

Chaudhuri:2020:STG

[CR20] Saraswata Chaudhuri and Eric Renault. Score tests in GMM: Why use implied probabilities? Journal of Econometrics, 219(2):260–280, December 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620300993.

Chen:2024:SEL

[CR24] Jiafeng Chen and David M. Ritzwoller. Semiparametric estimation of long-term treatment effects. Journal of Econometrics, 237(2A):??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0304407623002610.

Callaway:2021:DDM

[CS21] Brantly Callaway and Pedro H. C. Sant’Anna. Difference-in-differences with multiple time periods. Journal of Econometrics, 225(2):200–230, December 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620303948.

Christensen:2022:MRI

[CS22] Jens H. E. Christensen and Mark M. Spiegel. Monetary reforms and inflation expectations in Japan: Evidence from inflationindexed bonds. Journal of Econometrics, 231(2):410–431, December 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621002542.

Card:2023:ISI

[CSV23] David Card, Ian Schmutte, and Lars Vilhuber. Introduction to the special issue: Models of linked employer-employee data: Twenty years after “High Wage Workers and High Wage Firms”. Journal of Econometrics, 233(2):333–339, April 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623000337.

Chao:2023:JEC

[CSW23] John C. Chao, Norman R. Swanson, and Tiemen Woutersen. Jackknife estimation of a cluster-sample IV regression model with many weak instruments. Journal of Econometrics, 235(2):1747–1769, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407623000271.

Chen:2022:EIH

[CSZ22] Jia Chen, Yongcheol Shin, and Chaowen Zheng. Estimation and inference in heterogeneous spatial panels with a multifactor error structure. Journal of Econometrics, 229(1):55–79, July 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001433.

Chen:2020:IAI

[CT20] Rong Chen and Ruey S. Tsay. Introduction of the annals issue: Statistical learning for dependent data — a celebration of the 85th birthday of Professor George C. Tiao. Journal of Econometrics, 216(1):1–3, May 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620300075.

Chen:2021:JEL

[CT21] Ruxin Chen and Rami V. Tabri. Jackknife empirical likelihood for inequality constraints on regular functionals. Journal of Econometrics, 221(1):68–77, March 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620300373.

Chong:2024:VVL

[CT24] Carsten H. Chong and Viktor Todorov. Volatility of volatility and leverage effect from options. Journal of Econometrics, 240 (1):??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407624000150.

Candelaria:2023:IIN

[CU23] Luis E. Candelaria and Takuya Ura. Identification and inference of network formation games with misclassified links.

Journal of Econometrics, 235(2):862–891, August 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001531.

Chen:2020:SEC

[CW20] Songnian Chen and Qian Wang. Semiparametric estimation of a censored regression model with endogeneity. Journal of Econometrics, 215(1):239–256, March 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619301848.

Chen:2023:QRC

[CW23] Songnian Chen and Qian Wang. Quantile regression with censoring and sample selection. Journal of Econometrics, 234(1): 205–226, May 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621003092.

Cheng:2024:TSD

[CWXZ24] Ming-Yen Cheng, Shouxia Wang, Lucy Xia, and Xibin Zhang. Testing specification of distribution in stochastic frontier analysis. Journal of Econometrics, 239(2):??, February 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622000677.

Chen:2022:CBT

[CXW22] Xiaohong Chen, Zhijie Xiao, and Bo Wang. Copula-based time series with filtered nonstationarity. Journal of Econometrics, 228(1):127–155, May 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620303808.

Chen:2021:AMM

[CXY21] Rong Chen, Han Xiao, and Dan Yang. Autoregressive models for matrix-valued time series. Journal of Econometrics, 222(1): 539–560, ???? 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407620302050.

Cui:2021:MLR

[CXZC21] Qiurong Cui, Yuqing Xu, Zhengjun Zhang, and Vincent Chan. Max-linear regression models with regularization. Journal of Econometrics, 222(1):579–600, ???? 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407620302074.

Choi:2022:APC

[CY22] Jungjun Choi and Xiye Yang. Asymptotic properties of correlation-based principal component analysis. Journal of Econometrics, 229(1):1–18, July 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407621002050.

Chen:2023:TSR

[CYX^{+}23] Xin Chen, Dan Yang, Yan Xu, Yin Xia, Dong Wang, and Haipeng Shen. Testing and support recovery of correlation structures for matrix-valued observations with an application to stock market data. Journal of Econometrics, 232(2):544–564, February 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621002281.

Cui:2023:SDC

[CYZ23] Yan Cui, Jun Yang, and Zhou Zhou. State-domain change point detection for nonlinear time series regression. Journal of Econometrics, 234(1):3–27, May 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407621002645.

Chen:2020:PGH

[CZ20] Songnian Chen and Hanghui Zhang. n-prediction of generalized heteroscedastic transformation regression models. Journal of Econometrics, 215(2):305–340, April 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619302003.

Dalderop:2020:NFC

[Dal20] Jeroen Dalderop. Nonparametric filtering of conditional stateprice densities. Journal of Econometrics, 214(2):295–325, February 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030440761930168X.

Dalderop:2023:SEL

[Dal23] Jeroen Dalderop. Semiparametric estimation of latent variable asset pricing models. Journal of Econometrics, 236(1):??, September 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623001598.

Delgado:2021:TCV

[DAM21] Miguel A. Delgado and Luis A. Arteaga-Molina. Testing constancy in varying coefficient models. Journal of Econometrics, 222(1):625–644, May 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0304407620302657.

Duan:2023:QML

[DBH23] Jiangtao Duan, Jushan Bai, and Xu Han. Quasi-maximum likelihood estimation of break point in high-dimensional factor models. Journal of Econometrics, 233(1):209–236, March 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622000379.

deChaisemartin:2023:TWF

[dCD23] Clément de Chaisemartin and Xavier D’Haultfoeuille. Two-way fixed effects and differences-in-differences estimators with several treatments. Journal of Econometrics, 236(2):??, October 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623001963.

DAmour:2021:OOS

[DDF^{+}21] Alexander D’Amour, Peng Ding, Avi Feller, Lihua Lei, and Jasjeet Sekhon. Overlap in observational studies with highdimensional covariates. Journal of Econometrics, 221(2):644– 654, April 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407620302694.

Delavande:2022:ANA

[DDH22] Adeline Delavande, Emilia Del Bono, and Angus Holford. Academic and non-academic investments at university: the role of expectations, preferences and constraints. Journal of Econometrics, 231(1):74–97, November 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762030275X.

DiTraglia:2023:ICE

[DGJOOSB23] Francis J. DiTraglia, Camilo García-Jimeno, Rossa O’Keeffe-O’Donovan, and Alejandro Sánchez-Becerra. Identifying causal effects in experiments with spillovers and non-compliance. Journal of Econometrics, 235(2):1589–1624, August 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623000210.

Dong:2023:HDS

[DGL23] Chaohua Dong, Jiti Gao, and Oliver Linton. High dimensional semiparametric moment restriction models. Journal of Econometrics, 232(2):320–345, February 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001883.

Dalla:2020:ATT

[DGR20] Violetta Dalla, Liudas Giraitis, and Peter M. Robinson. Asymptotic theory for time series with changing mean and variance. Journal of Econometrics, 219(2):281–313, December 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301007.

Demetrescu:2022:TEP

[DGRT22] Matei Demetrescu, Iliyan Georgiev, Paulo M. M. Rodrigues, and A. M. Robert Taylor. Testing for episodic predictability in stock returns. Journal of Econometrics, 227(1):85–113, March 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620300026.

Demetrescu:2024:EIM

[DGRT24] Matei Demetrescu, Iliyan Georgiev, Paulo M. M. Rodrigues, and A. M. Robert Taylor. Extensions to IVX methods of inference for return predictability. Journal of Econometrics, 237(2C):??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407622000586.

Daouia:2021:EEE

[DGS21] Abdelaati Daouia, Stéphane Girard, and Gilles Stupfler. ExpectHill estimation, extreme risk and heavy tails. Journal of Econometrics, 221(1):97–117, March 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620300543.

Dominicy:2020:FMH

[DHIV20] Yves Dominicy, Matias Heikkilä, Pauliina Ilmonen, and David Veredas. Flexible multivariate Hill estimators. Journal of Econometrics, 217(2):398–410, August 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619302568.

Dovonon:2020:ISO

[DHK20] Prosper Dovonon, Alastair R. Hall, and Frank Kleibergen. Inference in second-order identified models. Journal of Econometrics, 218(2):346–372, October 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301408.

Dierkes:2024:MTR

[DHPW24] Maik Dierkes, Fabian Hollstein, Marcel Prokopczuk, and Christoph Matthias Würsig. Measuring tail risk. Journal of Econometrics, 241(2):??, April 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407624001155.

DHaultfoeuille:2024:TRE

[DHS24] Xavier D’Haultfoeuille, Stefan Hoderlein, and Yuya Sasaki. Testing and relaxing the exclusion restriction in the control function approach. Journal of Econometrics, 240(2):??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621000439.

Ding:2023:SJM

[Din23] Yashuang (Dexter) Ding. A simple joint model for returns, volatility and volatility of volatility. Journal of Econometrics, 232(2):521–543, February 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407621002268.

Dai:2021:WCP

[DJK21] Min Dai, Yanwei Jia, and Steven Kou. The wisdom of the crowd and prediction markets. Journal of Econometrics, 222(1):561– 578, ???? 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407620302062.

DiAddario:2023:IAW

[DKSS23] Sabrina Di Addario, Patrick Kline, Raffaele Saggio, and Mikkel Sølvsten. It ain’t where you’re from, it’s where you’re at: Hiring origins, firm heterogeneity, and wages. Journal of Econometrics, 233(2):340–374, April 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622000641.

DInnocenzo:2024:DPC

[DL24] Enzo D’Innocenzo and Andre Lucas. Dynamic partial correlation models. Journal of Econometrics, 241(2):??, April 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407624000939.

Dostie:2023:EPI

[DLCP23] Benoit Dostie, Jiang Li, David Card, and Daniel Parent. Employer policies and the immigrant-native earnings gap. Journal of Econometrics, 233(2):544–567, April 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002293.

Ding:2024:SCJ

[DLLZ24] Yi Ding, Yingying Li, Guoli Liu, and Xinghua Zheng. Stock cojump networks. Journal of Econometrics, 239(2):??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762300057X.

Dong:2021:WSE

[DLP21] Chaohua Dong, Oliver Linton, and Bin Peng. A weighted sieve estimator for nonparametric time series models with nonstationary variables. Journal of Econometrics, 222(2):909–932, June 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303079.

Ding:2021:HDM

[DLZ21] Yi Ding, Yingying Li, and Xinghua Zheng. High dimensional minimum variance portfolio estimation under statistical factor models. Journal of Econometrics, 222(1):502–515, ???? 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-

6895 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0304407620302037.

Dominitz:2022:MRS

[DM22] Jeff Dominitz and Charles F. Manski. Minimax-regret sample design in anticipation of missing data, with application to panel data. Journal of Econometrics, 226(1):104–114, January 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620304000.

DeLuca:2022:SPB

[DMP22] Giuseppe De Luca, Jan R. Magnus, and Franco Peracchi. Sampling properties of the Bayesian posterior mean with an application to WALS estimation. Journal of Econometrics, 230(2):299– 317, October 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621001482.

Davis:2023:TSE

[DN23] Richard Davis and Serena Ng. Time series estimation of the dynamic effects of disaster-type shocks. Journal of Econometrics, 235(1):180–201, July 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622000665.

Dong:2022:EVC

[DOT22] Hao Dong, Taisuke Otsu, and Luke Taylor. Estimation of varying coefficient models with measurement error. Journal of Econometrics, 230(2):388–415, October 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001615.

Daouia:2024:EEE

[DPS24] Abdelaati Daouia, Simone A. Padoan, and Gilles Stupfler. Extreme expectile estimation for short-tailed data. Journal of Econometrics, 241(2):??, April 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407624001167.

Dufays:2020:RPC

[DR20] Arnaud Dufays and Jeroen V. K. Rombouts. Relevant parameter changes in structural break models. Journal of Econometrics, 217(1):46–78, July 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407619302441.

Demetrescu:2022:RAI

[DR22a] Matei Demetrescu and Paulo M. M. Rodrigues. Residualaugmented IVX predictive regression. Journal of Econometrics, 227(2):429–460, April 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440762030395X.

Diebold:2022:PAI

[DR22b] Francis X. Diebold and Glenn D. Rudebusch. Probability assessments of an ice-free Arctic: Comparing statistical and climate model projections. Journal of Econometrics, 231(2):520–534, December 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620304012.

Diebold:2023:WWA

[DRG^{+}23] Francis X. Diebold, Glenn D. Rudebusch, Maximilian Göbel, Philippe Goulet Coulombe, and Boyuan Zhang. When will Arctic sea ice disappear? Projections of area, extent, thickness, and volume. Journal of Econometrics, 236(2):??, October 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623001951.

Diebold:2024:RWW

[DRG^{+}24] Francis X. Diebold, Glenn D. Rudebusch, Maximilian Göbel, Philippe Goulet Coulombe, and Boyuan Zhang. Reprint of: When will Arctic sea ice disappear? Projections of area, extent, thickness, and volume. Journal of Econometrics, 239(??):??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623003615.

Demetrescu:2024:TRB

[DRT24] Matei Demetrescu, Paulo M. M. Rodrigues, and A. M. Robert Taylor. Transformed regression-based long-horizon predictability tests. Journal of Econometrics, 237(2C):??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-

6895 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0304407622001294.

Davis:2020:NVA

[DS20] Richard A. Davis and Li Song. Noncausal vector AR processes with application to economic time series. Journal of Econometrics, 216(1):246–267, May 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620300221.

Dhaene:2021:SOC

[DS21] Geert Dhaene and Yutao Sun. Second-order corrected likelihood for nonlinear panel models with fixed effects. Journal of Econometrics, 220(2):227–252, February 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301196.

DeVos:2024:CSB

[DS24] Ignace De Vos and Ovidijus Stauskas. Cross-section bootstrap for CCE regressions. Journal of Econometrics, 240(1):??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623003640.

Diebold:2024:APA

[DSZ24] Francis X. Diebold, Minchul Shin, and Boyuan Zhang. On the aggregation of probability assessments: Regularized mixtures of predictive densities for Eurozone inflation and real interest rates. Journal of Econometrics, 237(2C):??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-

6895 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0304407622001464.

Davidson:2020:FIB

[DT20] Russell Davidson and Mirza Trokić. The fast iterated bootstrap. Journal of Econometrics, 218(2):451–475, October 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301457.

Dogan:2021:BRC

[DTB21] Osman Dogan, Süleyman Taspinar, and Anil K. Bera. A Bayesian robust chi-squared test for testing simple hypotheses. Journal of Econometrics, 222(2):933–958, June 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303018.

Dellaportas:2023:SIF

[DTPP23] Petros Dellaportas, Michalis K. Titsias, Katerina Petrova, and Anastasios Plataniotis. Scalable inference for a full multivariate stochastic volatility model. Journal of Econometrics, 232(2):501–520, February 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440762100227X.

Dovonon:2022:TES

[DTW22] Prosper Dovonon, Abderrahim Taamouti, and Julian Williams. Testing the eigenvalue structure of spot and integrated covariance. Journal of Econometrics, 229(2):363–395, August 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621000579.

Delavande:2022:IJE

[DvdKWZ22] Adeline Delavande, Wilbert van der Klaauw, Joachim Winter, and Basit Zafar. Introduction to the Journal of Econometrics annals issue on “Subjective Expectations and Probabilities in Economics”. Journal of Econometrics, 231(1):1–2, November 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622000045.

Dhaene:2020:IOI

[DW20] Geert Dhaene and Jianbin Wu. Incorporating overnight and intraday returns into multivariate GARCH volatility models. Journal of Econometrics, 217(2):471–495, August 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761930260X.

Drautzburg:2023:RSI

[DW23] Thorsten Drautzburg and Jonathan H. Wright. Refining set-identification in VARs through independence. Journal of Econometrics, 235(2):1827–1847, August 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623000325.

Enache:2024:QAH

[EF24] Andreea Enache and Jean-Pierre Florens. Quantile analysis of “hazard-rate” game models. Journal of Econometrics, 238(2): ??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623002981.

Enache:2023:FEA

[EFS23] Andreea Enache, Jean-Pierre Florens, and Erwann Sbai. A functional estimation approach to the first-price auction models. Journal of Econometrics, 235(2):1564–1588, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623000192.

Eliason:2023:SCS

[EHKS23] Marcus Eliason, Lena Hensvik, Francis Kramarz, and Oskar Nordström Skans. Social connections and the sorting of workers to firms. Journal of Econometrics, 233(2):468–506, April 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407622000896.

Escanciano:2021:OLI

[EL21] Juan Carlos Escanciano and Wei Li. Optimal linear instrumental variables approximations. Journal of Econometrics, 221(1):223–246, March 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620302153.

Elliott:2020:TTP

[Ell20] Graham Elliott. Testing for a trend with persistent errors. Journal of Econometrics, 219(2):314–328, December 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301019.

Engbom:2023:FPD

[EMS23] Niklas Engbom, Christian Moser, and Jan Sauermann. Firm pay dynamics. Journal of Econometrics, 233(2):396–423, April 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622000653.

Ergemen:2023:PEL

[Erg23] Yunus Emre Ergemen. Parametric estimation of long memory in factor models. Journal of Econometrics, 235(2):1483– 1499, August 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030440762200210X.

Escanciano:2023:IIS

[Esc23] Juan Carlos Escanciano. Irregular identification of structural models with nonparametric unobserved heterogeneity. Journal of Econometrics, 234(1):106–127, May 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621003055.

Feng:2024:MPS

[Fen24] Junlong Feng. Matching points: Supplementing instruments with covariates in triangular models. Journal of Econometrics, 238(1):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407623002956.

Ferman:2021:MEF

[Fer21] Bruno Ferman. Matching estimators with few treated and many control observations. Journal of Econometrics, 225(2):295–307, December 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621001895.

Fallahgoul:2024:APN

[FFL24] Hasan Fallahgoul, Vincentius Franstianto, and Xin Lin. Asset pricing with neural networks: Significance tests. Journal of Econometrics, 238(1):??, January 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002907.

Fan:2020:PBC

[FFX20] Jianqing Fan, Yang Feng, and Lucy Xia. A projectionbased conditional dependence measure with applications to high-dimensional undirected graphical models. Journal of Econometrics, 218(1):119–139, September 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620300403.

Feng:2022:IPD

[FGP22a] Guohua Feng, Jiti Gao, and Bin Peng. An integrated panel data approach to modelling economic growth. Journal of Econometrics, 228(2):379–397, June 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621000014.

Firpo:2022:GQR

[FGP^{+}22b] Sergio Firpo, Antonio F. Galvao, Cristine Pinto, Alexandre Poirier, and Graciela Sanroman. GMM quantile regression. Journal of Econometrics, 230(2):432–452, October 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001299.

Firpo:2023:UIV

[FGP23] Sergio Firpo, Antonio F. Galvao, and Thomas Parker. Uniform inference for value functions. Journal of Econometrics, 235(2):1680–1699, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407623000246.

Feng:2022:FIB

[FH22] Guanhao Feng and Jingyu He. Factor investing: a Bayesian hierarchical approach. Journal of Econometrics, 230(1):183–200, September 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030440762100258X.

Fan:2023:VC

[FH23] Yanqin Fan and Marc Henry. Vector copulas. Journal of Econometrics, 234(1):128–150, May 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002803.

Fulop:2022:BEL

[FHLL22] Andras Fulop, Jeremy Heng, Junye Li, and Hening Liu. Bayesian estimation of long-run risk models using sequential Monte Carlo. Journal of Econometrics, 228(1):62–84, May 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621000531.

Fan:2020:REI

[FHLZ20] Yanqin Fan, Fang Han, Wei Li, and Xiao-Hua Zhou. On rank estimators in increasing dimensions. Journal of Econometrics, 214(2):379–412, February 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619301678.

Fan:2023:EIH

[FHP23] Yanqin Fan, Fang Han, and Hyeonseok Park. Estimation and inference in a high-dimensional semiparametric Gaussian copula vector autoregressive model. Journal of Econometrics, 237(1): ??, November 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623002294.

Fu:2023:STT

[FHSW23] Zhonghao Fu, Yongmiao Hong, Liangjun Su, and Xia Wang. Specification tests for time-varying coefficient models. Journal of Econometrics, 235(2):720–744, August 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001440.

Fu:2023:TSC

[FHW23] Zhonghao Fu, Yongmiao Hong, and Xia Wang. Testing for structural changes in large dimensional factor models via discrete Fourier transform. Journal of Econometrics, 233(1):302– 331, March 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407622001270.

Fisher:2022:BNL

[FJ22] Mark Fisher and Mark J. Jensen. Bayesian nonparametric learning of how skill is distributed across the mutual fund industry. Journal of Econometrics, 230(1):131–153, September 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001147.

Fan:2022:BFA

[FJS22] Jianqing Fan, Bai Jiang, and Qiang Sun. Bayesian factoradjusted sparse regression. Journal of Econometrics, 230 (1):3–19, September 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407621000828.

Frazier:2021:IIL

[FK21] David T. Frazier and Bonsoo Koo. Indirect inference for locally stationary models. Journal of Econometrics, 223(1):1–27, July 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303031.

Fan:2021:AFM

[FKL21] Jianqing Fan, Yuan Ke, and Yuan Liao. Augmented factor models with applications to validating market risk factors and forecasting bond risk premia. Journal of Econometrics, 222(1):269– 294, ???? 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407620301925.

Fan:2020:FAR

[FKW20] Jianqing Fan, Yuan Ke, and Kaizheng Wang. Factor-adjusted regularized model selection. Journal of Econometrics, 216(1): 71–85, May 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620300117.

Friedrich:2024:SBI

[FL24] Marina Friedrich and Yicong Lin. Sieve bootstrap inference for linear time-varying coefficient models. Journal of Econometrics, 239(??):??, February 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001701.

Feng:2022:HDT

[FLLM22] Long Feng, Wei Lan, Binghui Liu, and Yanyuan Ma. Highdimensional test for alpha in linear factor pricing models with sparse alternatives. Journal of Econometrics, 229(1):152– 175, July 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621001962.

Farbmacher:2022:EAN

[FLS22] Helmut Farbmacher, Leander Löw, and Martin Spindler. An explainable attention network for fraud detection in claims management. Journal of Econometrics, 228(2):244–258, June 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620302852.

Fan:2024:PQR

[FLS24] Rui Fan, Ji Hyung Lee, and Youngki Shin. Predictive quantile regression with mixed roots and increasing dimensions: the ALQR approach. Journal of Econometrics, 237(2C):??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407622002111.

Fang:2022:SMA

[FLX22] Fang Fang, Jialiang Li, and Xiaochao Xia. Semiparametric model averaging prediction for dichotomous response. Journal of Econometrics, 229(2):219–245, August 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303882.

Feng:2024:EBU

[FLZ24] Xingdong Feng, Wenyu Li, and Qianqian Zhu. Estimation and bootstrapping under spatiotemporal models with unobserved heterogeneity. Journal of Econometrics, 238(1):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002750.

Fakih:2022:SDT

[FMM^{+}22] Ali Fakih, Paul Makdissi, Walid Marrouch, Rami V. Tabri, and Myra Yazbeck. A stochastic dominance test under survey nonresponse with an application to comparing trust levels in Lebanese public institutions. Journal of Econometrics, 228(2): 342–358, June 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621002311.

Fitzenberger:2023:ETI

[FOP23] Bernd Fitzenberger, Aderonke Osikominu, and Marie Paul. The effects of training incidence and planned training duration on labor market transitions. Journal of Econometrics, 235(1):256– 279, July 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407622000690.

Forneron:2024:DIF

[For24a] Jean-Jacques Forneron. Detecting identification failure in moment condition models. Journal of Econometrics, 238(1):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623002683.

Forneron:2024:EIS

[For24b] Jean-Jacques Forneron. Estimation and inference by stochastic optimization. Journal of Econometrics, 238(2):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623003548.

Freyaldenhoven:2022:FML

[Fre22] Simon Freyaldenhoven. Factor models with local factors — determining the number of relevant factors. Journal of Econometrics, 229(1):80–102, July 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407621001275.

Fiorentini:2021:NTA

[FS21] Gabriele Fiorentini and Enrique Sentana. New testing approaches for mean-variance predictability. Journal of Econometrics, 222(1):516–538, ???? 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0304407620302049.

Fan:2023:WQS

[FS23a] Yanqin Fan and Xuetao Shi. Wald, QLR, and score tests when parameters are subject to linear inequality constraints. Journal of Econometrics, 235(2):2005–2026, August 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623000787.

Fiorentini:2023:DMN

[FS23b] Gabriele Fiorentini and Enrique Sentana. Discrete mixtures of normals pseudo maximum likelihood estimators of structural vector autoregressions. Journal of Econometrics, 235(2):643– 665, August 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407622001269.

Fan:2023:PII

[FST23] Yanqin Fan, Xuetao Shi, and Jing Tao. Partial identification and inference in moment models with incomplete data. Journal of Econometrics, 235(2):418–443, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001002.

Friedrich:2020:AWB

[FSU20] Marina Friedrich, Stephan Smeekes, and Jean-Pierre Urbain. Autoregressive wild bootstrap inference for nonparametric trends. Journal of Econometrics, 214(1):81–109, January 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619301095.

Funovits:2024:IEP

[Fun24] Bernd Funovits. Identifiability and estimation of possibly noninvertible SVARMA models: the normalised canonical WHF parametrisation. Journal of Econometrics, 241(2):??, April 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762400112X.

Fusejima:2024:IMV

[Fus24] Koki Fusejima. Identification of multi-valued treatment effects with unobserved heterogeneity. Journal of Econometrics, 238(1):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407623002798.

Fernandez-Val:2024:NSS

[FVvVV24] Ivan Fernández-Val, Aico van Vuuren, and Francis Vella. Nonseparable sample selection models with censored selection rules. Journal of Econometrics, 240(2):??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621000567.

Freeman:2023:LPR

[FW23] Hugo Freeman and Martin Weidner. Linear panel regressions with two-way unobserved heterogeneity. Journal of Econometrics, 237(1):??, November 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002142.

Fan:2024:TVM

[FWYZ24] Qingliang Fan, Ruike Wu, Yanrong Yang, and Wei Zhong. Time-varying minimum variance portfolio. Journal of Econometrics, 239(2):??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407622001646.

Francq:2020:VHS

[FZ20] Christian Francq and Jean-Michel Zakoïan. Virtual Historical Simulation for estimating the conditional VaR of large portfolios. Journal of Econometrics, 217(2):356–380, August 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619302544.

Francq:2022:TEM

[FZ22] Christian Francq and Jean-Michel Zakoïan. Testing the existence of moments for GARCH processes. Journal of Econometrics, 227(1):47–64, March 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620302268.

Gao:2020:SIJ

[GAL20] Jiti Gao, Heather Anderson, and Tong Li. Special issue of the Journal of Econometrics on “Econometric Estimation and Testing: Essays in Honour of Maxwell King”. Journal of Econometrics, 219(2):201–203, December 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620300968.

Gallant:2022:NBS

[Gal22] A. Ronald Gallant. Nonparametric Bayes subject to overidentified moment conditions. Journal of Econometrics, 228(1):27– 38, May 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621000555.

Gallant:2023:VCM

[Gal23] A. Ronald Gallant. Variance-covariance from a Metropolis chain on a curved, singular manifold. Journal of Econometrics, 235(2):843–861, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407622001506.

Gao:2020:NII

[Gao20] Wayne Yuan Gao. Nonparametric identification in index models of link formation. Journal of Econometrics, 215(2):399– 413, April 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407619302039.

Goodman-Bacon:2021:DDV

[GB21a] Andrew Goodman-Bacon. Difference-in-differences with variation in treatment timing. Journal of Econometrics, 225 (2):254–277, December 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407621001445.

Guethmundsson:2021:DGL

[GB21b] Gumundur Stefán Gumundsson and Christian Brownlees. Detecting groups in large vector autoregressions. Journal of Econometrics, 225(1):2–26, November 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762100124X.

Guo:2023:ICL

[GCT23] Xiao Guo, Yu Chen, and Cheng Yong Tang. Information criteria for latent factor models: a study on factor pervasiveness and adaptivity. Journal of Econometrics, 233(1):237–250, March 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622000707.

Graham:2022:SEE

[GdXP22] Bryan S. Graham and Cristine Campos de Xavier Pinto. Semiparametrically efficient estimation of the average linear regression function. Journal of Econometrics, 226(1):115–138, January 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621001925.

Gelman:2023:WSE

[Gel23]  Andrew Gelman. What is a standard error?  Jour-

nal of Econometrics, 237(1):??, November 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002324. See discussion [Pow23].

Rivas:2020:TDC

[GG20a] María Dolores Gadea Rivas and Jesús Gonzalo. Trends in distributional characteristics: Existence of global warming. Journal of Econometrics, 214(1):153–174, January 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619301125.

Gimenes:2020:NII

[GG20b] Nathalie Gimenes and Emmanuel Guerre. Nonparametric identification of an interdependent value model with buyer covariates from first-price auction bids. Journal of Econometrics, 219(1):1–18, November 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440762030110X.

Gimenes:2022:QRM

[GG22] Nathalie Gimenes and Emmanuel Guerre. Quantile regression methods for first-price auctions. Journal of Econometrics, 226(2):224–247, February 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407621001524.

Gaglianone:2022:IDI

[GGIS22] Wagner Piazza Gaglianone, Raffaella Giacomini, João Victor Issler, and Vasiliki Skreta. Incentive-driven inattention. Journal of Econometrics, 231(1):188–212, November 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620302736.

Giannerini:2024:VBT

[GGR24] Simone Giannerini, Greta Goracci, and Anders Rahbek. The validity of bootstrap testing for threshold autoregression. Journal of Econometrics, 239(??):??, February 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623000040.

Galvao:2020:UAN

[GGV20] Antonio F. Galvao, Jiaying Gu, and Stanislav Volgushev. On the unbiased asymptotic normality of quantile regression with fixed effects. Journal of Econometrics, 218(1):178–215, September 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620300415.

Graham:2022:IAI

[GH22] Bryan Graham and Keisuke Hirano. Introduction to the annals issue in honor of Gary Chamberlain. Journal of Econometrics, 226(1):1–3, January 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407621002128.

Gribisch:2023:MRC

[GH23] Bastian Gribisch and Jan Patrick Hartkopf. Modeling realized covariance measures with heterogeneous liquidity: a generalized matrix-variate Wishart state-space model. Journal of Econometrics, 235(1):43–64, July 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407622000392.

Ghanem:2024:CAB

[GHKOB24] Dalia Ghanem, Sarojini Hirshleifer, Désiré Kédagni, and Karen Ortiz-Becerra. Correcting attrition bias using changes-in-changes. Journal of Econometrics, 241(2):??, April 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407624000836.

Goncalves:2021:IRA

[GHKP21] Sílvia Gonçalves, Ana María Herrera, Lutz Kilian, and Elena Pesavento. Impulse response analysis for structural dynamic models with nonlinear regressors. Journal of Econometrics, 225(1):107–130, November 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001810.

Gallant:2022:CEU

[GHLL22] A. Ronald Gallant, Han Hong, Michael P. Leung, and Jessie Li. Constrained estimation using penalization and MCMC. Journal of Econometrics, 228(1):85–106, May 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762100052X.

Ghysels:2020:TLS

[GHM20] Eric Ghysels, Jonathan B. Hill, and Kaiji Motegi. Testing a large set of zero restrictions in regression models, with an application to mixed frequency Granger causality. Journal of Econometrics, 218(2):633–654, October 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301524.

Gourieroux:2023:TVM

[GJ23] C. Gourieroux and J. Jasiak. Time varying Markov process with partially observed aggregate data: an application to coronavirus. Journal of Econometrics, 232(1):35–51, January 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303791.

Gourieroux:2020:SBE

[GJM20] C. Gourieroux, J. Jasiak, and A. Monfort. Stationary bubble equilibria in rational expectation models. Journal of Econometrics, 218(2):714–735, October 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762030155X.

Gorgi:2024:BOD

[GK24] P. Gorgi and S. J. Koopman. Beta observation-driven models with exogenous regressors: a joint analysis of realized correlation and leverage effects. Journal of Econometrics, 237(2B):??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407621002165.

Giraitis:2021:TVI

[GKM21] Liudas Giraitis, George Kapetanios, and Massimiliano Marcellino. Time-varying instrumental variable estimation. Journal of Econometrics, 224(2):394–415, October 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303754.

Guggenberger:2023:TKP

[GKM23] Patrik Guggenberger, Frank Kleibergen, and Sophocles Mavroeidis. A test for Kronecker Product Structure covariance matrix. Journal of Econometrics, 233(1):88–112, March 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622000203.

Giacomini:2022:RBI

[GKR22] Raffaella Giacomini, Toru Kitagawa, and Matthew Read. Robust Bayesian inference in proxy SVARs. Journal of Econometrics, 228(1):107–126, May 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407621000518.

Gu:2021:AAP

[GKX21] Shihao Gu, Bryan Kelly, and Dacheng Xiu. Autoencoder asset pricing models. Journal of Econometrics, 222(1):429–450, ???? 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-

6895 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0304407620301998.

Gungor:2020:SST

[GL20] Sermin Gungor and Richard Luger. Small-sample tests for stock return predictability with possibly non-stationary regressors and GARCH-type effects. Journal of Econometrics, 218(2):750– 770, October 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620301573.

Ge:2023:NIL

[GLL23] Shuyi Ge, Shaoran Li, and Oliver Linton. News-implied linkages and local dependency in the equity market. Journal of Econometrics, 235(2):779–815, August 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001488.

Guo:2023:SIL

[GLLZ23] Xu Guo, Runze Li, Jingyuan Liu, and Mudong Zeng. Statistical inference for linear mediation models with high-dimensional mediators and application to studying stock reaction to COVID-19 pandemic. Journal of Econometrics, 235(1):166–179, July 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622000598.

Guo:2024:RSI

[GLLZ24] Xu Guo, Runze Li, Jingyuan Liu, and Mudong Zeng. Reprint: Statistical inference for linear mediation models with highdimensional mediators and application to studying stock reaction to COVID-19 pandemic. Journal of Econometrics, 239(2): ??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623003664.

Giraitis:2024:RIC

[GLP24] Liudas Giraitis, Yufei Li, and Peter C. B. Phillips. Robust inference on correlation under general heterogeneity. Journal of Econometrics, 240(1):??, March 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762400037X.

Gao:2023:BRM

[GLPY23] Jiti Gao, Fei Liu, Bin Peng, and Yayi Yan. Binary response models for heterogeneous panel data with interactive fixed effects. Journal of Econometrics, 235(2):1654–1679, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623000234.

Gold:2020:IHD

[GLT20] David Gold, Johannes Lederer, and Jing Tao. Inference for high-dimensional instrumental variables regression. Journal of Econometrics, 217(1):79–111, July 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619302386.

Grewenig:2022:ISE

[GLWW22] Elisabeth Grewenig, Philipp Lergetporer, Katharina Werner, and Ludger Woessmann. Incentives, search engines, and the elicitation of subjective beliefs: Evidence from representative online survey experiments. Journal of Econometrics, 231(1):304–326, November 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620302797.

Gao:2023:LDD

[GLX23] Wayne Yuan Gao, Ming Li, and Sheng Xu. Logical differencing in dyadic network formation models with nontransferable utilities. Journal of Econometrics, 235(1):302–324, July 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622000884.

Gospodinov:2021:GAM

[GM21] Nikolay Gospodinov and Esfandiar Maasoumi. Generalized aggregation of misspecified models: With an application to asset pricing. Journal of Econometrics, 222(1):451–467, ???? 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-

6895 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0304407620302001.

Giustinelli:2022:TCR

[GMM22] Pamela Giustinelli, Charles F. Manski, and Francesca Molinari. Tail and center rounding of probabilistic expectations in the health and retirement study. Journal of Econometrics, 231(1):265–281, November 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620302761.

Graham:2024:KDE

[GNP24] Bryan S. Graham, Fengshi Niu, and James L. Powell. Kernel density estimation for undirected dyadic data. Journal of Econometrics, 240(2):??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407622001610.

Goff:2024:VMA

[Gof24] Leonard Goff. A vector monotonicity assumption for multiple instruments. Journal of Econometrics, 241(1):??, April 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407624000812.

Goncalves:2020:BFM

[GP20] Sílvia Gonçalves and Benoit Perron. Bootstrapping factor models with cross sectional dependence. Journal of Econometrics, 218(2):476–495, October 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620301469.

Guay:2023:SVM

[GP23] Alain Guay and Florian Pelgrin. Structural VAR models in the frequency domain. Journal of Econometrics, 236(1):??, September 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623001604.

Gao:2024:TVM

[GPWY24] Jiti Gao, Bin Peng, Wei Biao Wu, and Yayi Yan. Time-varying multivariate causal processes. Journal of Econometrics, 240(1): ??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407624000174.

Gu:2023:PIN

[GR23] Jiaying Gu and Thomas M. Russell. Partial identification in nonseparable binary response models with endogenous regressors. Journal of Econometrics, 235(2):528–562, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001166.

Guay:2021:EEF

[GS21] François Guay and Gustavo Schwenkler. Efficient estimation and filtering for multivariate jump-diffusions. Journal of Econometrics, 223(1):251–275, July 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303511.

Gong:2022:MCL

[GSS22] Yifan Gong, Ralph Stinebrickner, and Todd Stinebrickner. Marriage, children, and labor supply: Beliefs and outcomes. Journal of Econometrics, 231(1):148–164, November 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620302803.

Gardner:2022:WSL

[GSV22] Ben Gardner, Chiara Scotti, and Clara Vega. Words speak as loudly as actions: Central bank communication and the response of equity prices to macroeconomic announcements. Journal of Econometrics, 231(2):387–409, December 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002530.

Gualdani:2021:EMN

[Gua21a] Cristina Gualdani. An econometric model of network formation with an application to board interlocks between firms. Journal of Econometrics, 224(2):345–370, October 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303729.

Guay:2021:ISV

[Gua21b] Alain Guay. Identification of structural vector autoregressions through higher unconditional moments. Journal of Econometrics, 225(1):27–46, November 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303651.

Gunsilius:2023:CIM

[Gun23] Florian F. Gunsilius. A condition for the identification of multivariate models with binary instruments. Journal of Econometrics, 235(1):220–238, July 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622000872.

Gupta:2023:ECF

[Gup23] Abhimanyu Gupta. Efficient closed-form estimation of large spatial autoregressions. Journal of Econometrics, 232(1):148– 167, January 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621001597.

Gao:2020:HPD

[GXZ20] Jiti Gao, Kai Xia, and Huanjun Zhu. Heterogeneous panel data models with cross-sectional dependence. Journal of Econometrics, 219(2):329–353, December 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301020.

Grundl:2023:RIF

[GZ23] Serafin Grundl and Yu Zhu. Robust inference in first-price auctions: Overbidding as an identifying restriction. Journal of Econometrics, 235(2):484–506, August 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001221.

Galbraith:2020:SRE

[GZW20] John W. Galbraith and Victoria Zinde-Walsh. Simple and reliable estimators of coefficients of interest in a model with high-dimensional confounding effects. Journal of Econometrics, 218(2):609–632, October 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620301512.

Han:2021:INM

[Han21] Sukjin Han. Identification in nonparametric models for dynamic treatment effects. Journal of Econometrics, 225(2):132–147, December 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620303717.

Han:2024:PIP

[HBB24] Kevin Han, Guillaume Basse, and Iavor Bojinov. Population interference in panel experiments. Journal of Econometrics, 238(1):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407623002816.

Hansen:2022:HSP

[HD22] Peter Reinhard Hansen and Elena-Ivona Dumitrescu. How should parameter estimation be tailored to the objective? Journal of Econometrics, 230(2):535–558, October 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001822.

Hsieh:2024:NRS

[HHK^{+}24] Chih-Sheng Hsieh, Yu-Chin Hsu, Stanley I. M. Ko, Jaromír Kovárík, and Trevon D. Logan. Non-representative sampled networks: Estimation of network structural properties by weighting. Journal of Econometrics, 240(1):??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407624000356.

Hahn:2020:STM

[HHL20] Jinyong Hahn, Jerry Hausman, and Josh Lustig. Specification test on mixed logit models. Journal of Econometrics, 219(1):19– 37, November 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620301184.

Han:2022:RPS

[HHLS22] Dongxiao Han, Jian Huang, Yuanyuan Lin, and Guohao Shen. Robust post-selection inference of high-dimensional mean regression with heavy-tailed asymmetric or heteroskedastic errors. Journal of Econometrics, 230(2):416–431, October 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001639.

Hafner:2022:ISM

[HHM22] Christian M. Hafner, Helmut Herwartz, and Simone Maxand. Identification of structural multivariate GARCH models. Journal of Econometrics, 227(1):212–227, March 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620302098.

Heiss:2022:NER

[HHO22] Florian Heiss, Stephan Hetzenecker, and Maximilian Osterhaus. Nonparametric estimation of the random coefficients model: an elastic net approach. Journal of Econometrics, 229(2):299–321, August 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621000178.

Harvey:2024:MCT

[HHPT24] Andrew Harvey, Stan Hurn, Dario Palumbo, and Stephen Thiele. Modelling circular time series. Journal of Econometrics, 239(??):??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407623001446.

Hu:2020:EPF

[HHS20] Yingyao Hu, Guofang Huang, and Yuya Sasaki. Estimating production functions with robustness against errors in the proxy variables. Journal of Econometrics, 215(2):375–398, April 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619302027.

Heiss:2022:DHS

[HHvR^{+}22] Florian Heiss, Michael Hurd, Maarten van Rooij, Tobias Rossmann, and Joachim Winter. Dynamics and heterogeneity of subjective stock market expectations. Journal of Econometrics, 231(1):213–231, November 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002232.

Harvey:2020:MTS

[HI20] Andrew Harvey and Ryoko Ito. Modeling time series when some observations are zero. Journal of Econometrics, 214(1):33–45, January 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030440761930106X.

Higgins:2023:IMD

[HJ23] Ayden Higgins and Koen Jochmans. Identification of mixtures of dynamic discrete choices. Journal of Econometrics, 237(1): ??, November 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623001562.

Herbst:2024:BLP

[HJ24] Edward P. Herbst and Benjamin K. Johannsen. Bias in local projections. Journal of Econometrics, 240(1):??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407624000010.

He:2023:MPT

[HJG23] Yi He, Sombut Jaidee, and Jiti Gao. Most powerful test against a sequence of high dimensional local alternatives. Journal of Econometrics, 234(1):151–177, May 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621003079.

Huang:2024:BRP

[HJL^{+}24] Dashan Huang, Fuwei Jiang, Kunpeng Li, Guoshi Tong, and Guofu Zhou. Are bond returns predictable with real-time macro data? Journal of Econometrics, 237(2C):??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-

6895 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0304407623001161.

Huang:2023:BAM

[HJLP23] Haitao Huang, Lei Jiang, Xuan Leng, and Liang Peng. Bootstrap analysis of mutual fund performance. Journal of Econometrics, 235(1):239–255, July 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622000951.

Huang:2021:NPM

[HJPS21] Wenxin Huang, Sainan Jin, Peter C. B. Phillips, and Liangjun Su. Nonstationary panel models with latent group structures and cross-section dependence. Journal of Econometrics, 221(1):198–222, March 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620302165.

Hou:2024:EVS

[HJW24] Li Hou, Baisuo Jin, and Yuehua Wu. Estimation and variable selection for high-dimensional spatial dynamic panel data models. Journal of Econometrics, 238(2):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623003214.

Heiler:2021:VIT

[HK21] Phillip Heiler and Ekaterina Kazak. Valid inference for treatment effect parameters under irregular identification and many extreme propensity scores. Journal of Econometrics, 222(2):1083–1108, June 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620303377.

Hwang:2022:DCR

[HKL22] Jungbin Hwang, Byunghoon Kang, and Seojeong Lee. A doubly corrected robust variance estimator for linear GMM. Journal of Econometrics, 229(2):276–298, August 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621000166.

Huitfeldt:2023:ILM

[HKNW23] Ingrid Huitfeldt, Andreas R. Kostøl, Jan Nimczik, and Andrea Weber. Internal labor markets: a worker flow approach. Journal of Econometrics, 233(2):661–688, April 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762200063X.

Huber:2023:NPU

[HKO^{+}23] Florian Huber, Gary Koop, Luca Onorante, Michael Pfarrhofer, and Josef Schreiner. Nowcasting in a pandemic using non-parametric mixed frequency VARs. Journal of Econometrics, 232(1):52–69, January 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620303936.

Han:2020:LEP

[HKR20] Hyojin Han, Stanislav Khrapov, and Eric Renault. The leverage effect puzzle revisited: Identification in discrete time. Journal of Econometrics, 217(2):230–258, August 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619302490.

He:2024:LMT

[HKST24] Changli He, Jian Kang, Annastiina Silvennoinen, and Timo Teräsvirta. Long monthly temperature series and the Vector Seasonal Shifting Mean and Covariance Autoregressive model. Journal of Econometrics, 239(??):??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002105.

Harris:2020:LSE

[HKT20] David Harris, Hsein Kew, and A. M. Robert Taylor. Level shift estimation in the presence of non-stationary volatility with an application to the unit root testing problem. Journal of Econometrics, 219(2):354–388, December 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301032.

He:2023:OWT

[HKTY23] Yong He, Xinbing Kong, Lorenzo Trapani, and Long Yu. Oneway or two-way factor model for matrix sequences? Journal of Econometrics, 235(2):1981–2004, August 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762300074X.

Hong:2021:OSE

[HKW21] Han Hong, Michael Keane, and Clifford Winston. Overview: Structural econometrics honoring Daniel McFadden. Journal of Econometrics, 222(1):1–3, ???? 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0304407620302451.

Hallin:2020:SEM

[HL20a] Marc Hallin and Davide La Vecchia. A simple R-estimation method for semiparametric duration models. Journal of Econometrics, 218(2):736–749, October 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0304407620301561.

Hong:2020:NEI

[HL20b] Seok Young Hong and Oliver Linton. Nonparametric estimation of infinite order regression and its application to the risk-return tradeoff. Journal of Econometrics, 219(2):389–424, December 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301044.

Hounyo:2023:EVC

[HL23] Ulrich Hounyo and Kajal Lahiri. Estimating the variance of a combined forecast: Bootstrap-based approach. Journal of Econometrics, 232(2):445–468, February 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002244.

Hong:2021:BEU

[HLL21] Han Hong, Huiyu Li, and Jessie Li. BLP estimation using Laplace transformation and overlapping simulation draws. Journal of Econometrics, 222(1):56–72, ???? 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407620302487.

Hou:2021:RLF

[HLLO21] Lei Hou, Kunpeng Li, Qi Li, and Min Ouyang. Revisiting the location of FDI in China: a panel data approach with heterogeneous shocks. Journal of Econometrics, 221(2):483– 509, April 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407620302426.

Ho:2023:HGV

[HLM23] Paul Ho, Thomas A. Lubik, and Christian Matthes. How to go viral: a COVID-19 model with endogenously time-varying parameters. Journal of Econometrics, 232(1):70–86, January 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621000105.

Hong:2024:KST

[HLM^{+}24] Yongmiao Hong, Oliver Linton, Brendan McCabe, Jiajing Sun, and Shouyang Wang. Kolmogorov–Smirnov type testing for structural breaks: a new adjusted-range based selfnormalization approach. Journal of Econometrics, 238(2):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623003196.

Hou:2024:PQR

[HLPZ24] Yanxi Hou, Xuan Leng, Liang Peng, and Yinggang Zhou. Panel quantile regression for extreme risk. Journal of Econometrics, 240(1):??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407624000204.

Horvath:2020:SMC

[HLRW20] Lajos Horváth, Zhenya Liu, Gregory Rice, and Shixuan Wang. Sequential monitoring for changes from stationarity to mild nonstationarity. Journal of Econometrics, 215(1):209–238, March 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619301964.

Hortacsu:2021:EFU

[HLS21] Ali Hortaçsu, Jiarui Liu, and Timothy Schwieg. Estimating the fraction of unreported infections in epidemics with a known epicenter: an application to COVID-19. Journal of Econometrics, 220(1):106–129, January 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762030302X.

Hafner:2020:EMC

[HLT20] Christian M. Hafner, Oliver B. Linton, and Haihan Tang. Estimation of a multiplicative correlation structure in the large dimensional case. Journal of Econometrics, 217(2):431–470, August 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407619302593.

Harvey:2021:STS

[HLT21] David I. Harvey, Stephen J. Leybourne, and A. M. Robert Taylor. Simple tests for stock return predictability with good size and power properties. Journal of Econometrics, 224(1):198–214, September 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621000270.

Hao:2024:DMM

[HLWZ24] Siteng Hao, Shu-Chin Lin, Jane-Ling Wang, and Qixian Zhong. Dynamic modeling for multivariate functional and longitudinal data. Journal of Econometrics, 239(2):??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002890.

Heiler:2021:SCR

[HM21] Phillip Heiler and Jana Mareckova. Shrinkage for categorical regressors. Journal of Econometrics, 223(1):161–189, July 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303407.

Higgins:2023:SEN

[HM23] Ayden Higgins and Federico Martellosio. Shrinkage estimation of network spillovers with factor structured errors. Journal of Econometrics, 233(1):66–87, March 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621003080.

Higbee:2024:CGS

[HM24] Joshua D. Higbee and James B. McDonald. A comparison of the GB2 and skewed generalized log-t distributions with an application in finance. Journal of Econometrics, 240(2):??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621000154.

Henry:2024:RMR

[HMM24] Marc Henry, Romuald Méango, and Ismaël Mourifié. Role models and revealed gender-specific costs of STEM in an extended Roy model of major choice. Journal of Econometrics, 238(2): ??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623002877.

Horowitz:2021:UPL

[HN21] Joel L. Horowitz and Lars Nesheim. Using penalized likelihood to select parameters in a random coefficients multinomial logit model. Journal of Econometrics, 222(1):44–55, ???? 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-

6895 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0304407620300427.

Hong:2022:AAF

[HNZ22] Zhiwu Hong, Linlin Niu, and Chen Zhang. Affine arbitragefree yield net models with application to the euro debt crisis. Journal of Econometrics, 230(1):201–220, September 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002591.

Ho:2023:GRB

[Ho23] Paul Ho. Global robust Bayesian analysis in large models. Journal of Econometrics, 235(2):608–642, August 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001257.

Horowitz:2021:BDB

[Hor21] Joel L. Horowitz. Bounding the difference between true and nominal rejection probabilities in tests of hypotheses about instrumental variables models. Journal of Econometrics, 222(2):1057–1082, June 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620303353.

Hoshino:2022:SIE

[Hos22] Tadao Hoshino. Sieve IV estimation of cross-sectional interaction models with nonparametric endogenous effect. Journal of Econometrics, 229(2):263–275, August 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620304036.

Harvey:2024:SDM

[HP24] Andrew Harvey and Dario Palumbo. Score-driven models for realized volatility. Journal of Econometrics, 237(2B):??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407623001422.

Hillebrand:2020:EMC

[HPP20] Eric Hillebrand, Felix Pretis, and Tommaso Proietti. Econometric models of climate change: Introduction by the guest editors. Journal of Econometrics, 214(1):1–5, January 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619301046.

Hetland:2024:DCE

[HPR24] Simon Hetland, Rasmus Søndergaard Pedersen, and Anders Rahbek. Dynamic conditional eigenvalue GARCH. Journal of Econometrics, 237(2B):??, December 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407621002141.

He:2023:SQR

[HPTZ23] Xuming He, Xiaoou Pan, Kean Ming Tan, and Wen-Xin Zhou. Smoothed quantile regression with large-scale inference. Journal of Econometrics, 232(2):367–388, February 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001950.

Hall:2021:EES

[HR21] George Hall and John Rust. Estimation of endogenously sampled time series: the case of commodity price speculation in the steel market. Journal of Econometrics, 222(1):219–243, ???? 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-

6895 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0304407620302554.

Hollstein:2020:VRB

[HS20] Fabian Hollstein and Chardin Wese Simen. Variance risk: a bird’s eye view. Journal of Econometrics, 215(2):517–535, April 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619302076.

Hansen:2021:MUP

[HS21a] Lars Peter Hansen and Thomas J. Sargent. Macroeconomic uncertainty prices when beliefs are tenuous. Journal of Econometrics, 223(1):222–250, July 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303419.

Hidalgo:2021:ISL

[HS21b] Javier Hidalgo and Marcia Schafgans. Inference without smoothing for large panels with cross-sectional and temporal dependence. Journal of Econometrics, 223(1):125–160, July 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303481.

Hajivassiliou:2024:SII

[HS24] Vassilis Hajivassiliou and Frédérique Savignac. Simultaneously incomplete and incoherent (SII) dynamic LDV models: With an application to financing constraints and firms’ decision to innovate. Journal of Econometrics, 238(1):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002622.

Hansen:2020:TPU

[HSHS20] Lars Peter Hansen, Bálint Szőke, Lloyd S. Han, and Thomas J. Sargent. Twisted probabilities, uncertainty, and prices. Journal of Econometrics, 216(1):151–174, May 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620300166.

Hong:2023:PGE

[HSJ23] Shengjie Hong, Liangjun Su, and Tao Jiang. Profile GMM estimation of panel data models with interactive fixed effects. Journal of Econometrics, 235(2):927–948, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001592.

Hsieh:2022:IED

[HSS22a] Yu-Wei Hsieh, Xiaoxia Shi, and Matthew Shum. Inference on estimators defined by mathematical programming. Journal of Econometrics, 226(2):248–268, February 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762100172X.

Hu:2022:INM

[HSS22b] Yingyao Hu, Susanne Schennach, and Ji-Liang Shiu. Identification of nonparametric monotonic regression models with continuous nonclassical measurement errors. Journal of Econometrics, 226(2):269–294, February 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001305.

Hsu:2024:TIC

[HSW24] Yu-Chin Hsu, Ji-Liang Shiu, and Yuanyuan Wan. Testing identification conditions of LATE in fuzzy regression discontinuity designs. Journal of Econometrics, 241(1):??, April 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407624000848.

Holt:2020:GHT

[HT20] Matthew T. Holt and Timo Teräsvirta. Global hemispheric temperatures and co-shifting: a vector shifting-mean autoregressive analysis. Journal of Econometrics, 214(1):198–215, January 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619301149.

Hubner:2023:IUD

[Hub23] Stefan Hubner. Identification of unobserved distribution factors and preferences in the collective household model. Journal of Econometrics, 234(1):301–326, May 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622000173.

Hounyo:2020:ILD

[HV20] Ulrich Hounyo and Rasmus T. Varneskov. Inference for local distributions at high sampling frequencies: a bootstrap approach. Journal of Econometrics, 215(1):1–34, March 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619301836.

Hwang:2023:FSC

[HV23] Jungbin Hwang and Gonzalo Valdés. Finite-sample corrected inference for two-step GMM in time series. Journal of Econometrics, 234(1):327–352, May 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622000069.

Hirano:2022:ACV

[HW22] Keisuke Hirano and Jonathan H. Wright. Analyzing crossvalidation for forecasting with structural instability. Journal of Econometrics, 226(1):139–154, January 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620304024.

Hafner:2024:DCS

[HW24] Christian M. Hafner and Linqi Wang. A dynamic conditional score model for the log correlation matrix. Journal of Econometrics, 237(2B):??, December 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407621002153.

Hwang:2021:STC

[Hwa21] Jungbin Hwang. Simple and trustworthy cluster-robust GMM inference. Journal of Econometrics, 222(2):993–1023, June 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303043.

Huang:2020:TMN

[HWZW20] Danyang Huang, Feifei Wang, Xuening Zhu, and Hansheng Wang. Two-mode network autoregressive model for largescale networks. Journal of Econometrics, 216(1):203–219, May 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620300191.

Hu:2022:IEG

[HY22] Yingyao Hu and Jiaxiong Yao. Illuminating economic growth. Journal of Econometrics, 228(2):359–378, June 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001767.

Hoshino:2023:TEM

[HY23] Tadao Hoshino and Takahide Yanagi. Treatment effect models with strategic interaction in treatment decisions. Journal of Econometrics, 236(2):??, October 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002117.

Han:2024:CAI

[HY24] Sukjin Han and Shenshen Yang. A computational approach to identification of treatment effects for policy evaluation. Journal of Econometrics, 240(1):??, March 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407624000265.

Inoue:2020:UVI

[IK20] Atsushi Inoue and Lutz Kilian. The uniform validity of impulse response inference in autoregressions. Journal of Econometrics, 215(2):450–472, April 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619302040.

Iskhakov:2021:ETS

[IK21] Fedor Iskhakov and Michael Keane. Effects of taxes and safety net pensions on life-cycle labor supply, savings and human capital: the case of Australia. Journal of Econometrics, 223(2):401– 432, August 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620303262.

Inoue:2022:JBI

[IK22] Atsushi Inoue and Lutz Kilian. Joint Bayesian inference about impulse responses in VAR models. Journal of Econometrics, 231(2):457–476, December 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002475.

Inoue:2022:ESI

[IKP22] Atsushi Inoue, Lutz Kilian, and Andrew Patton. Editorial for special issue in honor of Francis X. Diebold. Journal of Econometrics, 231(2):327–328, December 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762100244X.

Ikefuji:2020:EUC

[ILMM20] Masako Ikefuji, Roger J. A. Laeven, Jan R. Magnus, and Chris Muris. Expected utility and catastrophic risk in a stochastic economy-climate model. Journal of Econometrics, 214(1):110– 129, January 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407619301101.

Ishihara:2020:IET

[Ish20] Takuya Ishihara. Identification and estimation of timevarying nonseparable panel data models without stayers. Journal of Econometrics, 215(1):184–208, March 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619301873.

Carrion-i-Silvestre:2021:STS

[iSK21] Josep Lluís Carrion i Silvestre and Dukpa Kim. Statistical tests of a simple energy balance equation in a synthetic model of cotrending and cointegration. Journal of Econometrics, 224(1):22–38, September 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303559.

Jarjour:2020:DCA

[JC20] Riad Jarjour and Kung-Sik Chan. Dynamic conditional angular correlation. Journal of Econometrics, 216(1):137–150, May 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620300154.

Jeong:2020:SDM

[JfL20] Hanbat Jeong and Lung fei Lee. Spatial dynamic models with intertemporal optimization: Specification and estimation. Journal of Econometrics, 218(1):82–104, September 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762030035X.

Jin:2024:MME

[JKL24] Jiashun Jin, Zheng Tracy Ke, and Shengming Luo. Mixed membership estimation for social networks. Journal of Econometrics, 239(2):??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407622002081.

Juodis:2021:RPC

[JKW21] Arturas Juodis, Hande Karabiyik, and Joakim Westerlund. On the robustness of the pooled CCE estimator. Journal of Econometrics, 220(2):325–348, February 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301834.

Jenkins:2021:BWA

[JLMM21] Mark Jenkins, Paul Liu, Rosa L. Matzkin, and Daniel L. McFadden. The browser war — analysis of Markov Perfect Equilibrium in markets with dynamic demand effects. Journal of Econometrics, 222(1):244–260, ???? 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407620302566.

Jiang:2020:TSH

[JLP20] Bibo Jiang, Ye Lu, and Joon Y. Park. Testing for stationarity at high frequency. Journal of Econometrics, 215(2):341–374, April 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619302015.

Joao:2024:DCM

[JLSS24] Igor Custodio João, André Lucas, Julia Schaumburg, and Bernd Schwaab. Dynamic clustering of multivariate panel data. Journal of Econometrics, 237(2B):??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407622000689.

Jiang:2020:NSI

[JLZ20] Feiyu Jiang, Dong Li, and Ke Zhu. Non-standard inference for augmented double autoregressive models with null volatility coefficients. Journal of Econometrics, 215(1):165–183, March 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619301885.

Jiang:2021:AIS

[JLZ21] Feiyu Jiang, Dong Li, and Ke Zhu. Adaptive inference for a semiparametric generalized autoregressive conditional heteroskedasticity model. Journal of Econometrics, 224(2):306– 329, October 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620303699.

Jentsch:2021:VAI

[JM21] Carsten Jentsch and Marco Meyer. On the validity of Akaike’s identity for random fields. Journal of Econometrics, 222(1):676– 687, May 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407620302347.

Jin:2021:FMR

[JMS21] Sainan Jin, Ke Miao, and Liangjun Su. On factor models with random missing: EM estimation, inference, and cross validation. Journal of Econometrics, 222(1):745–777, May 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-

6895 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0304407620302815.

Jin:2022:IMP

[JMY22] Xin Jin, John M. Maheu, and Qiao Yang. Infinite Markov pooling of predictive distributions. Journal of Econometrics, 228(2): 302–321, June 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621002578.

Jochmans:2023:PEE

[Joc23] Koen Jochmans. Peer effects and endogenous social interactions. Journal of Econometrics, 235(2):1203–1214, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001853.

Jun:2020:CPC

[JP20] Sung Jae Jun and Joris Pinkse. Counterfactual prediction in complete information games: Point prediction under partial identification. Journal of Econometrics, 216(2):394–429, June 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619302258.

Jun:2024:ITA

[JP24] Sung Jae Jun and Joris Pinkse. An information-theoretic approach to partially identified auction models. Journal of Econometrics, 238(2):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407623002828.

Jiao:2024:TCD

[JPS24] Xiyu Jiao, Felix Pretis, and Moritz Schwarz. Testing for coefficient distortion due to outliers with an application to the economic impacts of climate change. Journal of Econometrics, 239(??):??, February 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002634.

Jiang:2023:RAE

[JPTZ23] Liang Jiang, Peter C. B. Phillips, Yubo Tao, and Yichong Zhang. Regression-adjusted estimation of quantile treatment effects under covariate-adaptive randomizations. Journal of Econometrics, 234(2):758–776, June 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001865.

Juodis:2022:IPF

[JS22] Arturas Juodis and Vasilis Sarafidis. An incidental parameters free inference approach for panels with common shocks. Journal of Econometrics, 229(1):19–54, July 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001135.

Janys:2024:MHA

[JS24] Lena Janys and Bettina Siflinger. Mental health and abortions among young women: time-varying unobserved heterogeneity, health behaviors, and risky decisions. Journal of Econometrics, 238(1):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407623002968.

Julliard:2023:SCL

[JSY23] Christian Julliard, Ran Shi, and Kathy Yuan. The spread of COVID-19 in London: Network effects and optimal lockdowns. Journal of Econometrics, 235(2):2125–2154, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623001288.

Jiang:2021:REL

[JYGH21] Bin Jiang, Yanrong Yang, Jiti Gao, and Cheng Hsiao. Recursive estimation in large panel data models: Theory and practice. Journal of Econometrics, 224(2):439–465, October 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303870.

Jun:2022:TRA

[JZ22] Sung Jae Jun and Federico Zincenko. Testing for risk aversion in first-price sealed-bid auctions. Journal of Econometrics, 226(2):295–320, February 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407621001469.

Jiang:2023:TSA

[JZS23] Feiyu Jiang, Zifeng Zhao, and Xiaofeng Shao. Time series analysis of COVID-19 infection curve: a change-point perspective. Journal of Econometrics, 232(1):1–17, January 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620302633.

Kamat:2024:IEP

[Kam24] Vishal Kamat. Identifying the effects of a program offer with an application to Head Start. Journal of Econometrics, 240(1): ??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407624000253.

Kasy:2022:WWW

[Kas22] Maximilian Kasy. Who wins, who loses? Identification of conditional causal effects, and the welfare impact of changing wages. Journal of Econometrics, 226(1):155–170, January 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621000452.

Koo:2020:HDP

[KASY20] Bonsoo Koo, Heather M. Anderson, Myung Hwan Seo, and Wenying Yao. High-dimensional predictive regression in the presence of cointegration. Journal of Econometrics, 219(2): 456–477, December 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620301068.

Kedagni:2023:ITE

[Kéd23] Désiré Kédagni. Identifying treatment effects in the presence of confounded types. Journal of Econometrics, 234(2):479– 511, June 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621001512.

Klein:2020:MRP

[KHK20] Nadja Klein, Helmut Herwartz, and Thomas Kneib. Modelling regional patterns of inefficiency: a Bayesian approach to geoadditive panel stochastic frontier analysis with an application to cereal production in England and Wales. Journal of Econometrics, 214(2):513–539, February 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619301587.

Kim:2023:PIC

[Kim23] Dongwoo Kim. Partially identifying competing risks models: an application to the war on cancer. Journal of Econometrics, 234(2):536–564, June 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407621001913.

Kitagawa:2021:IRP

[Kit21] Toru Kitagawa. The identification region of the potential outcome distributions under instrument independence. Journal of Econometrics, 225(2):231–253, December 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621000968.

Kitazawa:2022:TMC

[Kit22] Yoshitsugu Kitazawa. Transformations and moment conditions for dynamic fixed effects logit models. Journal of Econometrics, 229(2):350–362, August 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407621000464.

Kiviet:2020:TII

[Kiv20] Jan F. Kiviet. Testing the impossible: Identifying exclusion restrictions. Journal of Econometrics, 218(2):294–316, October 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762030138X.

Kociecki:2023:SGI

[KK23] Andrzej Kociecki and Marcin Kolasa. A solution to the global identification problem in DSGE models. Journal of Econometrics, 236(2):??, October 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407623001938.

Keane:2021:OIS

[KKIS21] Michael Keane, Dennis Kristensen, Fedor Iskhakov, and Bertel Schjerning. Overview: Implementation of structural dynamic models: Methodology and applications. Journal of Econometrics, 223(2):277–279, August 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001019.

Keane:2021:ECC

[KKKN21] Michael Keane, Jonathan Ketcham, Nicolai Kuminoff, and Timothy Neal. Evaluating consumers’ choices of Medicare Part D plans: a study in behavioral welfare economics. Journal of Econometrics, 222(1):107–140, ???? 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407620302517.

Konstantinidi:2023:STR

[KKS23] Antri Konstantinidi, Andros Kourtellos, and Yiguo Sun. Social threshold regression. Journal of Econometrics, 235(2):2057– 2081, August 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623001045.

Khalaf:2021:DPM

[KKSV21] Lynda Khalaf, Maral Kichian, Charles J. Saunders, and Marcel Voia. Dynamic panels with MIDAS covariates: Nonlinearity, estimation and fit. Journal of Econometrics, 220(2):589–605, February 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620301342.

Kleibergen:2021:ESC

[Kle21] Frank Kleibergen. Efficient size correct subset inference in homoskedastic linear instrumental variables regression. Journal of Econometrics, 221(1):78–96, March 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762030052X.

Kline:2024:CVB

[Kli24] Brendan Kline. Classical p-values and the Bayesian posterior probability that the hypothesis is approximately true. Journal of Econometrics, 240(1):??, March 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762400023X.

Koo:2021:ENM

[KLL21] Bonsoo Koo, Davide La Vecchia, and Oliver Linton. Estimation of a nonparametric model for bond prices from cross-section and time series information. Journal of Econometrics, 220(2):562– 588, February 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620301330.

Kolokolov:2020:SIP

[KLP20] Aleksey Kolokolov, Giulia Livieri, and Davide Pirino. Statistical inferences for price staleness. Journal of Econometrics, 218(1):32–81, September 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620300270.

Kueck:2023:EIT

[KLSW23] Jannis Kueck, Ye Luo, Martin Spindler, and Zigan Wang. Estimation and inference of treatment effects with L_{2}-boosting in high-dimensional settings. Journal of Econometrics, 234(2): 714–731, June 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407622000471.

Kim:2020:VRF

[KM20] Jihyun Kim and Nour Meddahi. Volatility regressions with fat tails. Journal of Econometrics, 218(2):690–713, October 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301548.

Kristensen:2021:SDD

[KMMS21] Dennis Kristensen, Patrick K. Mogensen, Jong Myun Moon, and Bertel Schjerning. Solving dynamic discrete choice models using smoothing and sieve methods. Journal of Econometrics, 223(2):328–360, August 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620303298.

Kojevnikov:2021:LTN

[KMS21] Denis Kojevnikov, Vadim Marmer, and Kyungchul Song. Limit theorems for network dependent random variables. Journal of Econometrics, 222(2):882–908, June 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620302402.

Kostakis:2024:TSL

[KMS24] Alexandros Kostakis, Tassos Magdalinos, and Michalis P. Stamatogiannis. Taking stock of long-horizon predictability tests: Are factor returns predictable? Journal of Econometrics, 237(2C):??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0304407623000052.

Keane:2021:CPC

[KN21] Michael Keane and Timothy Neal. Consumer panic in the COVID-19 pandemic. Journal of Econometrics, 220(1):86–105, January 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620302840.

Keane:2023:ISI

[KN23] Michael Keane and Timothy Neal. Instrument strength in IV estimation and inference: a guide to theory and practice. Journal of Econometrics, 235(2):1625–1653, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623000222.

Khan:2024:UIN

[KN24] Shakeeb Khan and Denis Nekipelov. On uniform inference in nonlinear models with endogeneity. Journal of Econometrics, 240(2):??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407622000409.

Knox:2022:ASR

[Kno22] Thomas A. Knox. Approximation of sign-regular kernels. Journal of Econometrics, 226(1):171–191, January 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001949.

Kim:2020:IRC

[KOEP20] Dukpa Kim, Tatsushi Oka, Francisco Estrada, and Pierre Perron. Inference related to common breaks in a multivariate system with joined segmented trends with applications to global and hemispheric temperatures. Journal of Econometrics, 214(1):130–152, January 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619301113.

Koh:2023:SOI

[Koh23] Paul S. Koh. Stable outcomes and information in games: an empirical framework. Journal of Econometrics, 237(1):??, November 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002154.

Kitagawa:2020:PDN

[KOPV20] Toru Kitagawa, José Luis Montiel Olea, Jonathan Payne, and Amilcar Velez. Posterior distribution of nondifferentiable functions. Journal of Econometrics, 217(1):161–175, July 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620300014.

Korolev:2021:IES

[Kor21] Ivan Korolev. Identification and estimation of the SEIRD epidemic model for COVID-19. Journal of Econometrics, 220 (1):63–85, January 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620302621.

Kheifets:2023:FML

[KP23] Igor L. Kheifets and Peter C. B. Phillips. Fully modified least squares cointegrating parameter estimation in multicointegrated systems. Journal of Econometrics, 232(2):300–319, February 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762100186X.

Kapetanios:2021:DUP

[KPR21] G. Kapetanios, M. H. Pesaran, and S. Reese. Detection of units with pervasive effects in large panel data models. Journal of Econometrics, 221(2):510–541, April 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407620302141.

Ke:2024:RIP

[KPS24] Shuyao Ke, Peter C. B. Phillips, and Liangjun Su. Robust inference of panel data models with interactive fixed effects under long memory: a frequency domain approach. Journal of Econometrics, 241(2):??, April 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407624001076.

Khan:2023:IDB

[KPT23a] S. Khan, M. Ponomareva, and E. Tamer. Identification of dynamic binary response models. Journal of Econometrics, 237(1): ??, November 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623002312.

Krampe:2023:SIS

[KPT23b] J. Krampe, E. Paparoditis, and C. Trenkler. Structural inference in sparse high-dimensional vector autoregressions. Journal of Econometrics, 234(1):276–300, May 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622000057.

Kock:2023:TRD

[KPV23] Anders Bredahl Kock, David Preinerstorfer, and Bezirgen Veliyev. Treatment recommendation with distributional targets. Journal of Econometrics, 234(2):624–646, June 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001518.

Kuersteiner:2023:EPE

[KPZ23] Guido M. Kuersteiner, Ingmar R. Prucha, and Ying Zeng. Efficient peer effects estimators with group effects. Journal of Econometrics, 235(2):2155–2194, August 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762300129X.

Kosar:2022:UMA

[KRvdK22] Gizem Kosar, Tyler Ransom, and Wilbert van der Klaauw. Understanding migration aversion using elicited counterfactual choice probabilities. Journal of Econometrics, 231(1):123–147, November 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621000415.

Karmakar:2022:SIT

[KRW22] Sayar Karmakar, Stefan Richter, and Wei Biao Wu. Simultaneous inference for time-varying models. Journal of Econometrics, 227(2):408–428, April 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407621000725.

Khalaf:2020:MCT

[KS20] Lynda Khalaf and Charles J. Saunders. Monte Carlo twostage indirect inference (2SIF) for autoregressive panels. Journal of Econometrics, 218(2):419–434, October 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301433.

Kojevnikov:2023:EIL

[KS23] Denis Kojevnikov and Kyungchul Song. Econometric inference on a large Bayesian game with heterogeneous beliefs.

Journal of Econometrics, 237(1):??, November 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762300218X.

Kojevnikov:2024:SIR

[KS24] Denis Kojevnikov and Kyungchul Song. Some impossibility results for inference with cluster dependence with large clusters. Journal of Econometrics, 237(2A):??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407623002403.

Kapetanios:2021:EIM

[KSS21] George Kapetanios, Laura Serlenga, and Yongcheol Shin. Estimation and inference for multi-dimensional heterogeneous panel datasets with hierarchical multi-factor error structure. Journal of Econometrics, 220(2):504–531, February 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301305.

Kalouptsidi:2021:LIR

[KSSR21] Myrto Kalouptsidi, Paul T. Scott, and Eduardo Souza-Rodrigues. Linear IV regression estimators for structural dynamic discrete choice models. Journal of Econometrics, 222(1): 778–804, May 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407620302578.

Krasnokutskaya:2022:EUI

[KST22] Elena Krasnokutskaya, Kyungchul Song, and Xun Tang. Estimating unobserved individual heterogeneity using pairwise comparisons. Journal of Econometrics, 226(2):477–497, February 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621000142.

Kolokotrones:2024:NWO

[KSW24] Thomas Kolokotrones, James H. Stock, and Christopher D. Walker. Is Newey–West optimal among first-order kernels? Journal of Econometrics, 240(2):??, March 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623000301.

Khismatullina:2023:NCE

[KV23] Marina Khismatullina and Michael Vogt. Nonparametric comparison of epidemic time trends: the case of COVID-19. Journal of Econometrics, 232(1):87–108, January 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762100155X.

Kole:2023:MSS

[KvD23] Erik Kole and Dick van Dijk. Moments, shocks and spillovers in Markov-switching VAR models. Journal of Econometrics, 236(2):??, October 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407623001902.

Kitagawa:2023:WSG

[KW23] Toru Kitagawa and Guanyi Wang. Who should get vaccinated? individualized allocation of vaccines over SIR network. Journal of Econometrics, 232(1):109–131, January 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002219.

Khalil:2022:TSO

[KY22] Umair Khalil and Nese Yildiz. A test of the selection on observables assumption using a discontinuously distributed covariate. Journal of Econometrics, 226(2):423–450, February 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002451.

Komunjer:2020:LRT

[KZ20] Ivana Komunjer and Yinchu Zhu. Likelihood ratio testing in linear state space models: an application to dynamic stochastic general equilibrium models. Journal of Econometrics, 218(2):561–586, October 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620301494.

Kaplan:2021:FPB

[KZ21] David M. Kaplan and Longhao Zhuo. Frequentist properties of Bayesian inequality tests. Journal of Econometrics, 221(1):312– 336, March 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620302359.

King:2020:HTB

[KZA20] Maxwell L. King, Xibin Zhang, and Muhammad Akram. Hypothesis testing based on a vector of statistics. Journal of Econometrics, 219(2):425–455, December 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301056.

Lange:2024:BFS

[Lan24] Rutger-Jan Lange. Bellman filtering and smoothing for statespace models. Journal of Econometrics, 238(2):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623003482.

Liu:2020:TFM

[LC20] Xialu Liu and Rong Chen. Threshold factor models for highdimensional time series. Journal of Econometrics, 216(1):53– 70, May 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620300105.

Li:2020:EEH

[LCL20] Kunpeng Li, Guowei Cui, and Lina Lu. Efficient estimation of heterogeneous coefficients in panel data models with common shocks. Journal of Econometrics, 216(2):327–353, June 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619302222.

Leng:2023:MDL

[LCW23] Xuan Leng, Heng Chen, and Wendun Wang. Multi-dimensional latent group structures with heterogeneous distributions. Journal of Econometrics, 233(1):1–21, March 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002177.

Lewbel:2023:IDR

[LCZ23] Arthur Lewbel, Jin Young Choi, and Zhuzhu Zhou. Overidentified Doubly Robust identification and estimation. Journal of Econometrics, 235(1):25–42, July 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0304407622000434.

Lewbel:2022:KFL

[Lew22] Arthur Lewbel. Kotlarski with a factor loading. Journal of Econometrics, 229(1):176–179, July 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001603.

Lin:2024:BCM

[LH24] Zhongjian Lin and Yingyao Hu. Binary choice with misclassification and social interactions, with an application to peer effects in attitude. Journal of Econometrics, 238(1):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002671.

Liao:2024:IRI

[Lia24] Moyu Liao. Identification of a rational inattention discrete choice model. Journal of Econometrics, 240(1):??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407624000162.

Li:2020:LTF

[LKLP20] Mengheng Li, Siem Jan Koopman, Rutger Lit, and Desislava Petrova. Long-term forecasting of El Niño events via dynamic factor simulations. Journal of Econometrics, 214(1):46–66, January 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407619301071.

Li:2020:UNI

[LL20] Jia Li and Zhipeng Liao. Uniform nonparametric inference for time series. Journal of Econometrics, 219(1):38–51, November 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301354.

Li:2021:WWC

[LL21] Shaoran Li and Oliver Linton. When will the Covid-19 pandemic peak? Journal of Econometrics, 220(1):130–157, January 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303055.

Lewbel:2022:ISM

[LL22] Arthur Lewbel and Xirong Lin. Identification of semiparametric model coefficients, with an application to collective households. Journal of Econometrics, 226(2):205–223, February 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001834.

Lee:2023:PPP

[LL23] Kwangmin Lee and Jaeyong Lee. Post-processed posteriors for sparse covariances. Journal of Econometrics, 236(1):??, September 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623001914.

Liu:2022:PFL

[LLCW22] Yanghui Liu, Yehua Li, Raymond J. Carroll, and Naisyin Wang. Predictive functional linear models with diverging number of semiparametric single-index interactions. Journal of Econometrics, 230(2):221–239, October 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001020.

Lee:2021:SHF

[LLSS21] Sokbae Lee, Yuan Liao, Myung Hwan Seo, and Youngki Shin. Sparse HP filter: Finding kinks in the COVID-19 contact rate. Journal of Econometrics, 220(1):158–180, January 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303365.

Li:2020:DMN

[LLV20] Z. Merrick Li, Roger J. A. Laeven, and Michel H. Vellekoop. Dependent microstructure noise and integrated volatility estimation from high-frequency data. Journal of Econometrics, 215(2):536–558, April 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407619302106.

Lee:2023:TTS

[LLW23] Kyungho Lee, Oliver Linton, and Yoon-Jae Whang. Testing for time stochastic dominance. Journal of Econometrics, 235(2):352–371, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407622000963.

Liu:2022:PBW

[LLYZ22] Xiaobin Liu, Yong Li, Jun Yu, and Tao Zeng. Posteriorbased Wald-type statistics for hypothesis testing. Journal of Econometrics, 230(1):83–113, September 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002608.

Li:2022:VVE

[LLZ22] Yingying Li, Guangying Liu, and Zhiyuan Zhang. Volatility of volatility: Estimation and tests based on noisy high frequency data with jumps. Journal of Econometrics, 229(2):422– 451, August 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621000701.

Lee:2024:LRI

[LM24] Adam Lee and Geert Mesters. Locally robust inference for non-Gaussian linear simultaneous equations models. Journal of Econometrics, 240(1):??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407623003639.

Loaiza-Maya:2024:HUL

[LMNZ24] Rubén Loaiza-Maya, Didier Nibbering, and Dan Zhu. Hybrid unadjusted Langevin methods for high-dimensional latent variable models. Journal of Econometrics, 241(2):??, April 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407624000873.

Liu:2024:ALF

[LMR24] Lin Liu, Rajarshi Mukherjee, and James M. Robins. Assumptionlean falsification tests of rate double-robustness of doublemachine-learning estimators. Journal of Econometrics, 240(2): ??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623002166.

Liu:2021:PFC

[LMS21] Laura Liu, Hyungsik Roger Moon, and Frank Schorfheide. Panel forecasts of country-level Covid-19 infections. Journal of Econometrics, 220(1):2–22, January 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762030347X.

LaVecchia:2023:HOC

[LMS23] Davide La Vecchia, Alban Moor, and Olivier Scaillet. A higher-order correct fast moving-average bootstrap for dependent data. Journal of Econometrics, 235(1):65–81, July 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622000422.

Loaiza-Maya:2022:FAV

[LMSND22] Rubén Loaiza-Maya, Michael Stanley Smith, David J. Nott, and Peter J. Danaher. Fast and accurate variational inference for models with many latent variables. Journal of Econometrics, 230(2):339–362, October 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001330.

Lachowska:2023:DFE

[LMSW23] Marta Lachowska, Alexandre Mas, Raffaele Saggio, and Stephen A. Woodbury. Do firm effects drift? Evidence from Washington administrative data. Journal of Econometrics, 233(2):375–395, April 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407622000604.

Lopes:2022:PIP

[LMT22] Hedibert F. Lopes, Robert E. McCulloch, and Ruey S. Tsay. Parsimony inducing priors for large scale state-space models. Journal of Econometrics, 230(1):39–61, September 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002621.

Lutkepohl:2020:IPI

[LMY20] Helmut Lütkepohl, George Milunovich, and Minxian Yang. Inference in partially identified heteroskedastic simultaneous equations models. Journal of Econometrics, 218(2):317–345, October 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301391.

Li:2024:PRN

[LNP24] Yifan Li, Ingmar Nolte, and Manh Cuong Pham. Parametric risk-neutral density estimation via finite lognormal-Weibull mixtures. Journal of Econometrics, 241(2):??, April 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407624000940.

Loh:2023:NIE

[Loh23] Isaac Loh. Nonparametric identification and estimation with discrete instruments and regressors. Journal of Econometrics, 235(2):1257–1279, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407622001907.

Lumsdaine:2023:EPG

[LOW23] Robin L. Lumsdaine, Ryo Okui, and Wendun Wang. Estimation of panel group structure models with structural breaks in group memberships and coefficients. Journal of Econometrics, 233(1): 45–65, March 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407622000033.

Lettau:2020:ELA

[LP20a] Martin Lettau and Markus Pelger. Estimating latent assetpricing factors. Journal of Econometrics, 218(1):1–31, September 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620300051.

Lieberman:2020:HSL

[LP20b] Offer Lieberman and Peter C. B. Phillips. Hybrid stochastic local unit roots. Journal of Econometrics, 215(1):257–285, March 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619301897.

Lieberman:2022:UTA

[LP22] Offer Lieberman and Peter C. B. Phillips. Understanding temporal aggregation effects on kurtosis in financial indices. Journal of Econometrics, 227(1):25–46, March 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762030258X.

Lamarche:2023:WBI

[LP23a] Carlos Lamarche and Thomas Parker. Wild bootstrap inference for penalized quantile regression for longitudinal data. Journal of Econometrics, 235(2):1799–1826, August 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623000313.

Lee:2023:NIE

[LP23b] Ji Hyung Lee and Byoung G. Park. Nonparametric identification and estimation of the extended Roy model. Journal of Econometrics, 235(2):1087–1113, August 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001774.

Liu:2023:RIS

[LP23c] Yanbo Liu and Peter C. B. Phillips. Robust inference with stochastic local unit root regressors in predictive regressions. Journal of Econometrics, 235(2):563–591, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001233.

Li:2020:KBI

[LPG20] Degui Li, Peter C. B. Phillips, and Jiti Gao. Kernelbased inference in time-varying coefficient cointegrating regression. Journal of Econometrics, 215(2):607–632, April 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619302210.

Lamy:2023:DIS

[LPV23] Laurent Lamy, Manasa Patnam, and Michael Visser. Distinguishing incentive from selection effects in auction-determined contracts. Journal of Econometrics, 235(2):1172–1202, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001828.

Lui:2024:RTE

[LPY24] Yiu Lim Lui, Peter C. B. Phillips, and Jun Yu. Robust testing for explosive behavior with strongly dependent errors. Journal of Econometrics, 238(2):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623003421.

Lu:2021:SEO

[LQ21] Junwen Lu and Zhongjun Qu. Sieve estimation of optionimplied state price density. Journal of Econometrics, 224 (1):88–112, September 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407621000737.

Lee:2020:AIP

[LR20] Jungyoon Lee and Peter M. Robinson. Adaptive inference on pure spatial models. Journal of Econometrics, 216(2):375– 393, June 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407619302234.

Laurent:2020:VEJ

[LS20a] Sébastien Laurent and Shuping Shi. Volatility estimation and jump detection for drift-diffusion processes. Journal of Econometrics, 217(2):259–290, August 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619302507.

Lu:2020:DIT

[LS20b] Xun Lu and Liangjun Su. Determining individual or time effects in panel data models. Journal of Econometrics, 215(1): 60–83, March 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407619301861.

Lavetti:2023:GDS

[LS23a] Kurt Lavetti and Ian M. Schmutte. Gender differences in sorting on wages and risk. Journal of Econometrics, 233(2):507– 523, April 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030440762200183X.

Lu:2023:UIL

[LS23b] Xun Lu and Liangjun Su. Uniform inference in linear panel data models with two-dimensional heterogeneity. Journal of Econometrics, 235(2):694–719, August 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001312.

Li:2024:IHE

[LS24] Tong Li and Yuya Sasaki. Identification of heterogeneous elasticities in gross-output production functions. Journal of Econometrics, 238(2):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407623003536.

Lee:2022:LPR

[LSG22] Ji Hyung Lee, Zhentao Shi, and Zhan Gao. On LASSO for predictive regression. Journal of Econometrics, 229(2):322–349, August 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030440762100049X.

Liu:2024:GKP

[LSK24] Jingyuan Liu, Ao Sun, and Yuan Ke. A generalized knockoff procedure for FDR control in structural change detection. Journal of Econometrics, 239(2):??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001567.

Lu:2023:SNE

[LST23] Zhentong Lu, Xiaoxia Shi, and Jing Tao. Semi-nonparametric estimation of random coefficients logit model for aggregate demand. Journal of Econometrics, 235(2):2245–2265, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623001458.

Lee:2024:TPF

[LSTW24] Ji Hyung Lee, Yuya Sasaki, Alexis Akira Toda, and Yulong Wang. Tuning parameter-free nonparametric density estimation from tabulated summary data. Journal of Econometrics, 238(1): ??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623002841.

Linton:2023:TSD

[LSW23] Oliver Linton, Myung Hwan Seo, and Yoon-Jae Whang. Testing stochastic dominance with many conditioning variables. Journal of Econometrics, 235(2):507–527, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001191.

Li:2024:CIT

[LSZ24] Xingyu Li, Yan Shen, and Qiankun Zhou. Confidence intervals of treatment effects in panel data models with interactive fixed effects. Journal of Econometrics, 240(1):??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407624000307.

Liu:2020:IEP

[LSZZ20] Ruiqi Liu, Zuofeng Shang, Yonghui Zhang, and Qiankun Zhou. Identification and estimation in panel models with overspecified number of groups. Journal of Econometrics, 215(2):574– 590, April 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407619302118.

Lin:2020:RIS

[LT20] Yingqian Lin and Yundong Tu. Robust inference for spurious regressions and cointegrations involving processes moderately deviated from a unit root. Journal of Econometrics, 219(1):52– 65, November 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620301585.

Lok:2021:IBT

[LT21] Thomas M. Lok and Rami V. Tabri. An improved bootstrap test for restricted stochastic dominance. Journal of Econometrics, 224(2):371–393, October 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620303730.

Lin:2020:EDN

[LTY20] Yingqian Lin, Yundong Tu, and Qiwei Yao. Estimation for double-nonlinear cointegration. Journal of Econometrics, 216(1):175–191, May 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620300178.

Lin:2021:UHS

[LTY21] Zhongjian Lin, Xun Tang, and Ning Neil Yu. Uncovering heterogeneous social effects in binary choices. Journal of Econometrics, 222(2):959–973, June 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303080.

Li:2023:MLE

[LTYZ23] Dong Li, Yuxin Tao, Yaxing Yang, and Rongmao Zhang. Maximum likelihood estimation for α-stable double autoregressive models. Journal of Econometrics, 236(1):??, September 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623001653.

Linton:2021:ESI

[LTZ21] Oliver Linton, Viktor Todorov, and Zhengjun Zhang. Editorial for the special issue on financial econometrics in the age of the digital economy. Journal of Econometrics, 222(1):265– 268, ???? 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407620301913.

Lu:2022:EMC

[Lu22] Zhentong Lu. Estimating multinomial choice models with unobserved choice sets. Journal of Econometrics, 226(2):368–398, February 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621001755.

Luo:2020:UHA

[Luo20] Yao Luo. Unobserved heterogeneity in auctions under restricted stochastic dominance. Journal of Econometrics, 216(2):354– 374, June 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407619302246.

Linton:2020:CCD

[LW20] Oliver Linton and Jianbin Wu. A coupled component DCS-EGARCH model for intraday and overnight volatility. Journal of Econometrics, 217(1):176–201, July 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620300038.

Lee:2023:TRN

[LW23a] Yoonseok Lee and Yulong Wang. Threshold regression with nonparametric sample splitting. Journal of Econometrics, 235(2):816–842, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440762200149X.

Li:2023:PPE

[LW23b] Hengxin Li and Ruodu Wang. PELVE: Probability equivalent level of VaR and ES. Journal of Econometrics, 234(1):353– 370, May 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407622000380.

Li:2023:IML

[LWY23] Yong Li, Nianling Wang, and Jun Yu. Improved marginal likelihood estimation via power posteriors and importance sampling. Journal of Econometrics, 234(1):28–52, May 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002736.

Lim:2024:CLC

[LWZ24] Dennis Lim, Wenjie Wang, and Yichong Zhang. A conditional linear combination test with many weak instruments. Journal of Econometrics, 238(2):??, January 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623003184.

Luo:2023:IAM

[LX23] Yao Luo and Ruli Xiao. Identification of auction models using order statistics. Journal of Econometrics, 236(1):??, September 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623001513.

Luo:2022:IDG

[LXX22] Yao Luo, Ping Xiao, and Ruli Xiao. Identification of dynamic games with unobserved heterogeneity and multiple equilibria. Journal of Econometrics, 226(2):343–367, February 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001470.

Li:2021:SDP

[LY21] Liyao Li and Zhenlin Yang. Spatial dynamic panel data models with correlated random effects. Journal of Econometrics, 221(2):424–454, April 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0304407620302372.

Liu:2022:SSM

[LY22] Ruixuan Liu and Zhengfei Yu. Sample selection models with monotone control functions. Journal of Econometrics, 226(2):321–342, February 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407621001342.

Li:2020:DIC

[LYZ20] Yong Li, Jun Yu, and Tao Zeng. Deviance information criterion for latent variable models and misspecified models. Journal of Econometrics, 216(2):450–493, June 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619302271.

Lu:2020:GFT

[LZ20] Xiaohui Lu and Xu Zheng. A goodness-of-fit test for copulas based on martingale transformation. Journal of Econometrics, 215(1):84–117, March 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440761930185X.

Liao:2021:MAP

[LZGZ21] Jun Liao, Guohua Zou, Yan Gao, and Xinyu Zhang. Model averaging prediction for time series models with a diverging number of parameters. Journal of Econometrics, 223(1):190–221, July 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303493.

MacKinnon:2023:ULS

[Mac23] James G. MacKinnon. Using large samples in econometrics. Journal of Econometrics, 235(2):922–926, August 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001580.

Manski:2023:PPB

[Man23] Charles F. Manski. Probabilistic prediction for binary treatment choice: With focus on personalized medicine. Journal of Econometrics, 234(2):647–663, June 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001579.

Marx:2024:SBL

[Mar24] Philip Marx. Sharp bounds in the latent index selection model. Journal of Econometrics, 238(2):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002774.

Mavroeidis:2021:ESI

[Mav21] Sophocles Mavroeidis. Editorial for special issue: Vector autoregressions. Journal of Econometrics, 225(1):1, November 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001871.

Mezza:2021:IDE

[MB21] Alvaro Mezza and Moshe Buchinsky. Illegal drugs, education, and labor market outcomes. Journal of Econometrics, 223(2):454–484, August 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620303316.

Miller:2024:BRM

[MB24] J. Isaac Miller and William A. Brock. Beyond RCP8.5: Marginal mitigation using quasi-representative concentration pathways. Journal of Econometrics, 239(??):??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001792.

McFadden:2021:E

[McF21] Daniel McFadden. Epilogue. Journal of Econometrics, 222(1): 261–263, ???? 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407621000592.

Mehrabani:2023:EIL

[Meh23] Ali Mehrabani. Estimation and identification of latent group structures in panel data. Journal of Econometrics, 235(2):1464– 1482, August 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030440762200207X.

MaCurdy:2024:PPD

[MGSN24] Thomas MaCurdy, David Glick, Sonam Sherpa, and Sriniketh Nagavarapu. Profiling the plight of disconnected youth in America. Journal of Econometrics, 238(2):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002737.

Ma:2023:SST

[MGW23] Yingying Ma, Shaojun Guo, and Hansheng Wang. Sparse spatio-temporal autoregressions by profiling and bagging. Journal of Econometrics, 232(1):132–147, January 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762100035X.

Martellosio:2020:AQS

[MH20] Federico Martellosio and Grant Hillier. Adjusted QMLE for the spatial autoregressive parameter. Journal of Econometrics, 219(2):488–506, December 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620301081.

Martinez-Iriarte:2024:UEG

[MIMRS24] Julián Martínez-Iriarte, Gabriel Montes-Rojas, and Yixiao Sun. Unconditional effects of general policy interventions. Journal of Econometrics, 238(2):??, January 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002865.

Martinez-Iriarte:2020:ATU

[MISW20] Julián Martínez-Iriarte, Yixiao Sun, and Xuexin Wang. Asymptotic F tests under possibly weak identification. Journal of Econometrics, 218(1):140–177, September 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620300063.

Miftakhova:2020:SAH

[MJLS20] Alena Miftakhova, Kenneth L. Judd, Thomas S. Lontzek, and Karl Schmedders. Statistical approximation of high-dimensional climate models. Journal of Econometrics, 214(1):67–80, January 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407619301083.

Ma:2021:EIS

[MLG21] Shujie Ma, Oliver Linton, and Jiti Gao. Estimation and inference in semiparametric quantile factor models. Journal of Econometrics, 222(1):295–323, ???? 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407620301937.

Miao:2020:PTM

[MLS20] Ke Miao, Kunpeng Li, and Liangjun Su. Panel threshold models with interactive fixed effects. Journal of Econometrics, 219(1):137–170, November 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620302396.

Manski:2021:ECI

[MM21a] Charles F. Manski and Francesca Molinari. Estimating the COVID-19 infection rate: Anatomy of an inference problem. Journal of Econometrics, 220(1):181–192, January 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301676.

Meyer:2021:ETS

[MM21b] Bruce D. Meyer and Nikolas Mittag. An empirical total survey error decomposition using data combination. Journal of Econometrics, 224(2):286–305, October 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303687.

Maneesoonthorn:2020:HFJ

[MMF20] Worapree Maneesoonthorn, Gael M. Martin, and Catherine S. Forbes. High-frequency jump tests: Which test should we use? Journal of Econometrics, 219(2):478–487, December 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762030107X.

Ma:2023:IIT

[MMY23] Jun Ma, Vadim Marmer, and Zhengfei Yu. Inference on individual treatment effects in nonseparable triangular models. Journal of Econometrics, 235(2):2096–2124, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623001276.

Martin:2020:IEM

[MNP20] Gael M. Martin, K. Nadarajah, and D. S. Poskitt. Issues in the estimation of mis-specified models of fractionally integrated processes. Journal of Econometrics, 215(2):559–573, April 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761930209X.

MacKinnon:2023:CRI

[MNW23a] James G. MacKinnon, Morten Ørregaard Nielsen, and Matthew D. Webb. Cluster-robust inference: a guide to empirical practice. Journal of Econometrics, 232(2):272–299, February 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622000781.

MacKinnon:2023:TAL

[MNW23b] James G. MacKinnon, Morten Ørregaard Nielsen, and Matthew D. Webb. Testing for the appropriate level of clustering in linear regression models. Journal of Econometrics, 235(2):2027–2056, August 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623001021.

Matsushita:2023:SOR

[MO23] Yukitoshi Matsushita and Taisuke Otsu. Second-order refinements for t-ratios with many instruments. Journal of Econometrics, 232(2):346–366, February 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001901.

Miller:2024:ITIa

[MP24a] J. Isaac Miller and Felix Pretis. Introduction to the Themed Issue on Climate Econometrics. Journal of Econometrics, 238(2):

??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623003445.

Miller:2024:ITIb

[MP24b] J. Isaac Miller and Felix Pretis. Introduction to the Themed Issue on Climate Econometrics. Journal of Econometrics, 239(??): ??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623003603.

Miao:2023:HDV

[MPS23] Ke Miao, Peter C. B. Phillips, and Liangjun Su. Highdimensional VARs with common factors. Journal of Econometrics, 233(1):155–183, March 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440762200032X.

Meitz:2021:TOD

[MS21a] Mika Meitz and Pentti Saikkonen. Testing for observationdependent regime switching in mixture autoregressive models. Journal of Econometrics, 222(1):601–624, May 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407620302438.

Mogliani:2021:BMP

[MS21b] Matteo Mogliani and Anna Simoni. Bayesian MIDAS penalized regressions: Estimation, selection, and prediction. Journal of Econometrics, 222(1):833–860, May 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407620302207.

Miao:2020:PTR

[MSW20] Ke Miao, Liangjun Su, and Wendun Wang. Panel threshold regressions with latent group structures. Journal of Econometrics, 214(2):451–481, February 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407619301708.

Ma:2023:GFL

[MT23a] Chenchen Ma and Yundong Tu. Group fused lasso for large factor models with multiple structural breaks. Journal of Econometrics, 233(1):132–154, March 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622000331.

Ma:2023:SEM

[MT23b] Chenchen Ma and Yundong Tu. Shrinkage estimation of multiple threshold factor models. Journal of Econometrics, 235(2):1876–1892, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407623000489.

Man:2024:RRE

[MTWZ24] Rebeka Man, Kean Ming Tan, Zian Wang, and Wen-Xin Zhou. Retire: Robust expectile regression in high dimensions. Journal of Econometrics, 239(2):??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623001537.

MacKinnon:2020:RID

[MW20] James G. MacKinnon and Matthew D. Webb. Randomization inference for difference-in-differences with few treated clusters. Journal of Econometrics, 218(2):435–450, October 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301445.

Mayer:2023:EIF

[MW23] Alexander Mayer and Dominik Wied. Estimation and inference in factor copula models with exogenous covariates. Journal of Econometrics, 235(2):1500–1521, August 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623000039.

Mykland:2021:OAV

[MZ21] Per A. Mykland and Lan Zhang. The observed asymptotic variance: Hard edges, and a regression approach. Journal of Econometrics, 222(1):411–428, ???? 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407620301986.

Nguyen:2020:LVU

[NEFG20] Giang Nguyen, Robert Engle, Michael Fleming, and Eric Ghysels. Liquidity and volatility in the U.S. Treasury market. Journal of Econometrics, 217(2):207–229, August 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619302465.

Narisetty:2022:CQR

[NK22] Naveen Narisetty and Roger Koenker. Censored quantile regression survival models with a cure proportion. Journal of Econometrics, 226(1):192–203, January 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303997.

Nagler:2022:SVC

[NKM22] Thomas Nagler, Daniel Krüger, and Aleksey Min. Stationary vine copula models for multivariate time series. Journal of Econometrics, 227(2):305–324, April 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621003043.

Norets:2022:ABE

[NP22] Andriy Norets and Justinas Pelenis. Adaptive Bayesian estimation of conditional discrete-continuous distributions with an application to stock market trading activity. Journal of Econometrics, 230(1):62–82, September 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762100261X.

Ng:2021:AIP

[NQV21] Serena Ng, Zhongjun Qu, and Timothy Vogelsang. Annals issue: PI-Day honoring Pierre Perron. Journal of Econometrics, 224(1):1–2, September 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407621001287.

Nicolau:2023:TIE

[NRS23] João Nicolau, Paulo M. M. Rodrigues, and Marian Z. Stoykov. Tail index estimation in the presence of covariates: Stock returns’ tail risk dynamics. Journal of Econometrics, 235(2):2266– 2284, August 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030440762300146X.

Newey:2021:CVD

[NS21] Whitney Newey and Sami Stouli. Control variables, discrete instruments, and identification of structural functions. Journal of Econometrics, 222(1):73–88, ???? 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407620302499.

Norets:2024:SBE

[NS24] Andriy Norets and Kenichi Shimizu. Semiparametric Bayesian estimation of dynamic discrete choice models. Journal of Econometrics, 238(2):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407623003585.

Norkute:2021:IVE

[NSYC21] Milda Norkute, Vasilis Sarafidis, Takashi Yamagata, and Guowei Cui. Instrumental variable estimation of dynamic linear panel data models with defactored regressors and a multifactor error structure. Journal of Econometrics, 220(2):416–446, February 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620301275.

Norkute:2021:FAA

[NW21] Milda Norkute and Joakim Westerlund. The factor analytical approach in near unit root interactive effects panels. Journal of Econometrics, 221(2):569–590, April 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407620302682.

Otneim:2020:PLF

[OJT20] Håkon Otneim, Martin Jullum, and Dag Tjøstheim. Pairwise local Fisher and naive Bayes: Improving two standard discriminants. Journal of Econometrics, 216(1):284–304, May 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620300257.

Olea:2021:MLP

[ON21] José Luis Montiel Olea and James Nesbit. (Machine) learning parameter regions. Journal of Econometrics, 222(1):716– 744, May 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407620302177.

Oh:2024:DFC

[OP24] Dong Hwan Oh and Andrew J. Patton. Dynamic factor copula models with estimated cluster assignments. Journal of Econometrics, 237(2C):??, December 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407622002135.

Odendahl:2024:EFP

[ORS24] Florens Odendahl, Barbara Rossi, and Tatevik Sekhposyan. Evaluating forecast performance with state dependence. Journal of Econometrics, 237(2C):??, December 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407621002657.

Olea:2021:ISV

[OSW21] José L. Montiel Olea, James H. Stock, and Mark W. Watson. Inference in structural vector autoregressions identified with an external instrument. Journal of Econometrics, 225(1):74–87, November 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620302311.

Okui:2021:HSB

[OW21] Ryo Okui and Wendun Wang. Heterogeneous structural breaks in panel data models. Journal of Econometrics, 220(2):447–473, February 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620301287.

Park:2020:VDM

[Par20] Yang-Ho Park. Variance disparity and market frictions. Journal of Econometrics, 214(2):326–348, February 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619301654.

Pang:2021:EMB

[PDC21] Tianxiao Pang, Lingjie Du, and Terence Tai-Leung Chong. Estimating multiple breaks in nonstationary autoregressive models. Journal of Econometrics, 221(1):277–311, March 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620302116.

Petrova:2022:AVB

[Pet22] Katerina Petrova. Asymptotically valid Bayesian inference in the presence of distributional misspecification in VAR models. Journal of Econometrics, 230(1):154–182, September 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621000865.

Pereda-Fernandez:2023:IET

[PF23] Santiago Pereda-Fernández. Identification and estimation of triangular models with a binary treatment. Journal of Econometrics, 234(2):585–623, June 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622000410.

Powell:2021:DMH

[PG21] David Powell and Dana Goldman. Disentangling moral hazard and adverse selection in private health insurance. Journal of Econometrics, 222(1):141–160, ???? 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407620302529.

Philip:2020:EPP

[Phi20] R. Philip. Estimating permanent price impact via machine learning. Journal of Econometrics, 215(2):414–449, April 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619302052.

Phillips:2024:HDI

[PK24] Peter C. B. Phillips and Igor L. Kheifets. High-dimensional IV cointegration estimation and inference. Journal of Econometrics, 238(2):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440762300338X.

Petukhina:2024:RM

[PKHZ24] Alla Petukhina, Yegor Klochkov, Wolfgang Karl Härdle, and Nikita Zhivotovskiy. Robustifying Markowitz. Journal of Econometrics, 239(2):??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0304407623000180.

Phillips:2020:EEE

[PLS20] Peter C. B. Phillips, Thomas Leirvik, and Trude Storelvmo. Econometric estimates of Earth’s transient climate sensitivity. Journal of Econometrics, 214(1):6–32, January 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619301058.

Proietti:2024:MCC

[PM24] Tommaso Proietti and Federico Maddanu. Modelling cycles in climate series: the fractional sinusoidal waveform process. Journal of Econometrics, 239(??):??, February 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622000987.

Pouliot:2023:SEM

[Pou23] Guillaume Allaire Pouliot. Spatial econometrics for misaligned data. Journal of Econometrics, 232(1):168–190, January 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001627.

Powell:2023:DWS

[Pow23] James L. Powell. Discussion of “What is a standard error?”. Journal of Econometrics, 237(1):??, November 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002348. See [Gel23].

Pei:2024:LCC

[PPX24] Youquan Pei, Heng Peng, and Jinfeng Xu. A latent class Cox model for heterogeneous time-to-event data. Journal of Econometrics, 239(2):??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407622001841.

Pretis:2020:EMC

[Pre20] Felix Pretis. Econometric modelling of climate systems: the equivalence of energy balance models and cointegrated vector autoregressions. Journal of Econometrics, 214(1):256–273, January 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407619301162.

Pathak:2021:HWD

[PS21a] Parag A. Pathak and Peng Shi. How well do structural demand models work? Counterfactual predictions in school choice. Journal of Econometrics, 222(1):161–195, ???? 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407620302530.

Perera:2021:BBP

[PS21b] Indeewara Perera and Mervyn J. Silvapulle. Bootstrap based probability forecasting in multiplicative error models. Journal of Econometrics, 221(1):1–24, March 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620300440.

Pellatt:2023:ATR

[PS23a] Daniel F. Pellatt and Yixiao Sun. Asymptotic F test in regressions with observations collected at high frequency over long span. Journal of Econometrics, 235(2):1281–1309, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001919.

Perera:2023:BST

[PS23b] Indeewara Perera and Mervyn J. Silvapulle. Bootstrap specification tests for dynamic conditional distribution models. Journal of Econometrics, 235(2):949–971, August 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001634.

Pettenuzzo:2022:CPS

[PST22] Davide Pettenuzzo, Yong Song, and Allan Timmermann. Corrigendum to “Predictability of stock returns and asset allocation under structural breaks” [j. econometrics 164 (2011) 60–78]. Journal of Econometrics, 227(2):513–517, April 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621000476.

Pettenuzzo:2023:DSC

[PST23] Davide Pettenuzzo, Riccardo Sabbatucci, and Allan Timmermann. Dividend suspensions and cash flows during the Covid-19 pandemic: a dynamic econometric model. Journal of Econometrics, 235(2):1522–1541, August 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762300012X.

Pettenuzzo:2011:PSR

[PT11] Davide Pettenuzzo and Allan Timmermann. Predictability of stock returns and asset allocation under structural breaks. Journal of Econometrics, 164(1):60–78, September 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611000479. See corrigendum [PST22].

Patnaik:2022:RHR

[PVWZ22] Arpita Patnaik, Joanna Venator, Matthew Wiswall, and Basit Zafar. The role of heterogeneous risk preferences, discount rates, and earnings expectations in college major choice. Journal of Econometrics, 231(1):98–122, November 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620302773.

Park:2021:NEJ

[PW21] Joon Y. Park and Bin Wang. Nonparametric estimation of jump diffusion models. Journal of Econometrics, 222(1):688– 715, May 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407620302189.

Phillips:2022:FCP

[PW22] Peter C. B. Phillips and Ying Wang. Functional coefficient panel modeling with communal smoothing covariates. Journal of Econometrics, 227(2):371–407, April 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621000944.

Phillips:2023:WBC

[PW23] Peter C. B. Phillips and Ying Wang. When bias contributes to variance: True limit theory in functional coefficient cointegrating regression. Journal of Econometrics, 232(2):469–489, February 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002190.

Pesaran:2020:EAP

[PY20] M. Hashem Pesaran and Cynthia Fan Yang. Econometric analysis of production networks with dominant units. Journal of Econometrics, 219(2):507–541, December 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301093.

Pesaran:2021:EIS

[PY21] M. Hashem Pesaran and Cynthia Fan Yang. Estimation and inference in spatial models with dominant units. Journal of Econometrics, 221(2):591–615, April 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0304407620302360.

Peng:2022:IMS

[PY22] Jingfu Peng and Yuhong Yang. On improvability of model selection by model averaging. Journal of Econometrics, 229(2):246– 262, August 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620303973.

Qu:2021:ESM

[QfLY21] Xi Qu, Lung fei Lee, and Chao Yang. Estimation of a SAR model with endogenous spatial weights constructed by bilateral variables. Journal of Econometrics, 221(1):180–197, March 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620302281.

Qu:2024:CFP

[QTZ24] Ritong Qu, Allan Timmermann, and Yinchu Zhu. Comparing forecasting performance in cross-sections. Journal of Econometrics, 237(2C):??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0304407621002256.

Rabovic:2023:ESS

[RC23] Renata Rabovic and Pavel Cízek. Estimation of spatial sample selection models: a partial maximum likelihood approach. Journal of Econometrics, 232(1):214–243, January 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002815.

Rossi:2023:SAE

[RL23] Francesca Rossi and Offer Lieberman. Spatial autoregressions with an extended parameter space and similarity-based weights. Journal of Econometrics, 235(2):1770–1798, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623000283.

Ren:2024:DEI

[RLZ^{+}24] Yimeng Ren, Zhe Li, Xuening Zhu, Yuan Gao, and Hansheng Wang. Distributed estimation and inference for spatial autoregression model with large scale networks. Journal of Econometrics, 238(2):??, January 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623003457.

Royer:2023:CAP

[Roy23] Julien Royer. Conditional asymmetry in Power ARCH(∞) models. Journal of Econometrics, 234(1):178–204, May 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621003031.

Rossi:2023:HOL

[RR23] Francesca Rossi and Peter M. Robinson. Higher-order least squares inference for spatial autoregressions. Journal of Econometrics, 232(1):244–269, January 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622000458.

Roth:2023:WTD

[RSBP23] Jonathan Roth, Pedro H. C. Sant’Anna, Alyssa Bilinski, and John Poe. What’s trending in difference-in-differences? A synthesis of the recent econometrics literature. Journal of Econometrics, 235(2):2218–2244, August 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623001318.

Rombouts:2020:DVR

[RSV20] Jeroen V. K. Rombouts, Lars Stentoft, and Francesco Violante. Dynamics of variance risk premia: a new model for disentangling the price of risk. Journal of Econometrics, 217(2):312– 334, August 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407619302520.

Rombouts:2020:NFE

[RSVZ20] Jeroen V. K. Rombouts, Olivier Scaillet, David Veredas, and Jean-Michel Zakoian. Nonlinear financial econometrics JoE special issue introduction. Journal of Econometrics, 217(2):203– 206, August 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407619302453.

Roth:2022:BAP

[RSW22] Christopher Roth, Sonja Settele, and Johannes Wohlfart. Beliefs about public debt and the demand for government spending. Journal of Econometrics, 231(1):165–187, November 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621000397.

Rho:2021:ITS

[RV21] Seunghwa Rho and Timothy J. Vogelsang. Inference in time series models using smoothed-clustered standard errors. Journal of Econometrics, 224(1):113–133, September 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620302608.

Rothe:2020:EDF

[RW20] Christoph Rothe and Dominik Wied. Estimating derivatives of function-valued parameters in a class of moment condition models. Journal of Econometrics, 217(1):1–19, July 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619302416.

Reuvers:2024:SGY

[RW24] Hanno Reuvers and Etienne Wijler. Sparse generalized Yule–Walker estimation for large spatio-temporal autoregressions with an application to NO_{2} satellite data. Journal of Econometrics, 239(??):??, February 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002361.

Sun:2021:EDT

[SA21] Liyang Sun and Sarah Abraham. Estimating dynamic treatment effects in event studies with heterogeneous treatment effects. Journal of Econometrics, 225(2):175–199, December 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762030378X.

So:2022:EEH

[SCC22] Mike K. P. So, Thomas W. C. Chan, and Amanda M. Y. Chu. Efficient estimation of high-dimensional dynamic covariance by risk factor mapping: Applications for financial risk management. Journal of Econometrics, 227(1):151–167, March 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301664.

Schmieder:2023:EAW

[Sch23] Johannes F. Schmieder. Establishment age and wages. Journal of Econometrics, 233(2):424–442, April 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623000350.

Semenova:2023:DML

[Sem23] Vira Semenova. Debiased machine learning of set-identified linear models. Journal of Econometrics, 235(2):1725–1746, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762300026X.

Sentana:2024:FU

[Sen24] Enrique Sentana. Finite underidentification. Journal of Econometrics, 240(1):??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407624000381.

Sonksen:2021:EAP

[SG21] Jantje Sönksen and Joachim Grammig. Empirical asset pricing with multi-period disaster risk: a simulation-based approach. Journal of Econometrics, 222(1):805–832, May 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407620302748.

Shi:2023:FSP

[SH23] Zhentao Shi and Jingyi Huang. Forward-selected panel data approach for program evaluation. Journal of Econometrics, 234(2):512–535, June 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407621001536.

Shimizu:2023:APB

[Shi23] Kenichi Shimizu. Asymptotic properties of Bayesian inference in linear regression with a structural break. Journal of Econometrics, 235(1):202–219, July 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762200077X.

Sun:2021:TVM

[SHL^{+}21] Yuying Sun, Yongmiao Hong, Tae-Hwy Lee, Shouyang Wang, and Xinyu Zhang. Time-varying model averaging. Journal of Econometrics, 222(2):974–992, June 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303067.

Sun:2023:PTV

[SHWZ23] Yuying Sun, Yongmiao Hong, Shouyang Wang, and Xinyu Zhang. Penalized time-varying model averaging. Journal of Econometrics, 235(2):1355–1377, August 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622002019.

Shin:2023:ARL

[SKF23] Minseok Shin, Donggyu Kim, and Jianqing Fan. Adaptive robust large volatility matrix estimation based on high-frequency financial data. Journal of Econometrics, 237(1):??, November 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002300.

Song:2021:VAR

[SKY^{+}21] Xinyu Song, Donggyu Kim, Huiling Yuan, Xiangyu Cui, Zhiping Lu, Yong Zhou, and Yazhen Wang. Volatility analysis with realized GARCH–Itô models. Journal of Econometrics, 222(1): 393–410, ???? 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407620301974.

She:2020:IHT

[SL20] Rui She and Shiqing Ling. Inference in heavy-tailed vector error correction models. Journal of Econometrics, 214(2):433–450, February 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407619301691.

Solvsten:2020:REM

[Søl20] Mikkel Sølvsten. Robust estimation with many instruments. Journal of Econometrics, 214(2):495–512, February 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619301721.

Schwartz:2024:LLN

[SS24a] Jacob Schwartz and Kyungchul Song. The law of large numbers for large stable matchings. Journal of Econometrics, 241(1): ??, April 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407624000885.

Startz:2024:VRC

[SS24b] Richard Startz and Douglas G. Steigerwald. The variance of regression coefficients when the population is finite. Journal of Econometrics, 240(1):??, March 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407624000277.

Schumann:2021:ILB

[SST21] Martin Schumann, Thomas A. Severini, and Gautam Tripathi. Integrated likelihood based inference for nonlinear panel data models with unobserved effects. Journal of Econometrics, 223(1):73–95, July 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620303432.

Schumann:2023:RSI

[SST23] Martin Schumann, Thomas A. Severini, and Gautam Tripathi. The role of score and information bias in panel data likelihoods. Journal of Econometrics, 235(2):1215–1238, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001889.

Shapiro:2022:MNS

[SSW22] Adam Hale Shapiro, Moritz Sudhof, and Daniel J. Wilson. Measuring news sentiment. Journal of Econometrics, 228(2):221– 243, June 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620303535.

Sasaki:2023:DDC

[STXH23] Yuya Sasaki, Yuya Takahashi, Yi Xin, and Yingyao Hu. Dynamic discrete choice models with incomplete data: Sharp identification. Journal of Econometrics, 236(1):??, September 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623001550.

Su:2021:MSU

[Su21] Jiun-Hua Su. Model selection in utility-maximizing binary prediction. Journal of Econometrics, 223(1):96–124, July 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303420.

Sasaki:2023:EIP

[SU23] Yuya Sasaki and Takuya Ura. Estimation and inference for policy relevant treatment effects. Journal of Econometrics, 234(2): 394–450, June 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621001494.

Sun:2024:IVH

[Sun24] Zhenting Sun. Instrument validity for heterogeneous causal effects. Journal of Econometrics, 237(2A):??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407623002397.

Sarafidis:2021:CYP

[SW21a] Vasilis Sarafidis and Tom Wansbeek. Celebrating 40 years of panel data analysis: Past, present and future. Journal of Econometrics, 220(2):215–226, February 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301822.

Smeekes:2021:AAT

[SW21b] Stephan Smeekes and Etienne Wijler. An automated approach towards sparse single-equation cointegration modelling. Journal of Econometrics, 221(1):247–276, March 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620302190.

Sabzikar:2020:ATN

[SWP20] Farzad Sabzikar, Qiying Wang, and Peter C. B. Phillips. Asymptotic theory for near integrated processes driven by tempered linear processes. Journal of Econometrics, 216(1):192– 202, May 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030440762030018X.

Su:2023:ILG

[SWX23] Liangjun Su, Wuyi Wang, and Xingbai Xu. Identifying latent group structures in spatial dynamic panels. Journal of Econometrics, 235(2):1955–1980, August 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623000738.

Sun:2024:MKQ

[SWZZ24] Yan Sun, Chuang Wan, Wenyang Zhang, and Wei Zhong. A multi-kink quantile regression model with common structure for panel data analysis. Journal of Econometrics, 239(2):??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001178.

Saart:2022:FTS

[SX22] Patrick W. Saart and Yingcun Xia. Functional time series approach to analyzing asset returns co-movements. Journal of Econometrics, 229(1):127–151, July 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621000804.

Sun:2023:ILF

[SXZ23] Yucheng Sun, Wen Xu, and Chuanhai Zhang. Identifying latent factors based on high-frequency data. Journal of Econometrics, 233(1):251–270, March 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622000938.

Sun:2020:TOK

[SY20] Yixiao Sun and Jingjing Yang. Testing-optimal kernel choice in HAR inference. Journal of Econometrics, 219(1):123–136, November 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030440762030213X.

SantAnna:2020:DRD

[SZ20] Pedro H. C. Sant’Anna and Jun Zhao. Doubly robust difference-in-differences estimators. Journal of Econometrics, 219(1):101–122, November 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620301901.

Shi:2024:EPM

[SZ24] Peng Shi and Zifeng Zhao. Enhanced pricing and management of bundled insurance risks with dependence-aware prediction using pair copula construction. Journal of Econometrics, 240(1): ??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407624000228.

Tamer:2021:IPE

[Tam21] Elie Tamer. Introduction to pandemic econometrics/Covid-19 pandemic. Journal of Econometrics, 220(1):1, January 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303663.

Tauchen:2022:NDN

[Tau22] George Tauchen. New directions in nonlinear structural estimation: Bayes and frequentist. Journal of Econometrics, 228(1): 1–3, May 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621002426.

Tchatoka:2020:ETI

[TD20] Firmin Doko Tchatoka and Jean-Marie Dufour. Exogeneity tests, incomplete models, weak identification and non-Gaussian distributions: Invariance and finite-sample distributional theory. Journal of Econometrics, 218(2):390–418, October 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301421.

Tu:2022:NIQ

[TLW22] Yundong Tu, Han-Ying Liang, and Qiying Wang. Nonparametric inference for quantile cointegrations with stationary covariates. Journal of Econometrics, 230(2):453–482, October 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001731.

Todorov:2022:NJV

[Tod22] Viktor Todorov. Nonparametric jump variation measures from options. Journal of Econometrics, 230(2):255–280, October 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001263.

Toulis:2021:ECP

[Tou21] Panos Toulis. Estimation of Covid-19 prevalence from serology tests: a partial identification approach. Journal of Econometrics, 220(1):193–213, January 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762030350X.

Tsionas:2023:BAN

[TPZ23] Mike Tsionas, Christopher F. Parmeter, and Valentin Zelenyuk. Bayesian Artificial Neural Networks for frontier efficiency analysis. Journal of Econometrics, 236(2):??, October 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002075.

Trapani:2021:ITH

[Tra21] Lorenzo Trapani. Inferential theory for heterogeneity and cointegration in large panels. Journal of Econometrics, 220(2):474– 503, February 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620301299.

Tsay:2020:TSC

[Tsa20] Ruey S. Tsay. Testing serial correlations in high-dimensional time series via extreme value theory. Journal of Econometrics, 216(1):106–117, May 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620300130.

Tuvaandorj:2020:RDD

[Tuv20] Purevdorj Tuvaandorj. Regression discontinuity designs, white noise models, and minimax. Journal of Econometrics, 218 (2):587–608, October 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620301500.

Tu:2022:SFC

[TW22] Yundong Tu and Ying Wang. Spurious functional-coefficient regression models and robust inference with marginal integration. Journal of Econometrics, 229(2):396–421, August 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621000713.

Tu:2023:PSR

[TX23] Yundong Tu and Xinling Xie. Penetrating sporadic return predictability. Journal of Econometrics, 237(1):??, November 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002257.

Todorov:2023:BRS

[TZ23] Viktor Todorov and Yang Zhang. Bias reduction in spot volatility estimation from options. Journal of Econometrics, 234(1): 53–81, May 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621002785.

Tommasi:2024:BPB

[TZ24] Denni Tommasi and Lina Zhang. Bounding program benefits when participation is misreported. Journal of Econometrics, 238(1):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407623002725.

uille:2023:NDD

[uHS23] Xavier D Haultf uille, Stefan Hoderlein, and Yuya Sasaki. Nonparametric difference-in-differences in repeated cross-sections with continuous treatments. Journal of Econometrics, 234(2): 664–690, June 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407622001452.

Umlandt:2024:SDA

[Uml24] Dennis Umlandt. Score-driven asset pricing: Predicting timevarying risk premia based on cross-sectional model performance. Journal of Econometrics, 237(2C):??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407623001641.

Urga:2024:EIH

[UW24] Giovanni Urga and Fa Wang. Estimation and inference for high dimensional factor model with regime switching. Journal of Econometrics, 241(2):??, April 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407624000988.

Ullah:2023:SPL

[UWY23] Aman Ullah, Tao Wang, and Weixin Yao. Semiparametric partially linear varying coefficient modal regression. Journal of Econometrics, 235(2):1001–1026, August 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762200166X.

Vazquez-Bare:2023:IES

[VB23a] Gonzalo Vazquez-Bare. Identification and estimation of spillover effects in randomized experiments. Journal of Econometrics, 237(1):??, November 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407621003067.

Viviano:2023:SLM

[VB23b] Davide Viviano and Jelena Bradic. Synthetic learner: Modelfree inference on treatments over time. Journal of Econometrics, 234(2):691–713, June 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440762200152X.

vandenBerg:2021:GSA

[vdBJMN21] Gerard. J. van den Berg, Lena Janys, Enno Mammen, and Jens Perch Nielsen. A general semiparametric approach to inference with marker-dependent hazard rate models. Journal of Econometrics, 221(1):43–67, March 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620300439.

vonGaudecker:2022:HHS

[vGW22] Hans-Martin von Gaudecker and Axel Wogrolly. Heterogeneity in households’ stock market beliefs: Levels, dynamics, and epistemic uncertainty. Journal of Econometrics, 231(1):232–247, November 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407621000403.

Vilhuber:2023:RTV

[Vil23] Lars Vilhuber. Reproducibility and transparency versus privacy and confidentiality: Reflections from a data editor. Journal of Econometrics, 235(2):2285–2294, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623001471.

Vogt:2020:MCN

[VL20] Michael Vogt and Oliver Linton. Multiscale clustering of nonparametric regression curves. Journal of Econometrics, 216(1): 305–325, May 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620300269.

VandeSijpe:2023:PCL

[VW23] Nicolas Van de Sijpe and Frank Windmeijer. On the power of the conditional likelihood ratio and related tests for weakinstrument robust inference. Journal of Econometrics, 235(1):

82–104, July 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407622000367.

Wang:2022:MLE

[Wan22] Fa Wang. Maximum likelihood estimation and inference for high dimensional generalized factor models with application to factor-augmented regressions. Journal of Econometrics, 229(1): 180–200, July 2022. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620303894.

Wang:2023:SBS

[Wan23] Ao Wang. Sieve BLP: a semi-nonparametric model of demand for differentiated products. Journal of Econometrics, 235(2):325–351, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407622000860.

Wang:2024:VPC

[WAZ24] Weichen Wang, Ran An, and Ziwei Zhu. Volatility prediction comparison via robust volatility proxies: an empirical deviation perspective. Journal of Econometrics, 239(2):??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623003494.

Wang:2022:ASB

[WCLC22] Christina Dan Wang, Zhao Chen, Yimin Lian, and Min Chen. Asset selection based on high frequency Sharpe ratio. Journal of Econometrics, 227(1):168–188, March 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620302244.

Wang:2020:UDP

[WCWL20] Luheng Wang, Zhao Chen, Christina Dan Wang, and Runze Li. Ultrahigh dimensional precision matrix estimation via refitted cross validation. Journal of Econometrics, 215(1):118– 130, March 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030440761930171X.

Wan:2022:GFT

[WD22] Phyllis Wan and Richard A. Davis. Goodness-of-fit testing for time series models via distance covariance. Journal of Econometrics, 227(1):4–24, March 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620302256.

Wan:2022:RTB

[WFL22] Runqing Wan, Andras Fulop, and Junye Li. Real-time Bayesian learning and bond return predictability. Journal of Econometrics, 230(1):114–130, September 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621000877.

Wagner:2020:FMO

[WGH20] Martin Wagner, Peter Grabarczyk, and Seung Hyun Hong. Fully modified OLS estimation and inference for seemingly unrelated cointegrating polynomial regressions and the environmental Kuznets curve for carbon dioxide emissions. Journal of Econometrics, 214(1):216–255, January 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619301150.

Williams:2020:NID

[Wil20] Benjamin Williams. Nonparametric identification of discrete choice models with lagged dependent variables. Journal of Econometrics, 215(1):286–304, March 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619301812.

Windmeijer:2024:TUL

[Win24] Frank Windmeijer. Testing underidentification in linear models, with applications to dynamic panel and asset pricing models. Journal of Econometrics, 240(2):??, March 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762100097X.

Wu:2024:BNI

[WLFF24] Yujia Wu, Wei Lan, Xinyan Fan, and Kuangnan Fang. Bipartite network influence analysis of a two-mode network. Journal of Econometrics, 239(2):??, February 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002786.

Wang:2024:RBM

[WLFM24] Hongfei Wang, Binghui Liu, Long Feng, and Yanyuan Ma. Rank-based max-sum tests for mutual independence of highdimensional random vectors. Journal of Econometrics, 238(1): ??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623002944.

Wan:2024:MFZ

[WLLS24] Runzhe Wan, Yingying Li, Wenbin Lu, and Rui Song. Mining the factor zoo: Estimation of latent factor models with sufficient proxies. Journal of Econometrics, 239(2):??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623000179.

Webb:2021:ACD

[WML21] Ryan Webb, Nitin Mehta, and Ifat Levy. Assessing consumer demand with noisy neural measurements. Journal of Econometrics, 222(1):89–106, ???? 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0304407620302505.

Woodcock:2023:DDW

[Woo23] Simon D. Woodcock. The determinants of displaced workers’ wages: Sorting, matching, selection, and the Hartz reforms. Journal of Econometrics, 233(2):568–595, April 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762200118X.

Wooldridge:2024:WSE

[Woo24] Jeffrey M. Wooldridge. What is a standard error? (And how should we compute it?). Journal of Econometrics, 237(2A):??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407623002336.

Wang:2021:NEL

[WPLL21] Hanchao Wang, Bin Peng, Degui Li, and Chenlei Leng. Nonparametric estimation of large covariance matrices with conditional sparsity. Journal of Econometrics, 223(1):53–72, July 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303456.

Wang:2024:PDM

[WPS24] Yiren Wang, Peter C. B. Phillips, and Liangjun Su. Panel data models with time-varying latent group structures. Journal of Econometrics, 240(1):??, March 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407624000319.

Wang:2021:ILG

[WS21] Wuyi Wang and Liangjun Su. Identifying latent group structures in nonlinear panels. Journal of Econometrics, 220(2):272– 295, February 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620301214.

Wei:2024:ECE

[WTH24] Bo Wei, Kean Ming Tan, and Xuming He. Estimation of complier expected shortfall treatment effects with a binary instrumental variable. Journal of Econometrics, 238(2):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002889.

Wang:2022:EIA

[WX22] Yulong Wang and Zhijie Xiao. Estimation and inference about tail features with tail censored data. Journal of Econometrics, 230(2):363–387, October 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001548.

Wang:2023:MFR

[WXY23] Xiaohu Wang, Weilin Xiao, and Jun Yu. Modeling and forecasting realized volatility with the fractional Ornstein–Uhlenbeck process. Journal of Econometrics, 232(2):389–415, February 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621002037.

Winston:2021:VSC

[WY21] Clifford Winston and Jia Yan. Vehicle size choice and automobile externalities: a dynamic analysis. Journal of Econometrics, 222(1):196–218, ???? 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0304407620302542.

Wang:2022:TPJ

[WZ22a] Bin Wang and Xu Zheng. Testing for the presence of jump components in jump diffusion models. Journal of Econometrics, 230(2):483–509, October 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407621001779.

Werker:2022:STH

[WZ22b] Bas J. M. Werker and Bo Zhou. Semiparametric testing with highly persistent predictors. Journal of Econometrics, 227(2):347–370, April 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407621001573.

Wang:2024:WBI

[WZ24] Wenjie Wang and Yichong Zhang. Wild bootstrap inference for instrumental variables regressions with weak and few clusters. Journal of Econometrics, 241(1):??, April 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407624000733.

Wang:2024:HDL

[WZL24] Di Wang, Yao Zheng, and Guodong Li. High-dimensional lowrank tensor autoregressive time series modeling. Journal of Econometrics, 238(1):??, January 2024. CODEN JECMB6.

ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0304407623002609.

Wang:2022:HQE

[WZLL22] Guochang Wang, Ke Zhu, Guodong Li, and Wai Keung Li. Hybrid quantile estimation for asymmetric power GARCH models. Journal of Econometrics, 227(1):264–284, March 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620302220.

Wei:2024:IBP

[WZZW24] Waverly Wei, Yuqing Zhou, Zeyu Zheng, and Jingshen Wang. Inference on the best policies with many covariates. Journal of Econometrics, 239(2):??, February 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623001549.

Xiong:2023:LDL

[XP23] Ruoxuan Xiong and Markus Pelger. Large dimensional latent factor modeling with missing observations and applications to causal inference. Journal of Econometrics, 233(1):271– 301, March 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407622000914.

Xu:2020:ILR

[Xu20] Ke-Li Xu. Inference of local regression in the presence of nuisance parameters. Journal of Econometrics, 218(2):532–560, October 2020. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620301482.

Yang:2020:TIR

[Yan20] Xiye Yang. Time-invariant restrictions of volatility functionals: Efficient estimation and specification tests. Journal of Econometrics, 215(2):486–516, April 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619302088.

Yang:2020:DML

[YCK20] Jui-Chung Yang, Hui-Ching Chuang, and Chung-Ming Kuan. Double machine learning with gradient boosting and its application to the big N audit quality effect. Journal of Econometrics, 216(1):268–283, May 2020. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620300245.

Yang:2021:EDP

[YfL21] Kai Yang and Lung fei Lee. Estimation of dynamic panel spatial vector autoregression: Stability and spatial multivariate cointegration. Journal of Econometrics, 221(2):337–367, April 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-

6895 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S030440762030227X.

Yu:2024:SCV

[YGV24] Lu Yu, Jiaying Gu, and Stanislav Volgushev. Spectral clustering with variance information for group structure estimation in panel data. Journal of Econometrics, 241(1):??, April 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407624000551.

Yu:2022:PEL

[YHKZ22] Long Yu, Yong He, Xinbing Kong, and Xinsheng Zhang. Projected estimation for large-dimensional matrix factor models. Journal of Econometrics, 229(1):201–217, July 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001123.

Yu:2023:MPR

[YLC^{+}23] Tao Yu, Pengfei Li, Baojiang Chen, Ao Yuan, and Jing Qin. Maximum pairwise-rank-likelihood-based inference for the semiparametric transformation model. Journal of Econometrics, 235(2):454–469, August 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407622001208.

Yousuf:2021:BHD

[YN21] Kashif Yousuf and Serena Ng. Boosting high dimensional predictive regressions with time varying parameters. Journal of Econometrics, 224(1):60–87, September 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620302827.

Yoon:2024:IES

[Yoo24] Jangsu Yoon. Identification and estimation of sequential games of incomplete information with multiple equilibria. Journal of Econometrics, 238(2):??, January 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002853.

Yang:2021:EAE

[YS21] Yimin Yang and Peter Schmidt. An econometric approach to the estimation of multi-level models. Journal of Econometrics, 220(2):532–543, February 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620301317.

Yu:2022:BME

[Yu22] Jun Yu. Bayesian methods in economics and finance: Editor’s introduction. Journal of Econometrics, 230(1):1–2, September 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621003109.

Yu:2023:WUT

[Yu23] Bin Yu. What is uncertainty in today’s practice of data science? Journal of Econometrics, 237(1):??, November 2023. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762300235X.

Yu:2024:PET

[YYX24] Xiufan Yu, Jiawei Yao, and Lingzhou Xue. Power enhancement for testing multi-factor asset pricing models via Fisher’s method. Journal of Econometrics, 239(2):??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623001525.

Yang:2021:THD

[YZC21] Xinxin Yang, Xinghua Zheng, and Jiaqi Chen. Testing highdimensional covariance matrices under the elliptical distribution and beyond. Journal of Econometrics, 221(2):409–423, April 2021. CODEN JECMB6. ISSN 0304-4076 (print), 1872-

6895 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0304407620302384.

Zhao:2021:CAF

[ZD21] Anqi Zhao and Peng Ding. Covariate-adjusted Fisher randomization tests for the average treatment effect. Journal of Econometrics, 225(2):278–294, December 2021. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621001457.

Zhao:2024:NSG

[ZD24] Anqi Zhao and Peng Ding. No star is good news: a unified look at rerandomization based on p-values from covariate balance tests. Journal of Econometrics, 241(1):??, April 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407624000708.

Zhou:2024:SPS

[ZGHK24] Weilun Zhou, Jiti Gao, David Harris, and Hsein Kew. Semiparametric single-index predictive regression models with cointegrated regressors. Journal of Econometrics, 238(1):??, January 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407623002932.

Zhang:2024:TES

[ZGZ24] Jin-Ting Zhang, Jia Guo, and Bu Zhou. Testing equality of several distributions in separable metric spaces: a maximum mean discrepancy based approach. Journal of Econometrics, 239(2): ??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407622000859.

Zhu:2020:MSA

[ZHPW20] Xuening Zhu, Danyang Huang, Rui Pan, and Hansheng Wang. Multivariate spatial autoregressive model for large scale social networks. Journal of Econometrics, 215(2):591–606, April 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761930212X.

Zhu:2023:NGE

[Zhu23] Ke Zhu. A new generalized exponentially weighted moving average quantile model and its statistical inference. Journal of Econometrics, 237(1):??, November 2023. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623002269.

Zhao:2020:SWE

[ZHW20] Puying Zhao, David Haziza, and Changbao Wu. Survey weighted estimating equation inference with nuisance functionals. Journal of Econometrics, 216(2):516–536, June 2020. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619302404.

Zincenko:2024:EIS

[Zin24] Federico Zincenko. Estimation and inference of seller’s expected revenue in first-price auctions. Journal of Econometrics, 241(1): ??, April 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407624000800.

Zhang:2023:MAP

[ZL23] Xinyu Zhang and Chu-An Liu. Model averaging prediction by K-fold cross-validation. Journal of Econometrics, 235(1):280– 301, July 2023. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407622000975.

Zhang:2022:ODE

[ZLB22] Congshan Zhang, Jia Li, and Tim Bollerslev. Occupation density estimation for noisy high-frequency data. Journal of Econometrics, 227(1):189–211, March 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030440762030230X.

Zou:2022:ICM

[ZLLT22] Tao Zou, Wei Lan, Runze Li, and Chih-Ling Tsai. Inference on covariance-mean regression. Journal of Econometrics, 230(2):318–338, October 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407621001585.

Zhang:2022:VEH

[ZLTT22] Congshan Zhang, Jia Li, Viktor Todorov, and George Tauchen. Variation and efficiency of high-frequency betas. Journal of Econometrics, 228(1):156–175, May 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407620303778.

Zhu:2024:SAM

[ZM24] Changbo Zhu and Hans-Georg Müller. Spherical autoregressive models, with application to distributional and compositional time series. Journal of Econometrics, 239(2):??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407623000209.

Zhu:2022:CRB

[ZT22] Yinchu Zhu and Allan Timmermann. Conditional rotation between forecasting models. Journal of Econometrics, 231 (2):329–347, December 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407621002505.

Zhong:2021:MRE

[ZZ21] Xiaohan Zhong and Lin Zhu. The medium-run efficiency consequences of unfair school matching: Evidence from Chinese college admissions. Journal of Econometrics, 224(2):271–285, October 2021. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407620303675.

Zhang:2024:OMA

[ZZ24a] Xiaomeng Zhang and Xinyu Zhang. Optimal model averaging based on forward-validation. Journal of Econometrics, 237(2C): ??, December 2024. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S030440762200094X.

Zhou:2024:NBC

[ZZ24b] He Zhou and Hui Zou. The nonparametric Box–Cox model for high-dimensional regression analysis. Journal of Econometrics, 239(2):??, February 2024. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407623000568.

Zhang:2022:LBI

[ZZLL22] Xingfa Zhang, Rongmao Zhang, Yuan Li, and Shiqing Ling. LADE-based inferences for autoregressive models with heavytailed G-GARCH(1, 1) noise. Journal of Econometrics, 227 (1):228–240, March 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407620303742.

Zhang:2024:PSD

[ZZZ24] Yaowu Zhang, Yeqing Zhou, and Liping Zhu. A postscreening diagnostic study for ultrahigh dimensional data. Journal of Econometrics, 239(2):??, February 2024. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407622001877.